A Bibliography of Publications in Parallel Processing Letters

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 February 2022 Version 1.31

Title word cross-reference

(n, k) [CL13, MMCW18]. 0 [GAL96b]. 1 [APY06, GAL96b, MG09, RS98b]. 2 [CFG94, DFRC01, MCDB12, MCCW15, Pan97, PND02, RS98b, ST97]. 3 [Kru98, WGM^{+}10, YKLD14]. $5 [SCF01]. 5 [QY20]. 6 [HY18]. ^{1} [Sar92]. ^{2} [GM92, TKE^{+}08, Ram94]. p̄_{5} [NP04]. PLS [DM12]. d [Bar93, HS96, Pel93]. ⋄P [KW19]. ♢P_{mute} [FMR05]. e [HK95b]. ℓ [Wu95a]. g [LYW20, LZC^{+}21]. K [Ste12, ACM19b, AS97, GGHJ04, Gal96a, KSOK07, NM96, San03, VP99]. L [Vid06]. m [WW19]. N [Ste12, AS97, IN99, KSOK07, NM96, Sto96, WW19, XUZ02]. O(m) [NTHK17]. Ω(log n − k log k) [Gal96a]. p [XUZ02, ZN13]. p_{5} [NP04]. Q_{n,k,m} [CS13, LZG^{+}19]. t [IR16]. X + Y [AG98]. Z [QRR97].

-Ary [AS97, WW19, Bar93, NM96, Pel93, Sto96, VP99, KSOK07, Ste12]. -based [FMR05]. -Bijective [MCCW15]. -CMP [TKE^{+}08]. -Complete [DM12]. -Connected [QY20]. -cube [HK95b, KSOK07]. -Cubes [AS97, NM96, Ste12]. -Cycles [HY18]. -D [APY06, MCDB12, ST97]. -Dependent [GGHJ04]. -Diameter [ACM19b]. -Dimensional [RS98b, San03, WW19, HS96]. -Exclusion [Vid06]. -Extra [LYW20]. -Free [NP04]. -Good-Neighbor [LZC^{+}21]. -Graph [LZG^{+}19]. -M [Ram94]. -nodes [Wu95a].

-Polyhedra [QRR97]. -Process [IN99]. -Processor [XUZ02]. -reachable [CFG94]. -Resilience [IR16]. -Separator [PND02]. -set-based [KCH92]. -Shredders [QY20]. -Size [XUZ02]. -Star [CL13, MMCW18]. -value [ZN13].

1-edge [CH94]. 1-Fault [RV09]. 1-Interval [MST21]. 1-Latent [KY06]. 11-Step [BB09].

2 [Hag92, Sch97]. 2-CNF [Hag92]. 2.0 [CG98a, DLPW08]. 2012 [DMVT14]. 2D [Fra93a, YTL92]. 2D-PARBS [Fra93a]. 2D-Torus [YTL92].

3-Connected [KQY21]. 3-Extra [ZXY21]. 3D [BKT92, CM95, SC96]. 3D-grid [BKT92].

5 [WLR95]. 512 [Nov20].

Abel [Gla93]. Abel-Poisson [Gla93]. Abstract [CFM^{+}99, Her05]. Accelerating [MCR^{+}17, RJ21]. Acceleration [DTST15, IKMH09, MV08b]. Accelerator [LCS^{+}14]. Accelerator-Based [LCS^{+}14]. Acceptance [GHF10]. Access [CmL21, DWH^{+}06, GRS07, HZW08, NA10d, TB94, VAP94]. Account [CQSY20a]. Achieve [VMA11]. Achieving [WJ16]. Active [FKB^{+}99, GEL03, MAP^{+}19]. acyclic [CDZ96]. Adaptable [GKSP99]. Adaptations [DPM^{+}10]. Adapting [TT05]. Adaptive [Agr99, BIW11, CBV^{+}05, DDT17, HW06, JL05, KBFB01, KB05, KY06, KW06, LL21, PAKM08, Ran05, SLG04, SS13, SZJK11, Tan09, WBUW14, ABIM93, Dua93]. ADD [KW19, VRR20]. Adders [TA09b]. Adding [SKL10]. Addition [EW10]. Additional [YL10]. Additive [Kru98, LS91]. Address [Ald03, BAS96b]. Adjusting [BF05]. Advances [NA05, Qiu10]. Affects [DW07b]. affine [BFCD94, CRF95, KP94, Len92]. Agent [PRH^{+}03, VMA11]. Agent-Based [PRH^{+}03]. Aggregating [YS05a]. Aggregation [PA99]. Agreement [Mal97, Pel95]. Aided [VKS99, AD92a]. Alberto [Ano04c]. Algebra [AB08, Sol09b]. Algebraic [HCD^{+}19, RR91]. Algorithm [AAH14, AD02, AG98, BE03, BGK^{+}98, BSM^{+}16, BOV15, BEGK00, CC17, CLM^{+}16, DF99, DDS17, Dev05, DWS15, DT00, EW11, FS99, GMCC05, HA10, HHL^{+}07, HL98b, HT99, HLJ97, IN99, JS06, KKS00, KOKM14, KSAOK05, LMS98, LPP01, LN05, Mar98, MMMT12, MM16, NF06, NTHK17, OV06, Pag13, Per21, RRM15, Ray21, ST12, SMK98, SCJ^{+}10, TTV00, Tsa04, VKS99, Vid06, XX20, AS92, BNKS93, BK92, CM96, DJM91, DRS96, GI94, JR96, KDK^{+}93, KESH95, LM97, MB95, OW92, PB96, Pan94, Per94, PS93, ROJ94, SF91, SS92, TM93, TH94, Wan96, Wu96, Wue09, YDL91, YDL94]. Algorithmic [BDR^{+}99, OLC^{+}00, ZKmST18]. Algorithms [APY06, Ano94, ACK99, AB16, BF02, CMS99, CJN99, CG98b, DDT99, DGT98, DFRC01, Don07, FDZ99, GGH^{+}04, GRS07, GHS97, GK03, HHKM13, HKP^{+}98, HHH^{+}99, JW07, KLN04, KN06, KU09, LSF14, LMT10, MS99, MRRV07, MGBG07, MHKT05, MMAL06, MB98, NP04, PL11, PND02, PADB03, Raj02, RDK00, SKN10, SDLM18, VP98, VSS00, WSK16, XUZ02, BM95, BCD95, CLM93, Chl92, CT94, CF95, DH94, DDTV92, Dua93, FD92, HK95a, IR95, Nak95b, NR95, OSW93, PA95, PD95, Qia95, RP93, RSS95, SC96, XWF93, Zar97, SW93]. Aligned [YWJ03]. Alignment [AP05, IKMH09, LMS98, DR94, KDK^{+}93]. Alignments [FAH03]. All-Optical [CKP00, PSSV02]. All-To-All [APY06, FV99, CMS99]. Allocation [BDG97, HW06, IÖ98, PK04, TE04, AER94, CC93, HS93, KY96, PPRZ93, RB95]. Allocator [VH01]. Almost [De 04, KN06, Pel95]. Almost-Optimal [KN06]. Almost-safe [Pel95]. Alternator [Had08, Kul07]. Amdahl [CWR21]. Amdahl-like [CWR21]. Among [CH03]. analyse [AR96]. Analysis [AD02, Agr99, CG97, GLMM^{+}04, GRS07, GS97, HC02, JSZM08, KOKM14, KBG^{+}13, Kra01, LCS^{+}14, MMJ^{+}03, RvGG01, SBR^{+}17, SMH03, WR97, WGM^{+}10, YGM97, ZHW09, ZXYY20, Lin92, VP96, Wan96]. Analytical [AH00a, AH00b, ZSLC18]. Analytics [GGS^{+}15]. Analyzing [ARA20]. Anarchy [GLMM06]. And-Parallelism [Mar98]. Annealing [Pet03]. Anniversary [Akl11b]. Anonymous [CLM^{+}16, DDLV17, Tur10]. Ant [DHBL06]. Anticipated [LM00]. API [CH03]. Apostolico [Ano04c]. Application [BLY^{+}08, GB98, GM04, HW06, KS97, KC02, NA10b, ZHW09, Lin92]. Applications [AB08, AKP99, BS01, BRT09, BCL97, BDG97, CMS99, CLT13, CGS15, CLZC11, CFG04, FOCK15, FC05, FAMP12, Gar14, GP03, Ger02, GG96, HM01, HL03, IDS^{+}05, KBFB01, KLJ08, KHAM04, LF14, LFJ99, MTD98, MMJ^{+}03, NKCS03, NTG99, PT07, PT98, RBS15, SB12, VD03, ZMD15, Alb92, DM96, LOSZ93]. Applied [AIL16, CmL19a]. Applying [Mac12, TFTY05, TB18]. Approach [EADN06, Gam07, GS15, KWDS03, VLR^{+}03, CU96, FFSY93, RP93, Xue96]. Approaches [AL05, BMSW04, DHBL06]. Approximability [CDP^{+}07, SX02]. Approximate [HRH18]. Approximating [Hag92]. Approximation [AB16, BGK^{+}06, Tur10, BCD95, DV97, FR96]. Approximations [BBL04]. Arbitrary [BD02, CM97, DWS15, Had08, VRR20, ES92]. arc [RB93]. arc-disjoint [RB93]. Archetypes [Mas99]. Architectural [Gam07, DRS96]. Architecture [ATM01, BLY^{+}08, DDS17, Fuj08, HL98b, MCDB12, MAP^{+}19, MCR^{+}17, MIJ16, PL11, ZS08, AD92b, GS95]. Architectures [ADKT12, AFN11, BGJ10, BCL97, CG98b, CRY^{+}03, Mic98, PAKM08, CLZ92, CF91, DM96, DAB^{+}11, Nak95a, RR91]. Area [PADB03, PT07, CQS09]. Areas [CQS11]. arithmetic [GW94]. ARM [TH13]. ARMI [TSS^{+}06]. arounds [KRS93]. Arrangement [AA21, ZXLY20]. Array [CG97, DP03, ER99, LTZ98, MCDB12, QM98, RS97, YKLD14, AQRW95, Kri91, Qia95, Sto96, TM93]. Arrays [DT00, FUV99, GS03, HQP98, KM09, Sib02, UKY09, GW94, Gla93, IK96, Len92, Myo92, Oks95, OSZ91, PS93, RR91, Xue91]. Art [KB19]. Artificial [Mac12, YF18]. Ary [AS97, KSOK07, Ste12, WW19, Bar93, NM96, Pel93, Sto96, TM93, VP99]. Asher [Wan96]. Aspects [NA07b]. Assessing [BMRGR13]. Assignment [EB98, KU09, KN92]. Assist [AB08]. associated [DT96]. Association [TH13, ZN13]. Associative [AAH14, SW08, THAJ15]. associativity [DGJS95]. Assumption [MMRT06]. Asymptotic [GGV12, PT07]. Asynchronism [RR08]. Asynchronous [BIW11, GSS08, IR16, Pel06, YS05a, AJ96, Dor92, FMR05]. ATM [EFZ98]. Atmospheric [LF14]. ATMR [CP98a, CP98b]. Atomic [BF97]. Attached [DWH^{+}06]. Attacks [LLC11]. Augmentation [JS15]. Augmented [ZXY20]. Authenticated [CLL16, NA07c]. Authentication [CLH13, LLC11]. Author [Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21]. Auto [CFL12, HlMD21]. Auto-Tuning [CFL12, HlMD21]. Automata [Car07, DTST15, TLHL17, Tor09, Wue09, BH96]. Automated [HC02, LLD^{+}03, TLHL17]. Automatic [AB08, AGMM00, BF97, CL95, CJ01, EW13, HBL03, KKKL18, MRS04, MMJ^{+}03, NvG95, WLW11, Fea94, GAL96b]. Automaton [MG09, MS09]. Availability [Wol99]. Average [CQS11, LP98, BHPS95, Kop96]. Averaging [RJ21]. avoidance [Dor92, Dua92]. AVX [Nov20]. AVX-512 [Nov20]. Aware [ACCLS14, BMRGR13, BK08, DDT17, FOCK15, LO09, LN05, SYR13a, ZS08]. Axiom [CBC03].

Back [HL98b]. Back-Propagation [HL98b]. Background [SCJ^{+}10]. backtrack [KCH92]. Bag [FOCK15, OKL11, SB12]. Bag-of-Task [FOCK15]. Bag-Of-Tasks [OKL11, SB12]. Balanced [ST14, SŠZ04]. Balancers [KMAG01]. Balancing [CLR11, DEKS04, EMPF04, GC06, GT00b, MRRV07, ST12, VH00, CLZ92, XL94]. Bandwidth [CRY^{+}03, BKT95, Pel93, VP96]. Barnes [BGLM05]. Barrier [APSF01]. Barrier-Lock [APSF01]. base [GM92]. Based [AAH14, Ahn19b, ARA20, AGLS12, AKP99, CLL16, CGW05, CJ01, Deb97, DHBL06, DEKS04, DBK^{+}09b, EW13, ESSP01, FC05, FGK15, HHL^{+}07, HL98b, KSAOK05, LCS^{+}14, LZC^{+}21, LMT10, MCDB12, MMT17, MR01, OLC^{+}00, PFM^{+}09, PRH^{+}03, Sch05, SFHW11, Sch15, THAJ15, WLW11, Wu98, YKLD14, ZJSY20, ZZL^{+}21, Ahn19a, AD92b, AQRW95, BNW07, DM96, FMR05, KS96, KP94, KCH92, KP96, Qia95, ROJ94]. Basic [Sol09b]. Batch [TLH11, FL96]. Batched [Nov20]. BaTS [FOCK15]. Bayesian [MMAL06]. Be [VMA11]. Behavior [Cho98]. Being [Akl10]. Ben [Wan96]. Ben-Asher [Wan96]. Benchmark [ZHW09]. Benchmarking [CG01, DM96]. Benefits [BK08, ZTS^{+}16]. Benes [CM97]. Best [BET04, HlMD21, KBH03, KY96]. better [Wan96]. Between [BFG^{+}08, DLP03, GM03, GVCS14, LLD^{+}03, Trä09, Mac95]. Beyond [AB99, MST21]. BFS [SS92]. bibliography [Nak95a]. Biconnectivity [JS15]. Bicriteria [BBL04]. Bicube [LZC^{+}21]. Bicube-Based [LZC^{+}21]. bidimensional [ABIM93]. Big [MDAT17, VDL^{+}15]. Bijective [MCCW15]. Binary [CGR16, HMA99, KR98, QM98, XU99, XUZ02, SW93, UN92]. Bioinformatics [CCCV04, CFG04]. Biological [Ada07, CC17, IKMH09]. Biomolecular [FGK15, NA07b]. BIP [TW01]. BIP/Myrinet [TW01]. Bipancyclic [LZY18]. Bipartite [BMFU18, CJN99, Hag95, TMW20]. Bird [CS95]. Bird-Meertens [CS95]. Bit [Gla93]. Bit-level [Gla93]. Black [MCI^{+}01]. BLAS [MCR^{+}17, SS05]. Block [BOV15, MCDB12, OV06, VRT97, WBUW14]. Block-Cyclic [VRT97]. Block-Jacobi [BOV15, OV06]. Block-Parallel [MCDB12]. Blocked [FDZ99]. Blocking [HT98, WHTW10]. Blog [BLL12]. Blue [KBG^{+}13]. Bluegene [ABB^{+}03]. Bluegene/L [ABB^{+}03]. Bluetooth [TB18]. Boltzmann [OTK15, PKW^{+}03, ZHW09]. Boolean [TB94]. bottleneck [NM96]. Bound [CG98a, HPP99, CT94, CF95, Gal96a, Kop96, Mac93]. Boundary [BFG^{+}08]. Bounded [BEGK00, KW19, Sin96, CK93]. Bounds [DR00, GR11, HT21, MV04, Pel93, SB12, CL93, MSP95, RLG91, Wu95a]. Box [MCI^{+}01]. Bracha [Ray21]. Bracket [HT99]. Braided [LG97]. Branch [CG98a, HPP99, CT94, CF95]. Branch-And-Bound [HPP99, CG98a, CT94, CF95]. Breadth [Gre93]. Breadth-depth [Gre93]. Bridges [Dev05]. Brief [CQSY20a]. Broadcast [Ahn19a, BM97, DVWZ99, IR16, LP98, LM00, Mic98, Ray21, RV09, SW08, HK95b, Mac93, SV93, WV95]. Broadcast/Reduction [SW08]. Broadcasting [AD02, CMS99, DR00, DKP98, DD99, DV02, HL98a, KCPP00, Pel92, Qiu07, BKK94, BOSZ92, BOSW94]. Broader [TBV98]. Bruijn [AH98, Bar93, DA93, MV04, Pel93, RB93]. BSP [CG01, CS01, DPM^{+}10, FS99, Gav03, GS99, Hin03, KH02, Kru08, Lou01, TLHL17, Zav01]. BSP/CGM [FS99]. BSPGRID [Vas03]. BSPlib [Sui17]. BSR [GS94b, XU99, XUZ02]. buddy [JD92]. Budget [OKL11, WSK16]. Buffer [AP05]. Buffering [BF02, BCS05]. Buffers [Gon09]. Building [Gam07]. Bulk [BVHR13, CS03, Gav08, Sui17]. Bulk-Synchronous [CS03, Gav08]. Bus [ER99, HL98b, KCPP00, Wu98, FY93, HS96, OSZ91, PS93, Qia95, Sah95, TM93]. Bus-Based [HL98b]. bused [CL93]. Buses [HQP98, KM02, LTZ98, QM98, RS97, KRS93, LOSZ94, Sue95]. Butterflies [KKC98]. Butterfly [BDDP98, FU92]. Buttery [CQS16]. Bytecode [Gam03]. Byzantine [ABO06, BIW11, FMR05, IR16, Mal97, MT16, Pel95, Pel06, Ray21].

C [BGL03, BGLM05]. Cache [AFN11, CmL21, CP98c, KS97, KET06, PKW^{+}03, DGJS95, KS96]. CacheFlow [KET06]. Caches [WHTW10]. Caching [SBR^{+}17]. Calculating [CS95, Cos15b, HT99, ZN13]. Calculus [DP03]. camera [Ano98g]. camera-ready [Ano98g]. Caml [CF03, CRV08]. Can [Pan03, VMA11]. Canonical [QRR97]. Capture [SBR^{+}17]. Card [WJ16]. Cards [SCF01]. Carlo [HUZ06, MHC^{+}04]. Carry [TA09b]. Cartesian [AA21, ACM19a, BA95, Wan96]. Cascades [BLL12]. Case [BGLM05, CFG04, DBK09a, HTHH05, MG11, NDFM07, VLR^{+}03, ZTS^{+}16, BHPS95]. Case-Study [HTHH05]. Causal [BM97]. CCCs [SŠ03]. Cell [Dec21, DBK^{+}09b]. Cellular [Car07, DTST15, GT00a, GMCC05, JW07, MG09, MS09, Tor09, Wue09]. census [BNKS93]. Center [HY19, LZS^{+}18, SDLM18, XLP12, ZMZZ20, ZMZZ21]. Certain [RRM15]. Certified [CLL16]. CFD [Tan09]. CFinder [PPV12]. CGM [FS99]. Chain [BRT09, Wue09, ZN13]. chain-rules [Wue09]. Chains [ACK99, SDLM18]. Challenges [LGHB07]. Channel [Dua92, GR11, GHS97]. Channels [BL08, KW19, VRR20]. Characteristics [HZW08, Vir93]. Characterization [EFZ98, DDR96]. Characterizing [CLT13]. Check [PT07]. Checking [RvW98]. Checkpoint [SWK^{+}13]. Checkpoint/Restart [SWK^{+}13]. Checkpointing [Ahn19a, CCR11]. Chemical [Cos15b, MBT15, ZKmST18]. Chief [Ano01c]. Chip [BLY^{+}08, CmL19a, TKE^{+}08, ZJ09]. Chip-Multiprocessor [TKE^{+}08]. Chips [Jes08, PAKM08, ZSLR21]. Cholesky [CHQ96, KESH95]. Chordal [NP04, PND02, KP96]. chordless [CLM93]. Christmas [RRM15]. Circuit [IÖ98, PS98, BHPS95]. Circuit-Switched [IÖ98, PS98]. Circuits [SX02, RB93]. Circulation [KOKM14]. CIRM [BET04]. Class [LSW97, XLP12, XX20, ZXLY20, BE96, HL96]. classes [Dua92, RP93]. Classical [NA07c]. Classification [AAH14, THAJ15]. Cleaner [JS18]. Client [GM04]. Client-Server [GM04]. Clients [ABO06]. Clique [AKPSR16]. Clock [APSF01, PT97]. clocks [ADG91]. close [GS94a]. Closed [ANS12]. Closure [PL11, Aln94, GMR92]. Cloud [DJMN13, KB19, WSK16]. Clouds [DT13, DT15, MJGR09, OKL11]. Cluster [DEKS04, GM03, GKSP99, PRH^{+}03, RšT06, VLR^{+}03]. Cluster-Based [DEKS04]. Clustered [Trä09, TDS99]. Clustering [BNW07, CC17, DWH^{+}06, Fox13, HA10, PPV12, KESH95]. Clusters [Aba04, BLMR03, CG01, Dan01, DT01, DT13, DT15, FOCK15, HCD^{+}19, Kra01, PA99, WHTW10]. CM [WLR95]. CM-5 [WLR95]. CML [Deb97]. CMP [TKE^{+}08]. CNF [Hag92]. Co [GLMM^{+}04, SYR13a]. Co-Regulated [GLMM^{+}04]. Co-Schedules [SYR13a]. Coarse [Box09, Box21, CD99, SSK^{+}18]. Coarse-Grid [SSK^{+}18]. coarsest [GI94, Gal96a, JR96]. Code [EW13, Gam07, KLH^{+}14, KKKL18, van94]. Codes [Agr99, GKS^{+}14, KLN04, KN06, KLH^{+}14, PT07, PKW^{+}03, WHTW10, Sto96, TM93]. Codeword [Gar14]. Coherence [CP98c]. Collaborative [RGM06]. Collecting [FDFR93]. Collection [AKP99, PK04]. Collections [ZTS^{+}16]. Collective [BIC05, HSL09, Trä09, WLR95]. Colony [DHBL06]. color [PB96]. Colored [ACM19a]. Coloring [ACM19a, FR96]. column [Sue95]. Columns [Jai99]. combination [Gri92]. combinations [ES92]. combine [BH93]. Combined [Mak09]. Combining [DSV97, Pet03]. Common [BFG^{+}08, CS13, Per21]. Communicating [BW16]. Communication [Ahn19a, Ahn19b, APY06, ATM01, AP05, BBF05, BIC05, BGPT00, BCS05, CKP00, CFM^{+}99, CCA01, CG98b, DFRC01, DGVG96, Gon09, IT02, IDS^{+}05, JSZM08, KM09, LDC08, NA07c, Nit05, PSSV02, PS98, PAKM08, Pur06, Res97, RšT06, ST12, SMK98, SFHW11, TB18, TSS^{+}06, Trä09, VRT97, YGM97, AER93, DDT95, GC95, LJ95, NPT95, Qia95, RLG91, WLR95]. Communication-Efficient [DFRC01]. communication-intensive [Qia95]. communication-time [LJ95]. Communications [BGK00, FV99, GM03, GBS^{+}07, HSL09, SŠZ04, Wu98]. Communities [BLL12, PPF12]. Community [FAMP12, NTV12, TMKS16]. Compact [BF02, CJ01]. Comparative [CDM^{+}03, Fra93a]. Comparing [DW07a]. Comparison [FV99, LL21, Mer15, MDAT17, KRS93]. Compensation [MS05]. Competitive [BDG97, KRS15, Fio93]. Compilation [BGJ10, Kre97]. Compile [KS97, KS96]. Compile-Time [KS97, KS96]. Compiled [JSZM08]. Compiler [Bra00, LF14, Qui00, RvW98]. Compiling [ABNP00, PP97, van94, Mon94]. Complementary [TMW20]. Complete [ACM19a, BGPT00, DM12, EFZ98, Sch09, BCD95, Gre93]. Complexity [AKPSR16, BBF05, BRT09, CCQ13, CDP^{+}07, GPS03, OLC^{+}00, Pur06, Ste12, TBV98, TX98, ZKmST18, AER93, CDZ96, NPT95]. Component [KBFB01]. Components [GSS08, SKN10]. Composition [FC05, Pag13, ZTP^{+}08]. Compressed [De 04, HlMD21]. Compression [HS97, KKS00]. Computation [Akl03, Alb92, ANS12, DDS17, FUV99, HW07, Jon15, Mac12, Nov20, OKH^{+}02, Qiu10, Sta12, Vas03, ZSLC18, BIL92, GS94b, LP94, NvG95, RV96, SBP94]. Computational [ABG02, AVD01, Cam08, CLR11, DJMN13, DT01, MKA98, Ste12, LOSZ94, HMA99, Wol99]. Computations [Akl06b, Bra00, BA01, GK14, HlMD21, KM09, LF97, SHBG14, WJ12, YWJ03, BOSW94, CC93, JJ96, Akl06a]. Compute [DBK^{+}09b, RGM06]. Computer [Akl06c, NA07a, AD92a, AD92b, Ano98g, Chl92, Sah95]. Computers [Hua06, KCPP00, LPP01, DJM91, FY93, MMMS94, Zia95]. Computing [AGLS12, BDH^{+}08, BCMC^{+}04, BDR^{+}99, CSK00, DMVT14, DT01, DT13, DT15, DM12, FVR20, FGK15, GLMM^{+}04, GS94b, GK13, HQP98, HT94, IDS^{+}05, KB19, KBGE07, KWDS03, LMZ99, LOSZ94, LY07, MG11, MB98, NA07b, SCF01, SS13, TDS99, VMA11, VDL^{+}15, BNKS93, NR95, Pan94]. concave [LP94]. concept [Dua92].

Concurrency [GRS07, Jes06]. Concurrent [BGJ10, Deb97, GR99, HHL^{+}07, Sch15, SCJ^{+}10, Vid06, ZTS^{+}16]. Condition [GR99]. Conditional [CP12, CL13, CQS16, LYW20, Wan96]. Conference [DMVT14]. Configuration [CSK00, Sch09]. Conflict [VLL^{+}91]. Conflict-free [VLL^{+}91]. Congestion [DM12, RS98b]. Conjugate [SSK^{+}18]. Connected [AKPSR16, CZ20, KQY21, LPL20, LZY18, MST21, QY20, SKN10, VRR20, Chl92, FU92, KY96, Kri91]. Connection [MCCW15]. Connectivity [AS19, Bur21, CQSY20a, HY19, kLCL20, LYW20, MWZ19, Nin20, TMW20, WW19, WXW20, ZXY20, ZXY21, Aln94, YDL94]. Conquer [HL00, Tis01, HL96]. Conqueror [GKS^{+}14]. Consensus [BIW11, MR01, MPR18, FMR05]. Consistency [GR99, HK16]. Console [ESSP01]. Constant [ADS91, DDLV17, DMS99, MB98, SDLM18, MB95, Pan94, PS93]. Constraint [HKS01, OII12]. Constraints [WSK16, CRF95]. construct [SS92]. Constructing [GM92, JS18, KW06]. Construction [BSM^{+}16, BHQ92, CRF95, DZ13, DVWZ99, EFZ98, KR98, KLN04, KN06, BH96]. constructions [CC92]. constructor [Sie96]. Consumers [LLD^{+}03]. Consumption [dSJR^{+}15]. Contention [AFN11, BK09, CGR16, NPT95]. Contention-Friendly [CGR16]. Context [BW16, Sch15]. Context-Based [Sch15]. Context-Free [BW16]. Continuous [FUV99]. Contour [MB98, Gla93]. Contractible [KQY21]. Contraction [MHKT05, QY20]. Contraction-Critical [QY20]. Contrast [MTD98]. Control [Cam08, CmL19a, Kme14, MAP^{+}19, NA10d, OKL11, FFSY93, SF91, Xue91]. Controlled [GRS07]. Controllers [SKL10]. Convection [ZLL04].

Convection-Diffusion [ZLL04]. Convergence [AVD01, DWS15]. Convertible [CLL16]. Convex [BOSZ92, RB95]. Convolution [YF18]. Cooperative [DHBL06, GK13, RR08, ZTP^{+}08]. Coordinates [PPF12]. Coordination [GSS08, ZTS^{+}16]. Coping [NA10a]. Copy [GS97, Yai08]. Copying [PK04]. Core [AFN11, BDH^{+}08, DZ13, Jes08, MIJ16, SKL10, Sol09a, Sol09b, UVJ11, DAB^{+}11]. Cores [BBG^{+}11]. Correctness [CP98b, O’D94]. Coscheduling [CCA01]. Cost [BGK03, HC02, LM00, NF06, SMH03, ZTS^{+}16, DGJS95, Gal96a, PRS95]. cost-effective [PRS95]. Costs [BCS05, Cam08, GS97]. Count [EFZ98]. Counterexamples [Akl06c]. counters [AR95, ADS91]. coupled [CF91]. Cover [Tur10, BCD95]. Coverability [BW16]. Coverage [DGB10]. CPU [GGV12, Wol99]. CPU/GPU [GGV12]. crash [GV93]. crash-tolerant [GV93]. Cray [DBK09a]. CRCW [Fra93a]. CRCW-PRAM [Fra93a]. Create [MCI^{+}01]. creation [SJL94]. CREW [LO93]. Critical [QY20]. Cross [SS13]. Cross-Paradigm [SS13]. crossbar [SF91]. Crossed [CP12, HLH^{+}17]. crosstalk [BAS96b]. Cryptographic [NA10d]. Cryptography [NNA10]. Cube [HLH^{+}17, KSOK07, LC14, LH02, Wu98, FU92, HK95b]. Cube-Based [Wu98]. cube-connected [FU92]. Cubes [AMC18, AS97, CP12, DZ13, Ste12, ZXY20, NM96]. Cubic [KQY21]. CUDA [Fuj08, WJ12]. Current [SFHW11]. Curves [ANS12]. Cut [Dev05, ABIM93]. Cut-Nodes [Dev05]. cut-through [ABIM93]. Cutting [BMSW04]. Cutwidths [MV04]. Cycle [BDDP98, MMT17]. Cycle-Based [MMT17]. Cycles [ACM19a, AS97, HY18, HKP^{+}98, CLM93, FU92]. Cyclic [VRT97].

D [Kru98, TKE^{+}08, APY06, DFRC01, MCDB12, MG09, Pan97, ST97, WGM^{+}10, YKLD14]. DAG [CJ01, DDT99, DT96]. DAGs [CSBS95]. DARPA [LY07]. Dashboard [RGM06]. Data [Agr99, ADKT12, Ser01, ABNP00, AVD01, BLMR03, BGL03, BGLM05, BIL92, Bra00, Cam08, CDM^{+}03, CSFK08, DDT17, DZ13, DT13, DT15, FKB^{+}99, Gav08, GGS^{+}15, GB98, GKS08, Guo01, GS97, HA10, HZW08, HY19, KKKL18, KET06, LMZ99, LF97, LTB01, LZS^{+}18, MP93, MRRV07, MDAT17, Mon94, NKCS03, PP01, RJ21, RBS15, Sah95, SBR^{+}17, TKE^{+}08, THAJ15, Tho03, XLP12, XWF93, YGM97, ZSLR21, ZMZZ20, ZMZZ21, CU96, GAL96b, GDC94, GHSJ94, HT94, LV96, NvG95]. Data-Center [XLP12]. Data-Driven [NKCS03, TKE^{+}08]. Data-Flow [Ser01]. Data-Parallel [ABNP00, BGL03, GB98, Guo01, GS97, RJ21, CU96, LV96]. Database [NKCS03]. Databases [DEKS04, FKB^{+}99]. Dataflow [CG97, MDAT17]. Datagrid [Sch05]. Dataset [PPF12]. DatTeL [BGL03]. DCell [LZS^{+}18]. Deadline [WSK16]. Deadlock [Di 00, KSAOK05, SV00, WS03, Dua92, Dua93, FD92]. Deadlock-Free [SV00, WS03, Dua93]. deadlocks [Dor92]. deBruijn [Sch93]. Debugging [Sui17]. December [Akl14a]. Decentralised [GGV12]. Decentralized [TT05]. Decidability [AMN12, KM09]. decision [SW93]. Declarative [GSS08]. Decoders [De 04]. Decoding [XU99, XUZ02]. Decoherence [NA10a]. Decomposable [SŠZ04]. Decomposition [BDDP98, DHBL06, LSF14, OK20, YKLD14, DV95, IR95, NvG95, Sch93]. Decompositions [Nov20]. Decoupled [SAJ10, Tho03]. Deep [YF18]. Defeats [NA07a]. defined [NR95]. Degree [CK93, DLP03, Had08]. Degree-bounded [CK93]. Delaunay [LPP01]. Delay [BNW07, GS15, MBT15, ADS91, BHPS95, NPT95]. Delayed [PK04]. Delays [BGPT00]. Demand [CRY^{+}03, YGM97]. Demand-Driven [YGM97]. demon [TH94]. Denotational [Guo01]. Dense [DP03, Fuj08, IT02, WJ16, MP93]. Density [DW07b, HTHH05, PT07]. Departmental [GL05]. Dependence [KBFB01]. dependencies [HT94]. Dependency [KRS15, JJ96]. Dependent [GGHJ04]. Deployment [CFG04]. depth [Gre93]. Derivation [GB98, HC02, MHKT05, SMK98]. Deriving [Nit05, RDK00, RR91]. Description [Sch05, WLW11, ROJ94]. Design [BM97, BLY^{+}08, CS03, CG98b, Gar14, GEL03, GM04, HlMD21, MCR^{+}17, SZH14, YKLD14, Dua93, FFSY93, PRS95, SC96]. Designing [PMW^{+}12, Tsa04, Qia95]. designs [CH94]. Desktop [BFG^{+}08, CSFK08]. Destinations [ST97]. Detection [DMS99, GKK14, HlMD21, Mar98, ZYPGD08, FD92]. Detector [VRR20]. Determination [CFG04]. Determining [Alb91]. Deterministic [AJ96, DZ12, DFRC01, GS99, HPP99, RS97, RvGG01, DH94]. Developing [Ada09, Pan03, VD03]. Development [ABG02, CQSY20a, ESSP01, Hin03]. Device [DWH^{+}06]. Devices [LC14, SMK98, SCF01]. devising [Fla97]. DHcube [ZZZY21]. Diagnosability [CQS16, LYW20, WXW20]. Diagnosis [EADN06, LL21, ZZL^{+}21, DTLA93]. Diagram [DFRC01]. Diagrams [Cos15b, SW93]. Diameter [ACM19b, HL99, SH99, Sri96]. Diamond [GVCS14]. Dictionary [HS97, CC95]. Difference [Pur06]. Different [CH03]. Differential [OK20]. Diffusion [BLL12, EMPF04, ZLL04, XL94]. digital [ADG91, Gla93]. Digraph [PL11]. digraphs [CFG94]. Dilation [EM98]. Dimension [BEGK00, KW06, SŠ03]. Dimension-Adaptive [KW06]. Dimensional [BCMC^{+}04, BMSW04, Car07, LSW97, RS98b, San03, Sib02, WW19, CL93, CR95, HS96, KW92, MB95, OSZ91]. Dimensions [YL10]. dining [Per94]. DIOGENES [PRS95]. Direct [KB11]. Directed [FVR20, Sri96]. Directions [CQSY20a, KB19]. Directives [LF14]. Directory [CP98c]. Discovery [ZTP^{+}08]. Discrete [CLL16, CHQ96]. discrete-time [CHQ96]. Discriminating [Lin92]. Disha [KSOK07]. Disjoint [HHKM13, Kru98, YMZL04a, YMZL04b, RB93]. disjunction [BHPS95]. Disk [DZ13]. Dispel [Akl06c]. dispersal [DMP94]. Distance [ACM19a, GGH^{+}04, GZL^{+}20, HHH^{+}99, BIL92, DV97, PA95]. Distance-Two [GGH^{+}04]. Distances [CQS11]. Distributed [Agu02, AL05, Ahn19a, Ahn19b, ATM01, ABNP00, AKPSR16, BCBB09, BCQO99, Cho98, CRY^{+}03, DTLA93, FD92, GP03, IT02, Jes06, KOKM14, KWDS03, LPP01, LO09, Lou01, MMJ^{+}03, NA05, NTG99, OL11, PADB03, SBR^{+}17, SWK^{+}13, SB15, SCF01, VD03, WLR95, ZSLC18, AER94, GHSJ94, JJ96, MNR96, NvG95, PPRZ93, RLG91, Sah95, SS92, TH94, van94]. Distributed-Memory [IT02, GHSJ94, NvG95]. Distribution [Agr99, CSFK08, LF97, NA07c, NA10b, NNA10, Fea94, GAL96b]. Distributions [DP03, VRT97]. Distributively [Aba04]. Distributively-Owned [Aba04]. Divide [HL00, Tis01, HL96]. Divide-and-Conquer [HL00, HL96]. Divisible [CWR21, Rob11]. division [AD92b]. DLP [Sol09a]. DNA [CGW05, Gar14]. DNA-Based [CGW05]. Do [CT94, LJ95, CRF95]. Documents [MRS04]. Does [Ada09, KE00]. Domain [KSHL14, OK20]. Domain-Specific [KSHL14]. domatic [RP93]. Dominating [DWS15, HHKM13]. Domination [LZG^{+}19]. Double [FNP17, FGK15, HLJ97]. Double-Ended [FNP17]. Doubly [BIW11]. Doubly-Expedited [BIW11]. Down [CS03]. Downloads [PADB03]. Drawing [XU99]. Drawings [TTV00]. Driven [KET06, NKCS03, PMW^{+}12, PT98, TKE^{+}08, Tho03, YGM97]. DSL [GDTF17]. DTML [Sar92]. Due [LM00]. duplex [KRS93]. Duplication [BGK00]. During [KHW05]. Dust [CDM^{+}03]. Dynamic [BOV15, BL08, BDG97, CmL19a, CLR11, DGB10, DR00, DT00, DV02, Dol98, GM03, Hua06, Jes08, KMAG01, MG11, MCI^{+}01, MMJ^{+}03, PAG08, Ran05, SB15, SLG04, TFTY05, FFSY93, Myo92, PLR^{+}95, VP96]. Dynamically [WBUW14, Nak95a].

E-BaTS [FOCK15]. ear [IR95]. Early [TE04]. ECM [KSHL14]. Eden [HHOM02]. Edge [ACM19a, DDLV17, HL99, LPL20, MWZ19, MCCW15, RS98b, VSS00, ZXY20, CH94, FR96]. Edge-Colored [ACM19a]. Edge-Congestion [RS98b]. Edge-Connected [LPL20]. Edge-Connectivity [MWZ19]. Edge-Faults [MCCW15]. Edges [FVR20, KQY21, Pan97, Kop96, BFG^{+}08]. Editor [FAMP12, Ano01c, Ano06g, Ano08h, Ano09f, JKRW08, JKRW09, Qiu10, Sta12, TA09a, WKRJ10, WKR11]. Editor-in-Chief [Ano01c]. Editorial [Akl11a, Akl11b, Akl12d, Akl12a, Akl12b, Akl12c, Akl13a, Akl13b, Akl13c, Akl13d, Akl14a, Akl14d, Akl14b, Akl14c, Akl15a, Akl15b, Akl15c, Akl15d, Akl16a, Akl16b, Akl16c, Akl16d, Akl17a, Akl17b, Ano97a, Ano97b, Ano97c, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano01b, Ano02b, Ano02c, Ano03b, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano06e, Ano06b, Ano06c, Ano06d, Ano06f, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano09b, Ano09c, Ano09d, Ano09e, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, BET04, Cos93, Cos95, Cos96a, Cos96b, Cos97, Cos99, Cos00, Cos01, Cos03, Cos04, KBH03, YS05b]. Editors [Ada15, DMVT14, DT11, DT13, DT15, GK17, GK14, HGH12, WKR13, WKR14]. Effect [HSL09]. Effective [CSK00, HKP^{+}98, CR95, Dua94, PRS95]. effectiveness [CF95]. Effects [HUZ06]. Efficiency [CRY^{+}03, IR16]. Efficient [Ahn19a, Ber92, BEGK00, Box09, BH93, CLM93, CDM^{+}03, Dan01, DFRC01, EB98, GM03, GS99, GS03, GMCC05, HPP06, HCD^{+}19, HK95b, HHH^{+}99, JS18, KLN04, KH02, Mak09, Mal97, MCR^{+}17, NRP^{+}94, OL11, PA95, PA99, Rag98, Raj02, RBS15, SYR13a, SV93, VH01, CH96, Fra93b, GV93, Pan94, PPRZ93, RSS95, YDL94, Zar97]. EGEE [NDFM07]. Elastic [KB19, MRS04]. Election [DDLV17, SM20, AJ96, Tel95]. Electromagnetic [MKA98]. Electron [BCMC^{+}04]. Element [RGM06]. Elements [BB09, GLMM^{+}04, MB98, ES92]. Elimination [SMH03, PS93]. Elliptic [OK20]. Elusive [HoHqZS20]. Embeddability [YL10, Lat91]. Embedded [KCPP00]. Embedding [AS97, Bar93, BBBL04, EM98, HP92, KMY97, LH02, MCDB12, MLY99, QM98, RRM15, RS98b]. Embeddings [KKC98, Lav02]. Embodied [HW07, Mac12]. empty [OSW93]. Emulating [MS01a]. emulation [LO93]. Enabled [JSYD07, SPA04]. Enabling [GM04]. encodings [BAS96b]. Encounter [NA10c]. Encryption [CLL16, Wue09]. Ended [FNP17]. Energy [AB16, BMRGR13, FOCK15, LO09, LN05, OL11, TB18, VDL^{+}15]. Energy-Aware [BMRGR13, FOCK15]. Energy-Efficient [OL11]. Engine [SZJK11]. Engineering [Tan09]. Enhanced [ZXY21, LOSZ93]. Enhancement [MTD98, NNA10]. Entanglement [NA10b]. Enumeration [TMKS16]. Environment [ABNP00, CCCV04, CSK00, ESSP01, GKSP99, MHC^{+}04, RGM06, Vio94].

Environments [Agu02, AKP99, HUZ06, IDS^{+}05, MS01a, MMMT12, MM16, Pel06]. Enzyme [FGK15]. Enzyme-Based [FGK15]. equation [Sta95]. Equations [BCQO99, OK20, ZLL04, CHQ96, Len92, Mon94, Oks95, Xue91]. Equilibria [BCBB09, DM12]. Erasure [PT07]. EREW [DH94, DF99, De 04]. Eridis [OL11]. Errata [Akl06a]. eskimo [Ald03]. Establishing [BS01]. Estimation [GV03, MMAL06, OKL11, KP94, PBKP92]. Euclidean [BIL92, PA95]. Euler [CDSS93]. Euro [KBH03]. Euro-Par [KBH03]. Evaluating [Cam08, FZL^{+}16]. Evaluation [BDH^{+}08, DW07a, Her05, HRH18, KSHL14, Kru08, LG97, LGCH99, LZC^{+}21, Lou01, LZS^{+}18, MMAL06, OTK15, SYL95, SS05, Sol09b, SAJ10, Tho03, ZZL^{+}21, Alb92, ABIM93, SSSM93, YTL92]. Evaluations [KW06]. Even [SŠ03]. Eventual [MMRT06]. Eventually [VRR20]. Evolution [SZH14]. Evolutionary [MGBG07, SZH14, GG96]. Evolve [GMCC05]. Exact [KU09, GM92]. Example [LF14]. ExaStencils [KKKL18]. Exchange [MV04, CC93]. Exchanged [Nin20, ZXYY20, ZZL^{+}21]. Exchanges [Box09]. Exclusion [IN99, KM02, KY06, Vid06, BK92, CM96]. Exclusive [BKK94]. Exclusive-Read [BKK94]. Execution [ATM01, FC05, Gam03, GPS03, GK13, GS03, GKS08, HRH18, LSW97, PT98, RBS15, SS13, GDC94]. Execution-Driven [PT98]. Exhaustive [TMKS16]. exist [CT94, LJ95]. expander [Hag95]. Expedited [BIW11]. Experiences [LFJ99]. Experiment [SS13]. Experimental [BIC05, GJQ09, Kru08, LO09]. Experimenting [Ald03]. Experiments [GKS^{+}14, Mas99]. Explicit [Pur06, SFHW11]. Explicitly [CG97]. Exploiting [CmL21, RRS12, Sol09a]. Exploration [PT07]. Explorations [Hea07]. Exponentiation [EW13]. Expression [JSZM08, Per21]. Extended [CG98a]. Extending [LCS^{+}14, Mar03]. extensible [Sri96]. Extensions [HLH^{+}17, KE06, FK96]. external [GS94b]. Extra [LYW20, ZXY21]. Extraction [CJ01]. Extrapolation [CLT13]. Extreme [CCR11, LBD^{+}14]. Extreme-Scale [LBD^{+}14].

FAB [YWJ03]. facility [LHCT96]. Factor [LLC11, BH96, CHQ96]. factorisation [PBKP92]. Factorization [FV16, GCP99, DDT95, KESH95, MP93]. Factorizations [ZSLC18]. Facts [Sza12]. Faculty [KMY97]. Failure [GRS07, NDFM07, VRR20, ZYPGD08]. Failure-Sensitive [GRS07]. Failures [MT16, MLY99]. Fair [CRV08]. Fairly [GGS^{+}15]. Fairness [GS15]. False [SX02]. Family [UKY09, Fio93, HP92]. Fantasy [Syr13b]. Far [Sch09]. Farm [PK08]. Fast [BM95, CH96, CCA01, CGR16, DDS17, DS97, HK95a, HPP99, HRH18, LSF14, RV96, Aln94, AS94, DRS96, GHSJ94, Sta95]. Fastest [CJN99]. Fat [ZJ09]. Fat-Meshe [ZJ09]. Fault [ARS93, CC11, EADN06, GKK14, GK13, GP96, HLH^{+}17, KHAM04, LH02, LNLE00, MLY99, MV08a, NTG99, Res97, RV98, RV09, SV00, SH99, VMA11, Wu98, ZXY20, BK92, CM95, CH94, DTLA93, RT92]. Fault-Hamiltonicity [HLH^{+}17]. Fault-Tolerance [CC11, NTG99]. Fault-Tolerant [LH02, MLY99, MV08a, Res97, RV98, SV00, Wu98, ZXY20, BK92, CM95, CH94]. Faults [DR00, DV02, Mal97, MCCW15, DMP94, Pel95]. Faulty [ABO06, DPS00, IÖ98, KOKM14, Ste12, Pel92, Wu95a]. Federated [GMK13]. Feedback [Akl03, TFTY05]. Feynman [FGK15]. FFT [APY06, EW11, GB98]. File [DWH^{+}06, GMK13, PADB03]. Filters [KE06]. Final [Kme14]. Finding [BGK^{+}98, Dev05, DW07a, DW07b, SDLM18, Wu00, CDSS93, CLM93]. fine [CSBS95, GG96]. fine-grained [CSBS95, GG96]. Firing [UKY09]. First [DMVT14]. fit [KY96]. Fixed [Kme14, LJ95]. fixed-processor [LJ95]. Flexible [GBS^{+}07]. Flight [CF03]. Floating [MCR^{+}17]. Flow [ADKT12, Ser01, DDT17, FGK15, YGM97]. Flows [ZSLR21]. Fluids [LBD^{+}14]. flush [AR95]. Folded [Nin20, Lat91]. Follow [Ada09]. Forced [SZH14]. forest [OW92]. Foreword [GL02]. Form [Rob11, RV96]. Formal [BA01, GP03, Gav03, GB98, Sch15]. formalism [CS95]. Formalisms [HL02]. Format [HlMD21]. Formed [Tsa04, VP98]. Fortran [Sch97, EB98]. Foundations [Deb97]. Four [EM98, UKY09]. Four-State [UKY09]. Fourier [NA10a, GHSJ94]. FPGA [ZJSY20]. FPGAs [Dec21, TKE^{+}08]. Fractional [MMCW18, ZZZY21, ZMZZ20, ZMZZ21]. Fragile [MCDB12]. Framework [DDT17, FC05, GDDM17, LC14, RBS15, TH13, Tor09, VD03]. Frameworks [KWDS03, MDAT17, Qui00, THAJ15]. Free [BW16, FNP17, KY96, Kme14, MMRT06, NP04, PT97, SV00, ST14, WS03, Dua93, VLL^{+}91]. Friendly [CGR16]. FT [LNLE00]. Fully [AKPSR16, TH94]. Fully-Connected [AKPSR16]. Function [KW06, GI94, JR96]. Functional [BL08, DRS96, Gav03, GL05, HL02, LCK02, Lou01]. Functions [AH00a, AH00b, Akl03, Nit02, BNKS93, Gal96a, NR95]. Further [HUZ06]. Fusing [ATM01]. Fusion [PMW^{+}12, WBUW14]. Fusion-Driven [PMW^{+}12]. Future [CQSY20a, KB19]. Fuzzy [AsS19].

G [LTB01]. Game [AsS19, BCBB09, YF18]. Games [DM12]. Garbage [AKP99, PK04]. Gate [FGK15]. Gates [SX02]. Gathering [BGK^{+}06, MST21, CF93]. Gene [SZH14, KBG^{+}13]. General [BGJ10, NTHK17, SAJ10]. General-Purpose [SAJ10]. Generalising [Xue96]. Generalization [CS13, KP96]. Generalized [BCQO99, HY19, ZXYY20, ZZL^{+}21, DA93]. generate [TM93]. Generating [BEGK00, GS94a, Her05, Sto96, Tsa04, AS92]. Generation [EW13, LBD^{+}14, Pag13, SSK^{+}18, TLHL17, VP99, VRT97, WLW11, CL95, ES92]. Generations [EW13, KBG^{+}13]. Generative [KB05]. Generator [KOKM14, KKKL18]. Generic [Ger02, GB98, Nit05]. Genes [GLMM^{+}04]. Genetic [CG98b, LN05, VKS99, DDTV92]. genome [ZN13]. genome-wide [ZN13]. Genomic [SZH14]. Gentle [GSS08]. Geometric [Akl06a, Akl06b, DPS00]. GFFS [GMK13]. Given [EFZ98]. Global [ATM01, CBC03, CLH13, PS98, AR95, BH93, GMK13]. Go [KE00]. Gödel [AMN12]. Gomoku [YF18]. Good [LZC^{+}21]. Gossiping [FPP98, HS96, HKP^{+}98, RS96, SŠ03, FY93, LR93]. Göttfert [ST12]. GPGPU [BCL12]. GPU [EW11, EW13, GZW16, GGV12, LMT10, MG11, MV08b, SKN10, SHBG14, WJ16]. GPU-Based [LMT10]. GPUs [DZ12, HCD^{+}19, IKMH09, LF14, Mer15, WJ12]. graded [FDFR93]. Gradient [SSK^{+}18]. Gradual [BM97]. Grained [Box09, Box21, CD99, CSBS95, GG96]. Grammar [BW16]. Granularity [LTB01]. Graph [AA21, Agu02, ARA20, BGK^{+}98, Bur21, CS13, DWS15, GHJ^{+}08, LZG^{+}19, LGHB07, SKN10, TMW20, ZMD15, Aln94, AR96, BKT92, CL95, GS95, IR95, PB96, Pel93, RB93, Sch93, Sri96, SS92]. Graph-Optimization [ZMD15]. Graphics [AIL16, EW10]. Graphs [Ada09, AH98, ACM19a, BMFU18, CZ20, CJN99, CL13, CQS16, DMNP11, EMPF04, FVR20, GZL^{+}20, HHH^{+}99, KRS15, KQY21, LP98, LYW20, MMCW18, MWZ19, MST21, MV04, NTHK17, NP04, PND02, PPF12, QY20, WXW20, ZXLY20, Bar93, CLM93, CQS09, GP96, HP92, Hsu93, Kop96, LH95, PS93, RP93, Wol95, YDL94]. Gravity [GT00a]. Gray [Sto96, TM93]. Greedy [ST97]. Grid [BCMC^{+}04, CFG04, GLMM^{+}04, GK13, GM04, KW06, LLD^{+}03, MHC^{+}04, Pan97, SSK^{+}18, SPA04, TTV00, Gri92, CFR09, CSFK08, Don07, FC05, HK03, HUZ06, Kru98, MJGR09, NKCS03, Pan03, TT05, Wol99, ZS08, ZTP^{+}08, BKT92]. Grid-Enabled [SPA04]. Grid-Enabling [GM04]. Grids [ABG02, AVD01, BF02, CCCV04, DT01, Gam07, RS98b, WBUW14, BFG^{+}08, EM98, HMA99, NDFM07, ZYPGD08]. Group [CDP^{+}07, Vid06]. Groups [GT00b]. Guarantees [YWJ03]. Guest [Ano06f, Ano07f, Ano07g, Ano08h, Ano08f, Ano08g, Ano09f, FAMP12, JKRW08, JKRW09, TA09a, Ada15, Ano06g, DMVT14, DT11, DT13, DT15, GK17, GK14, HGH12, Qiu10, Sta12, WKRJ10, WKR11, WKR13, WKR14]. Guided [TFTY05].

H [LTB01]. H2O [KWDS03]. Hadoop [GS15]. Half [JS18, KRS93]. half-duplex [KRS93]. Halos [Bra00]. Hamilton [BDDP98, SŠZ04]. Hamiltonian [LL21, PND02, RB93, Wu00]. Hamiltonicity [HLH^{+}17]. Hand [San03]. Handling [Cam08]. Hard [Kre97, Mak09, Sar92]. Hardness [BGK^{+}06, Fla97]. Hardware [DTST15, Mar98, MAP^{+}19, SW08, SSK^{+}18, ROJ94]. Harness [MS01a, GKSP99]. Harnessing [IKMH09]. Hash [PFM^{+}09]. Hash-Based [PFM^{+}09]. Haskell [HBL03]. Heap [Wu00]. heaps [CC92, SSSM93]. Helmholtz [Sta95]. Help [Pan03]. Here [AKPSR16]. Hereditary [HHH^{+}99]. Heterogeneous [ADKT12, BLMR03, BRT09, BDR^{+}99, CLR11, FZL^{+}16, GGV12, GKSP99, GDDM17, HGH12, IDS^{+}05, LMZ99, LN05, RBS15, VDL^{+}15, DAB^{+}11]. Heuristic [GS15, AS94]. Heuristics [FV99, SYL95]. Hexagonal [GVCS14]. Hiding [AD02]. Hierarchical [BGK00, CLZC11, GEBR^{+}03, HA10, HGH12, Mic98, Wol95]. Hierarchy [Ada09, FDZ99, MPR18, NA10d, RRS12]. High [AFN11, EW10, EW11, GEL03, GBS^{+}07, GS03, GDTF17, GK14, HM01, HL03, HGH12, Hin03, KB19, LCS^{+}14, LY07, MG11, PAG08, PA99, PADB03, SHBG14, TSS^{+}06, UVJ11, GC95, GV96, Sch97]. High-Level [GS03, GDTF17, HM01, HL03, HGH12, Hin03, SHBG14]. High-Performance [GK14, LCS^{+}14]. High-Productivity [LY07]. Higher [HL00]. Higher-Order [HL00]. Highly [CLR11, RV98, Vid06]. HirondML [CRV08]. HMC [LC14]. HMC-SIM [LC14]. Holes [MCI^{+}01]. Homomorphic [HT99]. homomorphisms [Col95, GDH96]. Honeycomb [BBBL04, CMS99, MLY99, YMZL04a, YMZL04b]. Hop [EFZ98]. Hosted [DJMN13]. Hot [ST97]. Hot-Potato [ST97]. Hough [Pan94]. HPC [CLT13, CLZC11, CLR11, FOCK15]. HPCS [LY07]. HPF [Bra00, Guo01]. HPF-Like [Guo01]. Huffman [KLN04, LP94]. Hut [BGLM05]. Hybrid [GKS08, KHAM04, LC14, SFHW11]. Hybrid-Parallel [SFHW11]. Hyper [CQS16]. Hyper-Buttery [CQS16]. Hypercomputation [GB12, Sta12, Syr13b]. Hypercube [IÖ98, KKS00, kLCL20, Nin20, RV98, WD03, ZXYY20, Ber92, RT92, RSS95, Wu95a, Man91]. Hypercube-Like [kLCL20]. Hypercubes [AH98, BBBL04, FPP98, HY18, KSAOK05, KMY97, RS98b, SV00, Ste12, WW19, YL10, ZXY21, ZZL^{+}21, CF93, DGVG96, GP96, HK95b, KN92, Lat91, LS91, PPRZ93, Tel95, UN92, Zia95]. hypercubic [KP96]. Hypergraphs [BEGK00, KBGE07, LPL20]. Hypermedia [MRS04]. Hypermeshes [KKC98].

I-P [XU99]. I/O

[Agu02, BKT95, Geo01, GKS08, LFC^{+}10]. ICGS [Yan05]. Identification [NTV12]. Idle [IKMH09]. II [Ano04d, Cos03, Cos04]. ILP [Sol09a, TE04]. Image [Chu96, DEKS04, MTD98, VLR^{+}03]. Images [MTD98, SPA04]. Impact [AP05, BL08, DPM^{+}10, SKL10]. imperfectly [Xue96]. Implement [MMRT06]. Implementation [Ada07, AKP99, CF03, Dec21, Gav08, GB98, GBS^{+}07, HKS01, KH02, LSF14, MMAL06, OTK15, Per21, SAJ10, YKLD14, AR95, GV96]. Implementations [CH03, HL02, Nit05, SCJ^{+}10]. Implementing [CG98a, GHSJ94, KW19, MG11, PK08, BA95]. implications [DRS96]. Importance [Akl10, TB18, ZS08]. Importance-Aware [ZS08]. Impossibility [RS98a]. Imprecise [ZZL^{+}21]. Improve [BNW07, GS15, Sol09a]. Improved [Gon09, JS06, JSYD07, LPP01, MV04, QM98, SCJ^{+}10, Yan05, YDL91]. Improvement [CP98c, BK92]. Improving [CRY^{+}03, IDS^{+}05, MGBG07]. Incomplete [GCP99]. Incompleteness [AMN12]. Incorporating [NTG99]. increase [Dua94]. Increasing [ACK99, HW06]. Incremental [BEGK00, JJ96, Wu95b]. incrementally [Sri96]. Independent [BEGK00, CGW05, FS99, KBGE07, LF97, LZG^{+}19]. Independent-set [CGW05]. Index [Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, GZL^{+}20]. indexing [BFCD94]. Induced [Ahn19a, RB13]. Inducing [Jon15]. Inductive [CCQ13]. InfiniBand [KLJ08]. Influence [KKFZ14]. Information [BKK94, GGH^{+}04, NA07a, ZS08, FDFR93].

Infrastructure [BS01, LO09, NDFM07, OL11, TDS99]. Inherently [Akl06a, Akl06b, CT94]. initialization [HS93]. Input [GHF10, BF94]. inputs [BM95]. Insensitive [BE03]. Insertion [CmL21]. Inspired [ZMD15]. Installation [KHW05]. Instruction [Jes06]. Instruction-Level [Jes06]. Instructions [Ano98g]. Instrumentation [MCI^{+}01]. Integer [BGPT00, CD99, EW10, EW11, Ger18, Pag13, BM95, GAL96b]. Integrated [MIJ16]. Integrating [CWR21, RR08]. Integration [HUZ06]. Intel [MIJ16, OTK15, Sol09b, ZHW09]. Intelligence [VMA11, YF18]. Intelligent [OLC^{+}00, OKH^{+}02]. Intensive [IDS^{+}05, ZSLR21, Qia95]. Interconnect [AD02]. Interconnection [CS13, CQSY20a, GP03, JS18, Lav02, Qiu07, RvGG01, RV98, San03, AD92b, FK96, Fio93, GS94b, Wu96]. Interconnects [BK08]. Interface [DBK^{+}09b]. Intermediate [BF02, GGL12]. International [DMVT14]. Internet [AGLS12, CSK00]. Internet-Based [AGLS12]. Interpolants [KW06]. Interval [AS19, MST21, AER93, Fla97, PB96]. intractability [AER94]. Introduction [Ano08h, GSS08]. Intuitive [Tor09]. Inversion [Rag98, DT96]. Invocation [TT05]. Invocations [RR08]. Irregular [Agr99, Bra00, CA99, RBS15, Trä09, WS03, CSBS95]. Island [MGBG07]. Iso [GGS^{+}15]. Iso-Quality [GGS^{+}15]. Isolating [BCS05]. Isomigration [ABNP00]. Issue [Akl11b, Akl14a, Ano06e, DMVT14, DT13, DT15, FAMP12, GK14, HL03, Sta12, WKR14]. Issues [BDR^{+}99, CDP^{+}07, KHAM04, Pan03]. item [SV93]. Iteration [FNP17, FV16]. Iterations [YS05a]. Iterative [CA99, CLR11, KM09].

Jacobi [BOV15, KSHL14, OV06, Sol09a].

Janus [Ger02]. Java

[AGMM00, CSK00, Gam03, GPS03]. Jini [BS01]. Job [Aba04]. jobs [MSP95]. Join [BE03, BF94]. joins [TL91]. Jordan [RV96]. Josephus [LH02]. JPVM [LGCH99].

K-selection [Ber92]. Karp [DV95]. Kautz [DA93]. kernel [Gla93, HK03]. Kernels [KSHL14]. Kerrighed [VLR^{+}03]. Key [NA07c, NA10b, NNA10]. knapsack [AQRW95]. Kolmogorov [Ada07]. Kronecker [MWZ19]. Krylov [GCP99].

L [ABB^{+}03]. L2 [AFN11]. L2-Cache [AFN11]. Label [AAH14, Bur21]. LAM [RšT06]. LAM/MPI [RšT06]. Language [BGJ10, Deb97, DGRD03, GL05, GKS08, Guo01, HL00, HHOM02, LY07, Sch05]. Languages [LY07, ROJ94]. LAPACK [MCR^{+}17]. Large [BCQO99, BK09, CLT13, CC17, EADN06, GJQ09, HSL09, JKRW08, JKRW09, KHW05, LF14, LO09, Man91, NA05, OL11, WKRJ10, WKR11, WKR13, WKR14, WJ16, WGM^{+}10, YKLD14, Yan05, ZS08, BHQ92, CFG94, HP92]. Large-Scale [CLT13, GJQ09, HSL09, JKRW08, KHW05, LO09, OL11, WKRJ10, WKR11, WKR13, WKR14, WGM^{+}10, ZS08]. largest [OSW93]. last [Tro93]. Latency [AD02, AFN11, Jes93, SW08]. Latent [KY06]. Lattice [PKW^{+}03, ZHW09, DM96, OTK15]. lattice-based [DM96]. Lattices [KLW10]. Laws [CWR21]. Layer [GHS97]. Layered [MDAT17]. Layout [EFZ98, KKKL18]. Layouts [DLP03, BKT95]. Lazy [HHL^{+}07]. Leader [DDLV17, MR01, SM20, AJ96]. Leader-Based [MR01]. Leadership [MMRT06]. Leaf [WXW20]. Leaf-Sort [WXW20]. Leaping [HUZ06]. Learning [BCBB09, CmL19a, HlMD21, HL98b, KKFZ14, YF18, ZJSY20]. Learning-Based [ZJSY20]. least [LP94]. Legacy [Gam07]. Legality [BF05]. Lemma [JS18]. Length [HoHqZS20, KLN04, KN06, Kop96]. Length-Limited [KLN04]. Length-Restricted [KN06]. Level [EADN06, GS03, GDTF17, GGL12, HM01, HL03, HGH12, Hin03, Jes06, PAG08, SHBG14, TSS^{+}06, UVJ11, Gla93]. Leveraging [WHTW10]. Libraries [CG01, Kru08, Mer15, TDS99]. Library [BGL03, Bra00, KH02, SHBG14, TSS^{+}06]. Lifting [Nit02]. Lightweight [YKLD14]. Like [Guo01, kLCL20, CWR21, LV96]. Limitations [BW16]. Limited [CE98, KLN04]. Line [CGS15, MBT15, Kri91]. Linear [AA21, BRT09, BW16, DKR91, ER99, GGHJ04, LPL20, RRM15, Sol09b, Tel95, TX98, Yan05, BE96, Gal96a, GMR92, HT94, MMMS94, Myo92, Oks95, Sto96, SJL94, TM93]. Linearizability [CBC03]. Linearizable [FNP17, MNR96]. Linearly [MCCW15]. Link [MS99]. Linked [WJ12, DH94]. Links [GLMM06]. Linux [RšT06]. List [HHL^{+}07, Sib02, WJ12, Alb92, Col95, DH94, KY96]. List-Based [HHL^{+}07]. Listing [VP98]. literal [KCH92]. Little [Ahn19b, VDL^{+}15]. Live [VRR20]. LLC [DGRD03]. Load [CWR21, CLR11, CLZ92, DLP03, EFZ98, EMPF04, GC06, GT00b, KMAG01, MRRV07, ST14, VH00, XL94]. Load-Balanced [ST14]. Load-Balancing [MRRV07, VH00]. Loads [Rob11]. Local [CC17, LMT10, MM16, NTV12, ZXY20]. Local-Edge-Connectivity [ZXY20]. Locality [LTB01, VH00]. Locality-Preserving [VH00]. Locating [Nag09, PPF12]. Location [DFRC01, SMK98]. Lock [APSF01]. Log [Bur21]. Logarithm [CLL16]. Logarithmic [SM20, OSW93]. Logging [Ahn19b]. Logic [FGK15, MMT17, CDZ96, KCH92]. Logical [APSF01]. logP [SV93]. logspace [Sar92]. Long [CmL21]. Long-Term [CmL21]. Longest [ACK99, Per21]. look [BOSW94]. Loop [AL05, AGMM00, DSV97, HLJ97, RDK00, TFTY05, Xue97, AS94, GS94a, HT94, ROJ94]. loop-based [ROJ94]. Loops [LSW97, BFCD94, BL92, CRF95, DV97, PPRZ93, Xue96, van94, GL94]. loosely [CF91]. Losses [KE00]. Low [BAS96b, GGL12, HTHH05, LBD^{+}14, LSF14, PT07, TB18, BKT92, DDT95, DGJS95]. low-cost [DGJS95]. Low-Density [PT07]. Low-Level [GGL12]. Lower [HT21, CL93, Gal96a, Kop96, Mac93, RLG91]. LRU [CmL21]. LSTM [Dec21]. LU [DDT95, GCP99, PBKP92]. Lyapunov [BCQO99, CHQ96]. LZ [De 04].

M [Ram94]. Mach [LBD^{+}14]. Machine [Ada07, CmL19a, GKSP99, HlMD21, OII12, Ram94]. Machines [BK09, BNW07, Pag13, Gri92, NvG95, SMS95, TB94]. Macro [ADKT12, DDT17]. Made [Sax20]. Maekawa [BK92]. Maintaining [AS19, ADG91]. Maintenance [DGB10, FL96]. make [BK92]. Makespan [AB16]. Malicious [AGLS12]. Malleable [VD03]. Management [CFR09, HMA99, Jes08, NDFM07, SMK98, SYR13a, JD92]. Managing [ARA20, DJMN13]. Manipulating [QRR97]. manipulation [Sah95]. manuscripts [Ano98g]. Many [AFN11, Jes08, KBGE07, MCCW15, UVJ11, MIJ16]. Many-Core [AFN11, Jes08, UVJ11]. Map [TH13]. Map-Reduce [TH13]. Mapped [FVR20]. Mapping [AB08, BLY^{+}08, BRT09, BK08, CLZC11, EB98, UN92, GL94, GS95, HL96]. Mappings [BMRGR13, LF97, DDR96, KP94]. MapReduce [THAJ15]. mark [KCH92]. Marrella [AR96]. Massive [MAP^{+}19]. Massively [BCL97, FUV99, LFC^{+}10, SBR^{+}17, CLZ92]. Master [AGLS12, MRRV07, SB12]. Master-Slave [MRRV07, SB12]. Master-Worker [AGLS12]. matched [VLL^{+}91]. Matching [AMC18, BGK^{+}98, BMFU18, CP12, CL13, CLM^{+}16, Hsu93, HT99, MMCW18, MS01b, Raj02, XUZ02, ZXLY20, ZZZY21, ZMZZ20, ZMZZ21, Chu96, Hag95, YDL91]. Materialized [JL05]. Mathematical [Per21, Sza12, Vio94]. Matrices [Jai99, DT96, RV96]. Matrix [ACCLS14, Box09, FV16, Fuj08, GZW16, HT98, LSF14, SFHW11, SAJ10, WJ16, Lin92, MP93]. Matrix-Vector [Fuj08, GZW16, SFHW11]. max [CC92, BCD95]. Maximal [BBF05, BEGK00, CLM^{+}16, FS99, GGHJ04, KBGE07, MB98, NTHK17, ZXY20, YDL91]. Maximum [HC02, Hag92, PD95]. MCMC [ZN13]. MD [HTHH05]. Measured [HRH18]. Mechanism [CFM^{+}99, GC06]. Mechanisms [JSYD07, Jon15, Pel06, VH00]. Median [Nag09]. Mediator [JL05]. Medical [MTD98, SPA04]. MEDINA [SBR^{+}17]. Medium [TW01, VDL^{+}15]. Meertens [CS95]. memories [VLL^{+}91]. Memory [AL05, ABO06, BF97, BCS05, Cho98, CA99, DDLV17, FDZ99, GRS07, GPS03, GHF10, GKS08, GS97, HKS01, HK16, IT02, KE00, LPP01, LC14, LM00, OLC^{+}00, OKH^{+}02, Pag13, RRS12, SKL10, SWK^{+}13, VH00, VH01, VP96, WR97, Yai08, CHQ96, DJM91, DGJS95, FFSY93, GHSJ94, HS93, JD92, MNR96, NvG95, PPRZ93, Sah95]. Memristors [FVR20]. Merge [ST14]. Mesh [AG98, BK08, DDS17, EWM04, HT98, KCPP00, MB98, Ran05, San03, ST97, Tan09, Chl92, CR95, Fra93b, FY93, HS96, HP92, KY96, Kri91, KN95, Mac95, MB95, SC96, XL94, JS06, LJ10]. mesh-bus [FY93, HS96]. mesh-connected [Chl92, Kri91]. Meshe [ZJ09]. Meshes [CQS11, FV99, PS98, YMZL04a, Bar93, BA95, BAS96a, BOSZ92, BOSW94, CM95, CL93, CH94, Chu96, KSS95, KRS93, KW92, LOSZ93, Nak95b, Sue95, WV95]. Message [Ahn19b, AP05, CA99, Her05, PA99, TW01, Yai08, BNKS93, BAS96b, BH93, PBKP92].

Message-Buffer [AP05].

Message-Passing [Her05, BNKS93, BH93]. Messages [KW19]. Messy [HL98a]. Metacomputing [Gam03, GL05, LFJ99, MS01a]. Metadata [KBFB01, ZYPGD08]. Metaheuristics [AIL16]. Method [AH00a, AH00b, CBV^{+}05, CSK00, LG97, OK20, Pur06, TFTY05, Tan09, Yan05, DRS96, Sta95, XL94, OTK15]. Methodologies [GJQ09]. methodology [PRS95]. Methods [FAMP12, GCP99, KB05]. Metric [Fox13]. Metrics [DPM^{+}10, DGT98]. MIC [MIJ16]. microprocessors [DGJS95]. Microscopy [BCMC^{+}04]. Microthreading [Jes06]. Middleware [JSYD07, SS13]. Migratable [VD03]. Migrating [GM03]. Migration [CCR11, RB13]. Migration-Induced [RB13]. Migrations [CRV08]. Miller [DV95]. Millions [BBG^{+}11]. MIMD [DDTV92, MMMS94]. min [CC92, BCD95]. min-max [CC92]. Minimal [DWS15, DVWZ99, HRH18, BH96, Oks95, SC96]. Minimizing [HL99, Hua06]. Minimum [AA21, DF99, DW07a, LP98, BM95, FL96, Zar97]. Mining [TH13, THAJ15]. Minla [Pet03]. Misses [RB13]. Mixed [BLMR03]. ML [Gav08]. Mobile [HT21, PMW^{+}12, SMK98]. Mobility [CLH13, DTL05]. Mock [Sui17]. mode [Alb91]. Model [Agu02, Ald03, BVHR13, BNW07, CGW05, Deb97, DTST15, GZW16, GS99, GKS08, Guo01, GMCC05, Jes06, Jon15, KSAOK05, KSOK07, KSHL14, LL21, MGBG07, MIJ16, MDAT17, Ran05, RvW98, WR97, AD92a, DH94, GAL96b, GDC94, Mac93, SV93, WY92, Sch09]. Model-Based [Deb97, BNW07]. Modeling [BKW98, CP98a, DBK09a, KHW05, KB11, KE00, LDC08, ZSLC18]. Modelling [DJM91, HK16, NA05]. Models [BSM^{+}16, BA01, JW07, KHAM04, TBV98, XLP12, Fra93a, GS94b, Mac95]. Modified [GS15]. Modular [EW13, Gav08, LF97, DDR96]. modulo [BFCD94]. Monitoring [DBK^{+}09b, GGV12, GKK14, RGM06, AD92a]. monoid [CC93]. Monte [HUZ06, MHC^{+}04]. Morphogenesis [Mac12]. morphological [DJM91]. Most [VSS00]. Moves [MG09]. MPI [BBG^{+}11, BIC05, CGS15, DBK^{+}09b, FC05, GB98, GBS^{+}07, LNLE00, Ran05, RšT06]. MPI-FT [LNLE00]. MPICH [RšT06]. MR [TH13]. MR-ARM [TH13]. Multi [AAH14, DDS17, DZ13, GZW16, LSW97, Len92, MGBG07, MKA98, OII12, RvGG01, SKL10, Sol09a, Sol09b, SHBG14, BH93, CL93, DGJS95, DAB^{+}11, MSP95, TL91, ZN13]. multi-chain [ZN13]. Multi-Core [DZ13, SKL10, Sol09a, Sol09b, DAB^{+}11]. Multi-Dimensional [LSW97, CL93]. Multi-Disk [DZ13]. Multi-GPU [GZW16, SHBG14]. multi-joins [TL91]. Multi-Label [AAH14]. Multi-Mesh [DDS17]. multi-microprocessors [DGJS95]. Multi-Objective [MGBG07]. Multi-Organization [OII12]. Multi-Plate [MKA98]. multi-port [BH93]. multi-proposal [ZN13]. Multi-rate [Len92]. Multi-Stage [RvGG01]. multi-task [MSP95]. Multicast [BF02, SV00, Dua93, YTL92]. Multicomputers [WD03, DDTV92, KY96, LH95, YTL92]. Multicore [ACCLS14, GR11, RRS12, SFHW11, WHTW10]. Multicore-Based [SFHW11]. Multicores [Ger18, GDDM17]. Multidimensional [FKB^{+}99, WV95]. multifolded [Zia95]. Multigrid [HCD^{+}19, SSK^{+}18]. Multiple [DPM^{+}10, DHBL06, EW11, EW13, FZL^{+}16, San03, TLH11, Wu98, BOSZ92, BOSW94, SV93]. Multiple-Bus [Wu98]. multiple-item [SV93]. multiplexed [AD92b]. Multiplication [ACCLS14, Box09, EW10, EW11, Fuj08, GZW16, SFHW11, WJ16, DS97, Lin92]. multiplicity [DA93]. Multipoint [BBF05]. multipole [DRS96, Sta95]. Multiprecision [EW13]. multiprocessing [SYL95]. Multiprocessor [AKP99, BGK00, BDG97, CSBS95, EADN06, KMAG01, LZC^{+}21, Mak09, MMMT12, MM16, RB13, TKE^{+}08, Tho03, VKS99, CHQ96, Gri92, MSP95]. Multiprocessors [PP97, Trä09, VH00, VH01, ZJ09, Ber92, DDT95, GHSJ94, PBKP92, VP96]. Multiprogrammed [Vas03]. Multisearch [Jai99]. Multitasking [RR08]. Multithreaded [AKP99, CmL19a, CmL19b, CGK06, KMAG01, RBAA05, WJ12, YWJ03]. Multithreading [KET06, SW08, SLG04]. Mutual [IN99, KM02, KY06, Vid06, BK92, CM96]. mutual-exclusion [CM96]. Myrinet [Geo01, TW01]. Myth [Akl06c, GB12].

n [SF91, Fra93b, TM93]. N-ary [TM93]. N-processor [Fra93b]. Nash [DM12]. Native [WJ16]. Natural [DMVT14]. NC [Hag95, Sar92, SW93, VSS00, dGP92]. NC-algorithms [SW93]. Near [Kre97]. Nearest [Sax20]. Nearly [HL98b, DKR91]. NEC [ZHW09]. Necessity [BA01]. Need [SLG04]. Negotiation [LLD^{+}03]. Nehalem [BDH^{+}08, ZHW09]. Neighbor [LZC^{+}21]. Neighborhood [DW07a, JADT02, LMT10]. Neighbourhood [FPP98, Qiu07]. Nested [HA10, RDK00, BFCD94, DV97, HT94, PPRZ93, Xue96]. Nesterov [OK20]. Net [ESSP01, ZTS^{+}16, GSS08]. Net-Console [ESSP01]. Nets [HPP06]. Network [Ahn19a, BB09, BLY^{+}08, BR19, BDDP98, BK09, BLL12, CLH13, DWH^{+}06, DM12, GPP09, HSL09, JSYD07, KW19, Lav02, LDC08, LZS^{+}18, MT16, MMAL06, MBT15, PMW^{+}12, RJ21, Res97, SBR^{+}17, VAP94, XLP12, Aln94, AD92b, FU92, GMR92, GC95, GS94b, MB95, Par92, Sri96, Wol99]. Network-based [Ahn19a].

Network-On-Chip [BLY^{+}08]. Networks [AS19, AKPSR16, ACM19b, BD02, BBBL04, BGK^{+}06, CKP00, CMS99, CC11, CM97, CDM^{+}03, CS13, CQS16, CQSY20a, CC17, CLM^{+}16, DDT99, DGB10, DR00, DPS00, DVWZ99, EFZ98, Far98, FVR20, FAMP12, GKMP05, GP03, HL99, HY19, HLJ97, JS18, JADT02, Jon15, KSOK07, Kme14, LG97, LLC11, kLCL20, LL21, LZY18, Mal97, MTD98, MCCW15, MV08a, NTV12, PV99, PV00, Qiu07, RvGG01, RV98, RV09, RDK00, Rob11, San03, SB15, SZH14, SŠZ04, Tur10, VRR20, WS03, ZKmST18, ZSLR21, ZMZZ20, ZMZZ21, ZTP^{+}08, ABIM93, BA91, BHQ92, CC93, DA93, Dua92, Dua94, FP93, FK96, FFSY93, Fio93, HS96, KP96, Man91, Pel92, Pel95, Sie96, Sin96, SF91, Wu96, XL94]. Neural [Kme14, MTD98, RJ21, Sie96]. Newton [FV16]. Next [Sch09]. NIL [Sie96]. NOC [BLY^{+}08]. Node [GEL03, MLY99]. Nodes [Dev05, DBK^{+}09b, KMY97, Wu95a]. Noise [HSL09, MAP^{+}19]. Non [Fox13, PT07, SPA04, ZJ09]. Non-Asymptotic [PT07]. Non-Rigid [SPA04]. Non-uniform [ZJ09]. Nondeterminism [Kra01]. nondeterministic [Vir93]. Nonlinear [Akl03]. Nonnegative [FV16]. NOR [SX02]. NOR-Circuits [SX02]. normal [RV96]. Normality [GR99]. Note [Ada15, Akl11a, Akl11b, Akl12d, Akl12a, Akl12b, Akl12c, Akl13a, Akl13b, Akl13c, Akl13d, Akl14a, Akl14d, Akl14b, Akl14c, Akl15a, Akl15b, Akl15c, Akl15d, Akl16a, Akl16b, Akl16c, Akl16d, Akl17a, Akl17b, Ano97a, Ano97b, Ano97c, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano01b, Ano01c, Ano02b, Ano02c, Ano03b, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano06e, Ano06b, Ano06c, Ano06d, Ano06g, Ano06f, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, ACM19b, AKP99, BBL04, BET04, CD99, CBC03, CE98, Cos93, Cos95, Cos96a, Cos96b, Cos97, Cos99, Cos00, Cos01, Cos03, Cos04, DMVT14, DFRC01]. Note [DT11, DT13, DT15, FAMP12, GK17, GK14, HGH12, JKRW08, JKRW09, KBH03, LZG^{+}19, Qiu10, Sta12, TA09a, WW19, WKRJ10, WKR11, WKR13, WKR14, YS05b, ZXLY20, CQS09, Kri91]. Notifications [LLD^{+}03]. Novel [CC93, GZW16, Mar98]. NP [CCQ13, Sch09]. NP-Complete [Sch09]. NTT [MCDB12]. NTT-Based [MCDB12]. NUMA [ACCLS14]. NUMA-Aware [ACCLS14]. Number [KOKM14, LZG^{+}19, DS97, RB93, Wu95a]. numbers [Fra93b]. Numerical [KB11, MV08b].

O [Agu02, BKT95, Geo01, GKS08, LFC^{+}10, SF91]. Object [BCL97, FKB^{+}99, MPR18, Qui00, Wan93]. Object-Oriented [BCL97, Qui00]. Object-Relational [FKB^{+}99]. Objective [MGBG07, CF03, CRV08]. Objects [GR99]. Observation [CFM^{+}99]. Observing [DPM^{+}10]. OCaml [BCL12]. OcamlP3L [DLPW08]. October [BET04]. octrees [SJL94]. Off [CGS15, Geo01, IR16, Kri91, LP94]. Off-Line [CGS15, Kri91]. Off-Processor [Geo01]. Offering [XLP12]. Offs [DLP03]. old [RR91]. Omega [Fio93]. OmpSs [DAB^{+}11]. On-Chip [CmL19a]. One [BIW11, BOV15, FV99, LMS98, NA10c, Pan97, Sib02, Sol09a, DDR96, KN95]. One-Dimensional [Sib02]. One-Phase [LMS98]. One-Sided [Sol09a]. One-Step [BIW11]. One-Time [NA10c]. One-To-All [FV99]. one-to-one [DDR96]. One-Way [Pan97]. Online [BMFU18, GKK14, PLR^{+}95, dSJR^{+}15]. Onto [BRT09, LH95, van94]. Ontology [Sza12]. Opacity [Hea07]. open [IR95, GBS^{+}07]. OpenCL [GDDM17, OTK15]. OpenMP [Car07, HA10, Mar03, Per21, Ran05, RBAA05]. Operating [Jes08, MJGR09, OKH^{+}02, VLR^{+}03]. Operational [HHOM02, RGM06]. Operations [Agu02, BIC05, Cho98, FNP17, Nit05, Trä09, BH93, CH96, SW93, TB94, RGM06]. OPIOM [Geo01]. Optical [AD02, CKP00, CC11, ER99, HQP98, KM02, LTZ98, PSSV02, QM98, RS97, AD92b, BAS96b, Qia95]. Optimal [AA21, BBL04, BBBL04, BH96, De 04, DWH^{+}06, DV02, Dol98, EMPF04, FY93, GW94, HL98b, HT99, IK96, IÖ98, JADT02, KR98, KN06, Kme14, Kre97, KN92, Kru98, MS99, Mic98, NF06, RS98b, Sib02, SŠ03, Tur10, WSK16, XUZ02, XL94, YL10, dGP92, AS92, BNKS93, CC95, DKR91, JR96, LM97, NR95, PB96, Per94]. Optimality [BGK03, PV99, PV00, Sch93]. Optimally [Far98]. optimistic [MNR96]. Optimization [AB99, AB16, BKW98, CmL19b, DHBL06, GZW16, KSHL14, PKW^{+}03, WJ12, XX20, YGM97, ZMD15, KS96]. Optimizations [GGL12, HZW08, KLJ08, KET06]. Optimizing [CSFK08, GKS^{+}14, LSW97, MIJ16, ZSLR21]. optimum [GS94a]. Opus [LMZ99]. Order [AA21, Had08, HL00, Nov20, Pag13, Mac93, VAP94]. order-preserving [Mac93]. ordered [FDFR93, LH95]. Orderings [BOV15]. orders [AER93]. Organization [ABB^{+}03, OII12, RGM06]. Organization-Specific [RGM06]. Organizing [KWDS03]. Oriented [APSF01, BCL97, Qui00]. ORNL [DBK09a]. OROW [NR95]. OROW-PRAM [NR95]. Orthogonal [TTV00]. OTIS [JS06, LJ10]. OTIS-Mesh [JS06, LJ10]. Out-of-Card [WJ16]. out-of-order [VAP94]. Output [MAP^{+}19]. Over-Subscribed [RB13]. Overcome [SW08]. Overhead [Ahn19b, HRH18, MS05, BKT92, DDT95]. Overlap [SFHW11]. Overlay [MV08a, PFM^{+}09]. Overview [ABB^{+}03]. Owned [Aba04].

P [LTB01, BCD95, CCQ13, Gre93, XU99]. P-complete [BCD95, Gre93]. P2P [ZTP^{+}08]. Packet [SBR^{+}17, BHQ92]. Pads [NA10c]. Page [OLC^{+}00]. Page-Based [OLC^{+}00]. Pair [Kru98]. Pairs [VSS00]. pairwise [Par92]. Pancake [Lav02]. Paper [Syr13b]. Papers [Ano04d, BET04, Cos03, Cos04, KBH03, Nak95a]. Paradigm [SS13, LOSZ94]. Parallel [Agu02, AL05, AB99, Akl03, Akl06a, Akl06b, APY06, ARA20, ABNP00, ACK99, BS01, BE03, BGK^{+}98, BVHR13, BOV15, BCQO99, BK09, BGK03, BGL03, BMSW04, BCL97, Box21, Bra00, BDG97, BNW07, BA01, BW16, BKW98, CBV^{+}05, CC92, CG01, CD99, CG98a, CJN99, CS03, CLR11, CC17, Col95, CG97, CFL12, CG98b, DDS17, DZ13, DMS99, DLPW08, DFRC01, DHBL06, DGRD03, EADN06, ESSP01, FUV99, FG02, FV16, FGK15, FR96, FAH03, GLMM06, GR11, Gam03, GZW16, GGV12, GV03, Gav08, GRS07, Ger02, GCP99, GB98, GHS97, GS03, GKS08, Guo01, GS97, GK03, GMCC05, HPP06, HM01, HL03, HGH12, HlMD21, HC02, HPP99, HHOM02, Hin03, HRH18, HHH^{+}99, Hua06, IT02, IDS^{+}05, JW07, Jai99, JS15, JS06, JD92, JKRW08, JKRW09, Jon15, KR98, KB19, KBFB01, KB11]. Parallel [KBGE07, KW06, LSW97, LMZ99, LCK02, LMS98, LPP01, LFC^{+}10, LTB01, LSF14, LGHB07, MCDB12, MGBG07, MKA98, MTD98, MAP^{+}19, MS01a, MS01b, NF06, NA05, NP04, Nit02, OK20, OV06, OII12, OKH^{+}02, Pag13, PND02, PAG08, Per21, PD95, Pet03, PK04, PA99, PKW^{+}03, PPV12, PRH^{+}03, PT98, Rag98, Raj02, RJ21, Res97, RBS15, RRS12, ST12, SFHW11, SX02, ST14, Sui17, SZJK11, SJL94, TTV00, Tan09, THAJ15, Tho03, TMKS16, TX98, VD03, VP98, VMA11, Vas03, Wan93, WKRJ10, WKR11, WKR13, WKR14, WHTW10, Wu95b, XX20, Yan05, ZN13, ZSLC18, ZLL04, ADS91, Alb91, AD92a, AD92b, BM95, BIL92, BH96, CDSS93, CU96, CLM93, CDZ96, CH96, CLZ92, CF95, CF91, DS97, DM96, Dor92, FL96, GI94, GS94a, GMR92, GS94b, GV96, GS95, HK95a]. parallel [Hsu93, IR95, IK96, JR96, KDK^{+}93, KESH95, LV96, LP94, LM97, NRP^{+}94, O’D94, OW92, PB96, PPRZ93, RP93, RV96, SBP94, Sta95, SF91, SMS95, TL91, TM93, Vio94, Wu96, XWF93, Zar97, Zia95, TB94]. Parallel-Based [ARA20]. Parallel/Distributed [NA05]. Parallelism [BLMR03, BGLM05, Cam08, GDH96, GDTF17, Mar03, Mar98, MG11, NA07a, PP01, RBAA05, Vas03, GDC94, KCH92, WY92]. Parallelization [AH00a, AH00b, Ano94, AGMM00, CJ01, DSV97, EB98, HA10, HTHH05, KE06, MHC^{+}04, MMT17, Mas99, MMAL06, TA09b, AS94, BL92]. Parallelizing [Car07, DV97, HW06, NA10a, RvW98]. parameter [Wue09]. parameters [XL94]. Parasites [SZH14]. PARBS [Fra93a]. Parentheses [VP98, XUZ02]. Parenthesis [Tsa04]. Parity [PT07, GM92]. PARM [DF99]. Parssse [SZJK11]. Part [Ano04c, Ano04d, BET04, CP98a, CP98b, Cos03, Cos04, KBH03]. Partial [FV99, Her05, OK20, SMH03]. partially [LH95]. Particles [MG09]. Partition [NTHK17, GI94, Gal96a, JR96, RP93]. partitionable [KN92]. Partitioned [LFC^{+}10]. Partitioning [PFM^{+}09, BF94, CR95, GS95, NM96, XWF93]. Partitions [HHKM13]. Passing [CA99, Her05, Yai08, BNKS93, BH93, PBKP92]. Path [AA21, BSM^{+}16, DLP03, EFZ98, FVR20, MS99, TFTY05, Wu00, YMZL04a, YMZL04b, RR91]. Paths [VSS00, DA93, LM97, TH94]. Patience [NF06]. Pattern [MS01b]. Patterns [CmL21, FG02]. PDEs [Gri92]. Peak [GGV12]. Peer [CDPT10, GEBR^{+}03, PFM^{+}09, TT05]. Peer-To-Peer [CDPT10, GEBR^{+}03, PFM^{+}09, TT05]. Pentagrid [MS09]. Perfect [BGK^{+}98, VRR20, RP93]. Perfectly [ST14]. Performance [ARA20, ABG02, AP05, BDH^{+}08, BMRGR13, DBK09a, DDT95, DBK^{+}09b, FDZ99, GR11, GV03, GEL03, GKK14, GBS^{+}07, GKS^{+}14, GK14, KS97, KB19, KB05, KHW05, KB11, KBG^{+}13, KSAOK05, KSOK07, KE00, KSHL14, LGCH99, LDC08, LCS^{+}14, MSP95, MS05, MGBG07, MMAL06, MG11, Mer15, MMJ^{+}03, OTK15, PA99, PADB03, PKW^{+}03, PBKP92, RvGG01, RšT06, SKL10, Sch97, SS05, Sol09a, Sol09b, SAJ10, Tho03, TW01, WJ16, WGM^{+}10, YTL92, YWJ03, Zav01, ABIM93, GC95, GV96, KS96, KP94]. Performing [GGL12]. Periodic [KW92, TE04]. Permutation [Box09, CJN99, HLJ97, PSSV02, Sue95, CL93, HK95a, Kri91]. Permutations [Pan97, AS92]. Perpetual [LR93]. Petri [HPP06]. Phase [Chi03, LMS98]. Phenomena [BLL12]. Phi [OTK15]. philosophers [Per94]. Photocomputing [Hea07]. Physarum [Ada09, DTST15, Jon15, ZMD15, Ada07]. Physarum}-Inspired [ZMD15]. Physical [HTHH05, Kop96]. Physicalist [Sza12]. Physics [Mac12, MIJ16, Sta12]. pipeline [DT96]. Pipelined [HQP98, GW94, WV95]. Pipelines [GGV12]. Pipelining [BE03, CC93, DGVG96]. Pitfalls [HK16]. Placement [GKMP05, MG09]. Planar [BGK^{+}98, DFRC01, TTV00]. Plane [ACK99]. Plasmodium [Ada09]. Plate [MKA98]. Platform [SBR^{+}17, LO93].

Platforms [BRT09, BDR^{+}99, CLR11, CSFK08, DZ13, FZL^{+}16, MRRV07, SB12]. Player [DM12]. Playground [MJGR09]. Playing [MCI^{+}01]. Point [DFRC01, GBS^{+}07, MCR^{+}17]. Point-to-Point [GBS^{+}07]. Pointer [CP98c]. Poisson [Gla93]. Polaris [PP97]. Policy [Aba04]. Polly [GGL12]. polycephalum} [DTST15, Jon15, Ada09]. Polygon [MS99, BOSZ92]. Polyhedra [QRR97]. Polyhedral [GGL12, WR97]. polynomial [AER93]. polynomials [Hsu93]. Popular [CS13]. Population [SM20]. Port [LF14, BH93]. Portability [Zav01]. Portable [CF03, LNLE00]. Portal [WLW11]. Position [Syr13b]. Positive [TX98]. Possibility [HT21]. postal [Mac93]. Potato [ST97]. Potential [BCMC^{+}04]. Potts [GMCC05]. Power [DDT17, HQP98, IKMH09, Mak09, CF91, OSZ91]. Power-Aware [DDT17]. Power-Share [Mak09]. PowerPack [LCS^{+}14]. powers [Fra93a]. Practical [DMNP11, LO93]. Practice [DMVT14]. Practices [TB18]. PRAM [AJ96, DH94, De 04, Fra93a, GS94b, HS93, HS97, Lin92, NR95]. PRAMs [BKK94]. PRAS [KDK^{+}93]. precedence [UN92]. Precision [EW10, EW11]. Preclusion [AMC18, CP12, CL13, MMCW18, ZXLY20, ZZZY21, ZMZZ20, ZMZZ21]. Preconditioned [OV06]. Preconditioner [GCP99]. Preconditioners [HCD^{+}19]. Predicated [SMH03]. Predictability [BGK03]. Predicting [Wol99]. Prediction [ABG02, DBK09a, Di 00, FDZ99, JSYD07, KS97, MV08b, dSJR^{+}15, KS96]. Predictions [BNW07]. Predictive [KHW05]. Preemption [CE98, RS98a]. Preemptive [Aba04]. Preface [Ano05d, CQSY20b, DT01, DT03, HM01, HL03]. Prefix [CDPT10, DDS17, JS06, Nak95b, WJ12, WS03, SBP94]. Prefix-sums [Nak95b]. PRES [BSM^{+}16]. Presence [AGLS12, DR00, GPP09, RS98a]. Preserving [VH00, Mac93]. Preventive [CCR11]. Price [GLMM06, OII12]. primitives [AR95]. Prior [NA10c]. Prioritization [CmL21]. priority [CH96]. Probabilistic [EADN06]. Problem [BGK00, BBF05, CCQ13, CCCV04, CLL16, CGW05, EFZ98, FS99, HKS01, KU09, LMS98, NA10d, Pet03, SSSM93, Sch09, TFTY05, VSS00, YMZL04a, YMZL04b, AQRW95, BA95, DR94, GI94, Gal96a, JR96, LP94, OSW93, Per94, RP93, RR91, SF91, TH94, Wan96, Zar97]. Problems [ACK99, AB16, CA99, GK03, Kme14, Kre97, MS99, MKA98, PP01, SX02, ZMD15, BM95, BOSZ92, Man91, Myo92]. Process [AB08, DPM^{+}10, IN99, RDK00, Sin96]. Processes [GM03, KKFZ14, Vir93]. Processing [AIL16, ATM01, BCL12, BKW98, EW10, GSS08, JKRW08, JKRW09, LGHB07, Mak09, MKA98, MTD98, NA07a, WKRJ10, WKR11, WKR13, WKR14, AD92a, TL91]. Processor [BDH^{+}08, Geo01, GT00b, HW06, LTZ98, SW08, SS05, SAJ10, XUZ02, CC93, Fra93b, GW94, Kri91, LJ95, OSZ91, PPRZ93, PS93, RR91, SC96, YDL94]. processor-time-minimal [SC96]. Processors [BGPT00, CmL19a, CmL19b, HQP98, SKL10, Sol09b, TE04, WHTW10, DTLA93, GS94a, GMR92, Sto96, RSS95]. Product [AA21, ACM19b, EMPF04, MWZ19, Rob11, BA95, Wan96]. production [BA95]. Productive [GDTF17]. Productivity [LY07]. Products [ACM19a]. Profiling [LCS^{+}14, MS05, PKW^{+}03]. Program [ABG02, BF05, HC02, KS96, Mas99, LH95, Wol95]. Programming [BGK03, BCL12, CG98b, DLPW08, DT00, FG02, Ger02, GS03, HM01, HL03, HGH12, KH02, KB05, Kru08, KC02, LCK02, MS01a, PAG08, SHBG14, Tor09, TX98, ZTS^{+}16, Col95, DAB^{+}11, GAL96b, Myo92, RB95, Vio94]. Programs [ABNP00, BGLM05, CS03, CG97, EB98, ESSP01, GV03, Gav03, GS97, HPP06, Her05, KE00, LTB01, Lou01, CU96, CDZ96, Dor92, HT94, KCH92, LV96]. Project [LY07]. Proof [CM96, CP98b, GP03, HoHqZS20, O’D94]. Proofs [Gav03, LV96]. Propagation [Bur21, CDM^{+}03, HL98b]. Proper [ACM19a]. Properties [CQSY20a]. Proportionality [VDL^{+}15]. proposal [DAB^{+}11, ZN13]. Protein [CFG04]. PROTEUS [CCCV04]. Protocal [CP98a]. Protocol [Ahn19a, Ahn19b, BM97, CP98a, CP98b, Chi03, KY06, TT05, MNR96]. Protocols [BF97, CDM^{+}03, GHJ^{+}08, NA10b, SM20, UKY09]. Prototype [GM04]. Prototyping [TKE^{+}08]. Provably [CJN99]. Provers [HoHqZS20]. Proving [CU96]. Proximity [Ada09]. Proxy [TLH11]. Pruning [ZYPGD08]. published [Nak95a]. Publishers [LLD^{+}03]. Pure [ST97, TT05]. Purpose [SAJ10]. PVM [LHCT96]. pyramids [Fio93]. Python [Hin03].

QAP [AIL16]. QoS [WD03]. QR [YKLD14, ZSLC18]. QRQW [AJ96]. Quad [BDH^{+}08]. Quad-Core [BDH^{+}08]. Quadtree [BKT95, SJL94]. Quality [GGS^{+}15]. Quantifying [BK09]. Quantitative [KE00]. Quantum [Akl10, AsS19, CBV^{+}05, NA07c, NA07a, NA10b, NNA10, NA10d, Qiu10, XX20, NA10a]. quarternary [DS97]. Quasi [HUZ06]. Queries [FKB^{+}99, JL05]. Query [BKW98]. Queue [GGV12]. Queueing [BNW07]. Queues [FNP17, CH96]. Queuing [SB15].

R [Ram94]. Radiative [MIJ16]. Radio [BGK^{+}06, BHQ92]. radios [RS96]. Radix [MG11]. Rainbow [HY18]. Random [DPS00, FVR20, KOKM14, Mal97, MT16, ST97, TB94, DMP94, Pel95, YDL94]. Randomized [FS99, KY06, LSF14, MMMT12, MM16, DH94, JR96]. Rank [LSF14]. Ranking [BA91, GGS^{+}15, Sib02, ZYPGD08, DH94, dGP92]. Rapid [TKE^{+}08]. rate [Len92, VP96]. Ratio [Hua06, Tur10]. Ratios [BBL04, BGPT00]. RCN [GM92]. reachable [CFG94]. Reaching [VDL^{+}15]. Reaction [MBT15]. Reactions [Cos15b]. Read [BKK94]. ready [Ano98g]. Real [AB99, Akl03, BA01, GGS^{+}15, Mak09, Nag09, PPF12, ZJSY20, AD92a, CM96, CF91, SYL95]. Real-Time [AB99, Akl03, BA01, GGS^{+}15, Mak09, ZJSY20, AD92a, CM96, SYL95]. Reality [Syr13b]. realization [Kop96]. Realized [FGK15]. recognition [NRP^{+}94, Wan93]. Recognizing [NP04]. Reconfigurable [AD02, AD92b, EWM04, MAP^{+}19, MB98, PL11, PAKM08, QM98, RS97, RSS95, SSK^{+}18, TBV98, Aln94, BA91, BA95, BAS96a, Chu96, Fra93b, GC95, GS95, KSS95, LOSZ94, Mac95, MB95, Nak95a, Nak95b, OSZ91, PS93, Ram94, TM93]. Reconfiguration [CM95, RT92]. reconstruction [OW92, XWF93]. Recovery [KSAOK05]. rectangle [OSW93]. Rectilinear [MS99, SC96]. recurrence [Len92, Mon94, Xue91]. recurrences [BFCD94, CS95]. Recurrent [Sch09]. recursions [HL96]. Recursive [AH00a, AH00b, FG02, LG97, FK96]. recursively [NR95]. Redistribution [MRRV07]. redistributions [GHSJ94]. Reduce [GS97, TH13]. Reducing [BVHR13, RB13]. Reduction [LSW97, SW08, Jes93, SMS95]. Reductions [HPP06, HL99, Sar92]. Redundancy [SMH03, Dua94, PS93]. redundant [DS97]. Refined [KCH92]. Refinement [KLH^{+}14, Mas99, Ran05, RRS12]. reflected [Sto96, TM93]. Reflection [NTG99]. Register [TE04, AQRW95]. register-based [AQRW95]. Registration [SPA04]. Regular [Ano04d, BMFU18, Cos03, Cos04, DT00, LZY18, MG09, Trä09]. Regulated [GLMM^{+}04]. Regulatory [GLMM^{+}04].

Related [ACK99, CQS16, PL11, BM95, Man91]. Relation [GVCS14]. Relational [BKW98, FKB^{+}99]. relations [BF94]. Relationships [Trä09]. Relativistic [ANS12]. Relativity [AMN12]. Relaxed [DVWZ99, MV08a, RV09]. Relaxed-Ring [MV08a]. Reliability [AB16, LG97, LZC^{+}21, LZS^{+}18, XLP12, ZXYY20, ZZL^{+}21]. Reliable [AGLS12, BS01, DMP94, IR16, Ray21]. Remark [HY18]. Remote [GS97, RR08]. repairing [FP93]. Reparameterisation [RJ21]. Repetitions [DMS99]. Replication [AKP99, IT02, MP93]. Replication-Based [AKP99]. reporting [LM97]. Repository [FKB^{+}99]. Representation [GGL12, QRR97, Wan93]. Representations [Pan03]. Reprogrammability [ZKmST18]. request [VP96]. Requests [BBF05]. Requires [SM20]. Rescheduling [DPM^{+}10]. Research [CQSY20a]. Reservation [LO09, OL11]. Resilience [IR16]. Resolution [HKS01, ZJSY20]. Resource [CFR09, CmL19a, CDP^{+}07, dSJR^{+}15]. Resources [CFR09, Jes08, Vas03]. Respect [VSS00]. Response [GS15]. Restart [SWK^{+}13]. Restricted [DM12, GLMM06, KN06, KM09, MWZ19, Pag13]. Restriction [LZC^{+}21]. restructure [Xue96]. Results [BIC05, HT21, TBV98, ZHW09, RR91]. Retiming [DSV97]. Retrieval [Per21]. Reuse [WR97]. reverse [Wue09]. Reversible [TA09b]. Revisiting [DV95]. rewrite [Ram94]. RFID [TB18]. Right [San03]. Rigid [SPA04]. Ring [KMY97, MLY99, MV08a, Mic98, RSS95, UKY09, DTLA93, KP96, RS96]. Rings [MST21, PSSV02]. Ripple [TA09b]. Ripple-Carry [TA09b]. Risk [CDP^{+}07]. road [AD92a]. Robin [CmL19b]. Robots [DP09, HT21]. Robust [DW07a, DW07b, Fox13]. Rooted [PV99, PV00]. ROSE [Qui00]. rotator [Sri96]. Round [CmL19b]. Routable [Far98]. Router [KCPP00]. routers [Jes93]. routes [GS94b]. Routines [CLR11]. Routing [BCBB09, BCL97, GPP09, GHS97, GK03, HLJ97, KSS95, KLJ08, KSAOK05, KSOK07, LJ10, Pan97, RS97, RV98, SV00, ST97, SŠZ04, Ste12, WS03, WD03, ABIM93, ARS93, CL93, Dua93, Fla97, GP96, Kri91, Man91, Sue95, YTL92]. routings [HK95b]. row [Sue95]. Rule [TH13]. Rules [AAH14, Wue09]. Run [Mar98]. Run-Time [Mar98]. Running [SB12]. Runtime [ABNP00, CH03, DPM^{+}10, JSYD07, PAKM08, YKLD14, ZSLC18, CSBS95].

s [KCH92, GSS08, ZTS^{+}16]. S-NET [ZTS^{+}16, GSS08]. SAC [GS03]. Safe [DWS15, KY06, Pel95]. Sample [DZ12]. Sandpile [Sch09]. SAT [BCD95]. Satisfaction [BBF05, HKS01]. satisfiability [Hag92]. Savings [DHBL06]. Scalability [SB12, XLP12]. Scalable [Ahn19b, APSF01, CGS15, Fox13, MG11, PP97, VH01, Zia95]. Scalasca [WGM^{+}10]. Scale [CCR11, CLT13, GJQ09, HSL09, JKRW08, JKRW09, KHW05, LO09, LBD^{+}14, OL11, WKRJ10, WKR11, WKR13, WKR14, WGM^{+}10, YKLD14, ZS08]. Scaling [MS01b, ZSLC18]. Scan [Mer15, O’D94]. Scatter [DP09]. Scattering [MKA98, CF93]. scattering-gathering [CF93]. schedule [BKT92]. Scheduler [CFR09, VDL^{+}15]. Schedules [BBL04, Gon09, SYR13a, GS94a]. Scheduling [Aba04, Agu02, AL05, BGK00, BLMR03, BGPT00, CFR09, CF93, CJ01, DSV97, DMNP11, DT96, Don07, FOCK15, GDDM17, GS15, IDS^{+}05, KRS15, KLW10, KHAM04, LN05, Mak09, MRRV07, MRS04, MMMT12, MM16, Nit05, OII12, OKL11, SLG04, TFTY05, VKS99, WSK16, YWJ03, ZSLC18, AER93, DKR91, GG96, LH95, LHCT96, MSP95, NPT95, PLR^{+}95, ROJ94, RB95, SYL95, WLR95, Wol95, Wu95b, Wu96]. Scheme [CLL16, CLH13, LNLE00, Qiu07, Sch93]. Schemes [CP98c, DW07a, EMPF04, LDC08, Fla97, Tro93]. Science [DJMN13, LF14]. Scientific [BDH^{+}08, DT01, DT13, DT15, KLJ08, LF14, LFJ99, NKCS03, dSJR^{+}15]. Search [BMSW04, CC17, CGR16, DHBL06, EW13, Jai99, KR98, LMT10, MM16, SZJK11, Gre93]. Search-Based [EW13]. Second [LBD^{+}14, Pag13]. Second-Generation [LBD^{+}14]. Second-Order [Pag13]. Security [MCI^{+}01, TB18]. Segment [CP98c, HC02]. Selecting [KP94]. Selection [AG98, Box21, CmL19b, CH03, ER99, JL05, Ber92, Chl92, KCH92]. Selective [Rag98]. Selectively [PMW^{+}12]. Self [CLL16, Chi03, CLM^{+}16, DDT17, DDT99, DGT98, DDLV17, Dev05, DWS15, Dol98, FP93, GGH^{+}04, GGHJ04, GHJ^{+}08, HHKM13, KY06, KWDS03, NTHK17, PT97, PV99, PV00, SDLM18, Tur10, ZTP^{+}08, SS92, TH94]. Self-Adaptive [DDT17]. Self-Certified [CLL16]. Self-Composition [ZTP^{+}08]. Self-Organizing [KWDS03]. Self-repairing [FP93]. Self-Stabilization [Dol98, PV99, PV00]. Self-Stabilizing [Chi03, CLM^{+}16, DDT99, DGT98, DDLV17, Dev05, DWS15, GGH^{+}04, GGHJ04, GHJ^{+}08, HHKM13, KY06, NTHK17, PT97, Tur10, SDLM18, SS92, TH94]. Selfish [GPP09]. Semantic [ARA20, Deb97]. Semantics [EB98, Guo01, HHOM02]. Semi [ABO06, Fox13, Wu00]. Semi-Byzantine [ABO06]. Semi-Heap [Wu00]. Semi-Metric [Fox13]. semigroup [BOSW94]. Sender [Ahn19b]. Sender-Based [Ahn19b]. Sensitive [GRS07]. Sensor [DGB10, LLC11, PMW^{+}12]. Sensors [HT21, KLW10]. Separation [MS99, Mac95]. Separator [PND02].

September [Ano06e]. Sequence

[IKMH09, LMS98, XU99, CF93]. sequences [KDK^{+}93]. Sequencing [Pet03]. Sequential [Gam03, JS15, LV96, Nit02, CT94, GV93]. Sequential-like [LV96]. series [Hsu93]. series-parallel [Hsu93]. Server [GM04]. Servers [ABO06, GGS^{+}15, JSYD07]. Service [BFG^{+}08, CGK06, GGS^{+}15, RvGG01, Wol99]. Services [FC05, HK03, TT05, ZS08, ZTP^{+}08]. Set [DWS15, FS99, HHL^{+}07, HRH18, CGW05, KCH92, BCD95]. Sets [BEGK00, GGHJ04, GLMM^{+}04, HHKM13, KBGE07, FDFR93]. Shallow [WBUW14]. Shaped [DMNP11]. Shapley [GZL^{+}20]. Share [Mak09]. Shared [AL05, Ald03, AFN11, BCS05, CA99, CDP^{+}07, FNP17, GPS03, GKS08, HKS01, KE00, Pag13, VH00, VH01, WHTW10, CHQ96, DJM91, DGJS95, MNR96, VP96]. Shared-Memory [KE00, VH00, VH01, VP96]. Sharing [ABO06]. shift [AQRW95]. Short [ZXLY20, CQS09]. Shortening [HoHqZS20]. Shortest [DA93, FVR20, TFTY05, VSS00, TH94]. Shredders [QY20]. Shuffle [AMC18, MV04, NRP^{+}94]. Shuffle-Cubes [AMC18]. Shuffle-Exchange [MV04]. Sided [BOV15, Sol09a]. Sides [San03]. signals [Xue91]. Signature [TLH11]. Silent [Dev05]. Silicon [Jes08]. SIM [LC14]. SIMD [DJM91, FUV99, KE06, SW08, SMS95]. Similarity [CFG04]. Simple [Cos15b, DH94, GR11, MS99, MPR18, Sax20, Tan09, Zar97, AS92, GI94, LM97]. Simpler [JS15]. Simulated [Pet03]. Simulating [LOSZ93]. Simulation [AsS19, BKW98, DDTV92, HTHH05, JW07, KB11, LC14, MMT17, NA05, PRH^{+}03, PT98, UVJ11, DRS96, RSS95]. Simulations [CBV^{+}05, CGS15, GMCC05, LBD^{+}14, MHC^{+}04, PA99, VH00, ZHW09].

Simultaneous [Agu02, RB95, SLG04, SMS95]. Single [EB98, GHS97, VLR^{+}03, VSS00, GI94, JR96]. Single-Layer [GHS97]. Singular [Nov20]. Sisal [CG98a]. Site [CFR09]. Sites [CFR09]. Size [BD02, CC11, CM97, DWH^{+}06, XUZ02, BAS96a]. Sizes [EW11]. SkelCL [SHBG14]. Skeletal [DLPW08, DGRD03]. Skeleton [GB98, Nit05, PK08]. Skeletons [Ald03, ABG02, Ser01, BL08, BGK03, CFL12, Dan01, DTL05, HBL03, KC02, Nit02, PAG08, Zav01]. Skew [BE03]. Skew-Insensitive [BE03]. skewed [BF94]. Slave [MRRV07, SB12]. slices [SJL94]. Sliding [KKS00]. Small [DWH^{+}06, Tan09, BM95]. Smallers [Sax20]. Smallest [UKY09]. Smart [CDM^{+}03, SCF01]. Smoother [KSHL14]. smoothing [Gla93]. Snap [CDPT10, JADT02]. Snap-Stabilizing [CDPT10, JADT02]. Social [NTV12]. Soft [Mak09]. Software [ABB^{+}03, Hin03, KSAOK05, Sui17, Ano98g, Sie96]. Software-Based [KSAOK05]. Solution [CA99, DW07b, NA10d, Rag98, Rob11, SSK^{+}18, YMZL04a, YMZL04b, Gri92]. Solutions [DW07a, DW07b, Kre97]. Solve [MKA98]. Solver [LFC^{+}10, ZLL04]. Solvers [BCQO99, IT02, San03, SSK^{+}18, BE96]. Solving [CCCV04, CGW05, CHQ96, Kme14, MMMS94, SF91]. Some [GZL^{+}20, LFJ99, SKN10, Vir93]. Sort [DZ12, Mer15, WXW20]. Sorted [Jai99]. Sorting [BB09, BR19, CG01, CD99, ER99, GS99, Ger18, JS18, LJ10, MG11, NF06, KSS95, KN95, KW92, Man91, MB95, Par92, Sue95]. Space [BW16, SDLM18, SZJK11, GL94, HL96]. space-time [GL94, HL96]. Spaces [Fox13]. Spacetree [WBUW14]. Spanners [Kru98, CK93, LS91]. Spanning [DF99, LZY18, QM98, SKN10, FL96, Qia95, SS92, Zar97]. Spans [MPR18]. SPar [GDTF17]. Sparse [CA99, FDZ99, GZW16, HlMD21, KW06, Rag98, SFHW11, Yan05, Gri92, KESH95, Pel92, Pel95]. Spatial [MG09]. Spatially [DW07a, DW07b]. Special [Ano06e, DMVT14, DT13, DT15, FAMP12, GK14, GZL^{+}20, HL03, Sta12, WKR14, RP93]. Specific [KSHL14, RGM06]. Specification [CP98a, CP98b, Cho98, Deb97]. Specifications [Her05]. Specifunctions [CS01]. Specifying [Xue91]. Spectral [Pet03]. Speculative [GV03, WY92]. Speed [BGPT00, Pur06]. Speed-Up [Pur06]. Speedup [AB99, IN99, KE00, SYR13a]. Speedup-Aware [SYR13a]. SPL [GKS^{+}14]. SPMD [van94]. SPOC [BCL12]. Spreading [BLL12]. SQL [CGK06]. Squad [UKY09]. Square [QM98, IK96]. SRS [VD03]. Stabilization [Dol98, PV99, PV00]. Stabilizing [CDPT10, Chi03, CLM^{+}16, DDT99, DGT98, DDLV17, Dev05, DWS15, GGH^{+}04, GGHJ04, GHJ^{+}08, HHKM13, JADT02, KY06, NTHK17, PT97, Tur10, SDLM18, SS92, TH94]. Stable [BCQO99, ST14]. Stage [RvGG01]. STAPL [TSS^{+}06]. Star [CL13, MMCW18, Res97, RV98, CQS09, GP96]. State [DMNP11, KB19, SZJK11, UKY09]. State-of-the-Art [KB19]. states [AR96]. Static [BGK^{+}06, GS97, Wol95, GAL96b, GDC94, GC95, KRS93, MSP95, SYL95]. Station [GKMP05]. Status [RGM06]. Steady [DMNP11]. Steady-State [DMNP11]. Steiner [ACM19b]. Stencil [GKS^{+}14, GK14, KLH^{+}14, SHBG14, WHTW10]. Step [BB09, BIW11, ADG91]. Steps [Bur21]. Stepwise [Mas99]. Storage [ABO06, DWH^{+}06, PT07, VH01]. Strassen [EW11, GV96]. Strategies [BLMR03, DEKS04, ZZL^{+}21, CLZ92, GG96]. Strategy [BCL97, GGV12, Mak09, CR95, CC93, KY96, LH95]. Stream [BCL12, GSS08, GDTF17]. streamlining [PRS95]. Streams [FZL^{+}16, GM03]. strides [VLL^{+}91]. Strings [Tsa04, VP98]. Strong [HLH^{+}17, ZXLY20, ZZZY21]. Structure [kLCL20, ZKmST18]. Structured [DDT99, MV08a, Ran05, WSK16, Wu96]. Structures [FAMP12, Gav08, LMT10, KP96]. Studies [ZTS^{+}16]. Study [BGLM05, CDM^{+}03, CFG04, DBK09a, Ger18, HTHH05, MG11, TW01, DDT95, ZN13]. Sub [OSW93]. Sub-logarithmic [OSW93]. subarray [PD95]. Subcube [IÖ98, Lat91, YL10, KN92]. subgraph [FU92]. Subgraphs [LZY18]. Sublinear [KR98]. sublogarithmic [Sri96]. submesh [KY96]. Submission [MHC^{+}04]. Subprograms [Sol09b]. Subscribed [RB13]. Subsequence [Per21, LP94, PD95]. Subspace [GCP99]. Substrate [Ada07]. Substructure [kLCL20]. Subtraction [EW10, SCJ^{+}10]. Suffices [EM98]. suffix [BH96]. Suitable [Tan09]. Sum [HC02]. Summation [Mic98, Fra93b]. sums [Nak95b]. Sun [HZW08]. Super [LPL20, Sri96, ZJSY20]. Super-Resolution [ZJSY20]. Supercomputer [CGW05]. Supercomputers [CCR11, Kra01]. Supercube [ARS93]. Supercubes [SH99]. superprimitivity [IK96]. Support [Bra00, Dan01, DGB10, Mar98, NKCS03, Qui00, CSBS95]. Surface [CQS11, CQS09]. Survey [AIL16, CZ20, GJQ09, SWK^{+}13, DV97]. Survivability [KHAM04]. SVD [BOV15, OV06, Sol09a]. SVP [UVJ11]. Sweep [WGM^{+}10]. switch [SF91]. Switched [IÖ98, PS98]. Switches [LTZ98]. switching [BAS96b]. SX [ZHW09]. Symbolic [CJ01, GV03, JSZM08]. Symmetric [CKP00, GP03, LTZ98, CFG94, SS92]. Symmetrical [UKY09]. Symmetry [XLP12]. Synchronization [APSF01, Chi03, GS97, PT97, RS98a, ST14, UKY09, VAP94]. Synchronization-Free [ST14]. Synchronization-Oriented [APSF01]. Synchronizer [JADT02]. Synchronous [BVHR13, CS03, Gav08, LTZ98, Sui17, VH00, CF95]. Syntenic [FAH03]. Synthesis [RR91, GW94]. Synthesizing [Myo92]. Synthetic [PPF12]. System [ABB^{+}03, ER99, EADN06, Gam03, HlMD21, HTHH05, KMAG01, KHW05, LZC^{+}21, MS01a, MJGR09, PAKM08, RB13, VLR^{+}03, ZJSY20, ZSLC18, DS97, MNR96, Oks95, TM93, GMK13]. System-Level [EADN06]. System-On-Chips [PAKM08]. Systematic [MHKT05]. SystemC [SAJ10]. Systems [AL05, Ahn19a, Ahn19b, BIW11, BW16, CDPT10, CCA01, Cho98, Dol98, EADN06, FGK15, GR11, GEBR^{+}03, GKK14, GJQ09, HZW08, HGH12, Jes08, KB19, KB11, LO09, LCS^{+}14, LN05, LFC^{+}10, Mak09, NA05, OL11, OKH^{+}02, PFM^{+}09, RBS15, RRS12, SFHW11, SWK^{+}13, SHBG14, UVJ11, VD03, VLR^{+}03, VMA11, Wu98, Yan05, ZHW09, AER94, BNKS93, BA95, BAS96b, BH93, CRF95, FMR05, MMMS94, Mon94, OSZ91, PS93, RLG91, TL91]. Systolic [ES92, EWM04, HKP^{+}98, MCDB12, Tsa04, VP99, YKLD14, AS92, AQRW95, BE96, CC95, Gla93, Myo92, Oks95, Xue91]. Systolic-Array [MCDB12].

T2 [HZW08]. Table [KLJ08]. Tabu

[BMSW04]. Tagged [Ser01].

Tagged-Token [Ser01]. Target [BGJ10]. Targeting [ADKT12]. Targets [PMW^{+}12]. Task [BLMR03, DMNP11, Don07, FOCK15, GDDM17, GK13, GS95, KRS15, KU09, LHCT96, LN05, LTB01, Mar03, dSJR^{+}15, AER94, CL95, GG96, MSP95, PS93]. Tasks [LMZ99, Mak09, OKL11, SB12]. Taxonomy [Don07]. technique [Gri92, RT92]. Techniques [ARA20, Ano94, CJ01, DSV97, NTG99, OKH^{+}02, SWK^{+}13, CL95, FL96, RB95].

Technological [SKL10]. Technology [KKFZ14]. Template [BGL03, BGLM05, GM04, Raj02, Chu96, HBL03]. Temporal [AS19, CmL21, WHTW10]. Tensor [ACM19b]. Term [CmL21]. Terminating [Kul07]. Terms [LYW20]. Testing [Sui17, IK96]. Text [De 04]. Their [BD02, KSHL14, YWJ03, FK96, KP96, SSSM93]. THEMIS [KBFB01]. Theorem [HoHqZS20]. Theoretical [Agu02]. Theories [AMN12, Sch15]. Theory [CWR21, DMVT14, Gar14, Dua94, Vio94]. these [BW16]. thinning [DJM91]. Thread [ABNP00, CmL19b, SLG04]. Threads [CRV08]. Three [Akl06c, BCMC^{+}04, Car07, FL96, KBG^{+}13, YMZL04a, YMZL04b, MB95]. Three-Dimensional [BCMC^{+}04]. Tight [Wu95a]. Tighter [DR00, TBV98]. Tiled [ZSLC18]. Tiling [DSV97, GVCS14, KS97, KLW10, Xue97]. Time [AB99, Akl03, BAS96a, BA01, CBC03, CL93, Cos15a, DR00, DMS99, DD99, DVWZ99, Dol98, GGHJ04, GGS^{+}15, GS15, HW06, JSYD07, KS97, KR98, KY06, Kme14, LP98, Mak09, Mar98, MMRT06, MB98, Nag09, NA10c, Oks95, RRM15, RV09, SM20, VRR20, XUZ02, ZJSY20, AD92a, AER93, CM96, CHQ96, Gal96a, GL94, HL96, KS96, LP94, LJ95, MB95, Pan94, Pel92, PS93, SC96, SYL95]. Time-Adaptive [KY06]. Time-Free [MMRT06]. Time-Increasing [HW06]. Time-minimal [Oks95]. Time-Relaxed [DVWZ99, RV09]. Time-size [BAS96a]. Time-To-Live [VRR20]. Timelike [ANS12]. Times [HRH18, MRS04, RvGG01]. timestamps [Sin96]. TLP [Sol09a]. Toeplitz [Oks95]. Token [Ser01, KOKM14, DMP94]. Tolerance [CC11, GK13, LNLE00, NTG99, RT92]. Tolerant [LH02, MLY99, MV08a, Res97, RV98, RV09, SV00, VMA11, Wu98, ZXY20, ARS93, BK92, CM95, CH94, GV93, GP96].

Tolerating [MT16]. Tool [MKA98, AR96]. Tools [Pan03]. Toolset [WGM^{+}10]. Top [CS03]. Top-Down [CS03]. Topologies [Had08, LDC08]. Topology [BL08, BK08, MV08a, VRR20]. Tori [BF02, CQS11, DKP98, DV02, YMZL04b]. Toroidal [PS98]. Torus [DR00, LZY18, MLY99, San03, YTL92]. Tournament [Wu00]. Tours [CDSS93]. Toussaint [Ada09]. TPNC [DMVT14]. Trace [CLT13]. Track [PMW^{+}12]. tracking [JJ96]. Trade [DLP03, LP94]. trade-off [LP94]. Trade-Offs [DLP03]. tradeoffs [BAS96a, LJ95]. Trading [IR16, IT02]. traffic [AD92a, SF91]. Training [RJ21]. Trajectory [CBV^{+}05]. Transactional [GHF10]. Transfer [JSYD07, MIJ16]. Transform [FUV99, Pan94, PA95, NA10a]. Transformation [BFCD94, BF05, Xue97]. Transformations [AGMM00, KKKL18]. transforms [BIL92, GHSJ94]. Transitive [GMR92, PL11, Aln94, JJ96]. Translation [BSM^{+}16]. Transmissions [HT98, Pel92]. Transparency [Hea07]. Transport [Jon15]. Transposition [HT98]. traversal [GV93]. Tree [CKP00, CDPT10, CGR16, DGT98, DMNP11, HL99, IR95, JADT02, MHKT05, Nag09, PV99, PV00, Rob11, SKN10, XU99, XUZ02, YKLD14, FL96, LM97, Wu96, Zar97, dGP92]. Tree-Based [YKLD14]. Tree-Shaped [DMNP11]. Trees [BR19, BGPT00, BW16, DF99, DDLV17, JS15, KR98, LH02, LJ10, QM98, RRM15, SDLM18, VP99, Bar93, LP94, PLR^{+}95, SS92, UN92]. Triangle [NTHK17]. Triangular [Rag98, DT96]. Triangulation [LPP01, RRS12]. Trident [SS05]. Tridiagonal [San03, BE96, MMMS94]. Trilinos [LBD^{+}14]. True [SX02]. Truthful [GC06]. Tuning [CFL12, LTB01, MMJ^{+}03, PA99, HlMD21]. Turing [FG02]. Twentieth [Akl11b]. Two [APY06, BGPT00, BMSW04, Chl92, DM12, GGH^{+}04, HHKM13, KSHL14, LLC11, Nov20, ZTS^{+}16, CR95, KW92, OSZ91]. Two-Dimensional [BMSW04, CR95, KW92, OSZ91]. Two-Factor [LLC11]. Two-Player [DM12]. Typed [GSS08]. typesetting [Ano98g].

UltraSPARC [HZW08]. Unbounded [MT16, AQRW95]. Uncertain [HT21]. Uncertainty [Cos15a, GPP09]. Undecidability [AMN12]. Unfriendly [HHKM13]. Unified [Qiu07, FFSY93, RP93]. Uniform [Ano94, BGPT00, Dol98, BF94, Mon94, Xue91, ZJ09]. Uniformly [CZ20]. unifying [BOSW94, CU96, RR91]. unimodular [Xue96]. Unimodularity [BL92]. Unit [EW10]. Units [AIL16, GW94]. Unity [CP98a, CP98b]. Universal [Akl06c, MS09, NA07a, ZLL04]. Universality [FG02]. Unlabeled [DKP98]. Unrelated [OII12]. unshuffle [CC93]. unshuffle-exchange [CC93]. Unstructured [PFM^{+}09, WLR95]. Unsymmetric [Yan05]. Until [MG09]. update [Tro93]. update-last [Tro93]. updates [FL96]. Updating [DF99]. Upgrade [DBK09a]. Usability [Pan03]. Use [CCA01, KHW05]. Used [VMA11]. User [LLC11]. Using [AB08, ABG02, BGK00, BSM^{+}16, BNW07, Bur21, CP98a, CP98b, Cos15b, DDLV17, EW11, Gam07, GAL96b, GT00b, HW06, KW06, KHAM04, KE06, LF14, LG97, MGBG07, MKA98, Mas99, MTD98, MCI^{+}01, MM16, NTG99, PPF12, PK04, QRR97, Rag98, RvW98, SW08, SMH03, SHBG14, VRR20, WJ12, Wol99, Wu00, WS03, YKLD14, YF18, AR95, Ano98g, CS95, DS97, GHSJ94, RB95, TKE^{+}08, VP96]. Uspensky [Ada07]. Utilisation [GGV12]. Utility [SS13].

Validation [BSM^{+}16]. Value [Nov20, ZN13]. Values [VRR20]. Variable [Vas03]. Variables [YS05a]. Various [CS13]. Vector [Fox13, Fuj08, GZW16, SFHW11, Lin92, ZHW09]. Vectorization [Nov20]. vectors [DV97, VLL^{+}91, VAP94]. Vehicle [GK03]. Verification [BF97, CP98a, CP98b, GC06, HPP06, NA10b, TLH11]. Verifying [AS19, GPS03]. versatile [Zia95]. Versatility [Ray21]. Version [HMA99, Sch97]. Versus [CCQ13, CA99, NNA10]. Vertex [Tur10, CFG94]. Via [ADKT12, DBK09a, SS13, FV16, GDH96, MCR^{+}17, RSS95, KH02]. Video [CRY^{+}03]. Video-On-Demand [CRY^{+}03]. Virtual [CFR09, DLP03, GKSP99, KCPP00, RGM06, YKLD14]. Virtualization [FC05]. Visibility [HT21]. vision [AD92a]. vision-aided [AD92a]. Visual [Pan03, XWF93]. Visualization [HK03]. Visualized [Fox13]. Vital [VSS00]. VLSI [Jes93, Sch93]. Volume [Ano98a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, RLG91]. Voronoi [Cos15b, DFRC01]. Voting [Pel06]. vs [CCR11]. Vulnerability [BD02].

Wait [FNP17, PT97]. Wait-Free [FNP17, PT97]. Waiting [Hua06]. Waksman [BD02]. watchman [GS94b]. Water [WBUW14]. Watermark [MCDB12]. wavelength [AD92b]. wavelength-division [AD92b]. Wavelet [FUV99, KE06]. Wavepacket [CBV^{+}05]. Waves [WBUW14]. Way [Pan97, Tis01]. WDM [CC11]. Weak [DDLV17, Pag13]. Weakest [CS01]. Weakly [NP04]. Weather [MV08b, MIJ16, Wol99]. Web [ESSP01]. Web-Based [ESSP01]. weight [LP94]. Weighted [FVR20, RJ21]. Weighted-Averaging [RJ21]. Well [Tsa04, VP98, WSK16]. Well-Formed [Tsa04, VP98]. Well-Structured [WSK16]. Where [KE00]. While [Cam08, GL94]. WHILE-loops [GL94]. Whole [MPR18]. Wide [PADB03, PT07, ZN13]. Wide-Area [PADB03, PT07]. Window [KKS00]. Winograd [DV95]. Wireless [LLC11, TB18]. Without [APY06, BF02, NA07c, NA10c, Yai08, TB94]. Work [De 04, KRS15, KLN04, LP94, Zar97]. Work-Competitive [KRS15]. Work-Efficient [KLN04, Zar97]. Work-Optimal [De 04]. work-time [LP94]. Worker [AGLS12]. Workers [AGLS12]. Workflow [WLW11]. Workflows [DJMN13, WSK16, dSJR^{+}15]. Workload [DEKS04, Hua06, HW06, SYR13a, BF94]. Workshop [BET04, Ano94]. Workstation [Aba04, Dan01]. Workstations [CG01, PRH^{+}03, Tan09]. Wormhole [Di 00, KCPP00, Dua92, Dua94, HK95b, YTL92]. wrap [KRS93]. wrap-arounds [KRS93].

Xeon [OTK15, Sol09b]. XML [Sch05]. XOmega [Fio93]. XSEDE [GMK13]. XT4 [DBK09a]. XtreemOS [MJGR09].

Z [Wue09]. Z-parameter [Wue09]. Zero [KN95]. Zero-one [KN95].

References

Abraham:2021:MLA

[AA21] Jessie Abraham and Micheal Arockiaraj. Minimum linear arrangement of the Cartesian product of optimal order graph and path. Parallel Processing Letters, 31(01):??, March 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500043.

Abdelhamid:2014:MLR

[AAH14] Neda Abdelhamid, Aladdin Ayesh, and Wael Hadi. Multilabel rules algorithm based associative classification. Parallel Processing Letters, 24(1): 1450001, March 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:1999:PRT

[AB99] S. G. Akl and S. D. Bruda. Parallel real-time optimization: Beyond speedup. Parallel Processing Letters, 9(4): 499–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

6264.  URL http://www.
wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Aldinucci:2008:AMA

[AB08] Marco Aldinucci and Anne Benoit. Automatic mapping of assist applications using process algebra. Parallel Processing Letters, 18(1):175–188, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Aupy:2016:AAE

[AB16] Guillaume Aupy and Anne Benoit. Approximation algorithms for energy, reliability, and makespan optimization problems. Parallel Processing Letters, 26(1):1650001, March 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Abawajy:2004:PJS

[Aba04] Jemal H. Abawajy. Preemptive job scheduling policy for distributively-owned workstation clusters. Parallel Processing Letters, 14(2):255–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Almasi:2003:OBS

[ABB^{+}03] George Almási, Ralph Bellofatto, José Brunheroto, Călin Caşcaval, José G. Castaños, Paul Crumley, C. Christopher Erway, Derek Lieber, Xavier Martorell, José E. Moreira, Ramendra Sahoo, Alda Sanomiya, Luis Ceze, and Karin Strauss. An overview of the Bluegene/L system software organization. Parallel Processing Letters, 13(4):561–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Alt:2002:PDC

[ABG02] M. Alt, H. Bischof, and S. Gorlatch. Program development for computational grids using skeletons and performance prediction. Parallel Processing Letters, 12(2):157–??, June 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Arruabarrena:1993:PEA

[ABIM93] A. Arruabarrena, R. Beivide, C. Izu, and J. Miguel. A performance evaluation of adaptive routing in bidimensional cut-through networks. Parallel Processing Letters, 3(4):469– 484, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Antoniu:2000:CDP

[ABNP00] G. Antoniu, L. Bougé, R. Namyst, and C. Pérez. Compiling data-parallel programs to a distributed runtime environment with thread isomigration. Parallel Processing Letters, 10(2/3):201–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1002_03/S0129626400000202.
                                        html.

Attiya:2006:SMS

[ABO06] Hagit Attiya and Amir Bar-Or. Sharing memory with semi-Byzantine clients and faulty storage servers. Parallel Processing Letters, 16(4):419–428, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Alkowaileet:2014:NAM

[ACCLS14] Wail Y. Alkowaileet, David Carrillo-Cisneros, Robert V. Lim, and Isaac D. Scherson. NUMA-aware multicore matrix multiplication. Parallel Processing Letters, 24(4): 1450006, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Atallah:1999:PAL

[ACK99] M. J. Atallah, D. Z. Chen, and K. S. Klenk. Parallel algorithms for longest increasing chains in the plane and related problems. Parallel Processing Letters, 9(4):511–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Arora:2019:PCD

[ACM19a] Ajay Arora, Eddie Cheng, and Colton Magnant. Proper coloring distance in edge-colored Cartesian products of complete graphs and cycles. Parallel Processing Letters, 29(04): ??, December 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500166.

Arunandhi:2019:NSK

[ACM19b] Pranav Arunandhi, Eddie Cheng, and Christopher Melekian. A note on the Steiner k-diameter of tensor product networks. Parallel Processing Letters, 29 (02):??, June 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500087.

Ali:1992:PPM

[AD92a] A. T. Ali and E. L. Dagless.

A parallel processing model for real-time computer vision-aided road traffic monitoring. Parallel Processing Letters, 2(2-3): 257–264, September 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Aly:1992:RPC

[AD92b] K. A. Aly and P. W. Dowd.

Reconfigurable parallel computer architecture based on wavelength-division multiplexed optical interconnection network. Parallel Processing Letters, 2 (2-3):117–127, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Afsahi:2002:ALH

[AD02] A. Afsahi and N. J. Dimopoulos. Analysis of a latency hiding broadcasting algorithm on a reconfigurable optical interconnect. Parallel Processing Letters, 12(1):41–??, March 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Adamatzky:2007:PMI

[Ada07] Andrew Adamatzky. Physarum machine: Implementation of a Kolmogorov–Uspensky machine on a biological substrate. Parallel Processing Letters, 17(4): 455–467, December 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Adamatzky:2009:DPG

[Ada09] Andrew Adamatzky. Developing proximity graphs by Physarum Polycephalum: Does the plasmodium follow the Toussaint hierarchy? Parallel Processing Letters, 19(1):105–127, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Adamatzky:2015:GEN

[Ada15] Andrew Adamatzky. Guest Editors’ note. Parallel Processing Letters, 25(1):1502001, March 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Arora:1991:MDC

[ADG91] A. Arora, S. Dolev, and M. Gouda. Maintaining digital clocks in step. Parallel Processing Letters, 1(1):11–18, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Aldinucci:2012:THA

[ADKT12] M. Aldinucci, M. Danelutto, P. Kilpatrick, and M. Torquati. Targeting heterogeneous architectures via macro data flow. Parallel Processing Letters, 22 (2):1240006, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:1991:CDP

[ADS91] S. G. Akl, T. Duboux, and I. Stojmenovic. Constant delay parallel counters. Parallel Processing Letters, 1(2):143– 148, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Ali:1993:TCS

[AER93] H. H. Ali and H. El-Rewini. The time complexity of scheduling interval orders with communication is polynomial. Parallel Processing Letters, 3(1):53–58, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Ali:1994:ITA

[AER94] Hesham H. Ali and Hesham El-Rewini. On the intractability of task allocation in distributed systems. Parallel Processing Letters, 4(1-2):149–157, June 1994. CODEN PPLTEE. ISSN 0129-

Alves:2011:HLC

[AFN11] Marco A. Z. Alves, Henrique C. Freitas, and Philippe O. A. Navaux. High latency and contention on shared L2-cache for many-core architectures. Parallel Processing Letters, 21(1): 85–106, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Anand:1998:SAX

[AG98] Tarun Anand and Phalguni Gupta. A selection algorithm for X + Y on mesh. Parallel Processing Letters, 8(3):363– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Anta:2012:RIB

[AGLS12] Antonio Fernández Anta, Chryssis Georgiou, Luis López, and Agustín Santos. Reliable Internet-based master-worker computing in the presence of malicious workers. Parallel Processing Letters, 22(1):1250002, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Artigas:2000:ALT

[AGMM00] P. V. Artigas, M. Gupta, S. P. Mikiff, and J. E. Moreira. Automatic loop transformations and parallelization for Java. Parallel Processing Letters, 10(2/3):153–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000160.
html.

Agrawal:1999:DDA

[Agr99] Gagan Agrawal. Data distribution analysis for irregular and adaptive codes. Parallel Processing Letters, 9(1):135–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Aguilar:2002:GTM

[Agu02] J. Aguilar. A graph theoretical model for scheduling simultaneous I/O operations on parallel and distributed environments. Parallel Processing Letters, 12(1):113–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Andreae:1998:HBG

[AH98] Thomas Andreae and Martin Hintz. On hypercubes in de Bruijn graphs. Parallel Processing Letters, 8(2):259–268, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Ahn:2000:AMPa

[AH00a] J. Ahn and T. Han. An analytical method for parallelization of recursive functions. Parallel Processing Letters, 10(1):87–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S012962640000010X.
html.

Ahn:2000:AMPb

[AH00b] J. Ahn and T. Han. An analytical method for parallelization of recursive functions. Parallel Processing Letters, 10(4):359–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000330.
html.

Ahn:2019:ECI

[Ahn19a] Jinho Ahn. Efficient communication induced checkpointing protocol for broadcast networkbased distributed systems. Parallel Processing Letters, 29 (01):??, March 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S012962641950004X.

Ahn:2019:SSB

[Ahn19b] Jinho Ahn. Scalable senderbased message logging protocol with little communication overhead for distributed systems. Parallel Processing Letters, 29(02):??, June 2019. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500051.

Abdelkafi:2016:SMA

[AIL16] Omar Abdelkafi, Lhassane Idoumghar, and Julien Lepagnot. A survey on the metaheuristics applied to QAP for the graphics processing units. Parallel Processing Letters, 26(3): 1650013, September 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Armen:1996:DLE

[AJ96] C. Armen and D. B. Johnson.

Deterministic leader election on the asynchronous QRQW PRAM. Parallel Processing Letters, 6(2):247–250, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Akl:2003:PRT

[Akl03] Selim G. Akl. Parallel real-time computation of nonlinear feedback functions. Parallel Processing Letters, 13(1):65–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Akl:2006:ESP

[Akl06a] Selim G. Akl. Errata: Inherently Parallel Geometric Computations. Parallel Processing Letters, 16(2):185–188, June 2006. CODEN PPLTEE. ISSN 0129-6264. See [Akl06b].

Akl:2006:IPG

[Akl06b] Selim G. Akl. Inherently parallel geometric computations. Parallel Processing Letters, 16(1): 19–37, March 2006. CODEN PPLTEE. ISSN 0129-6264. See errata [Akl06a].

Akl:2006:TCD

[Akl06c] Selim G. Akl. Three counterexamples to dispel the myth of the universal computer. Parallel Processing Letters, 16(3): 381–403, September 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Akl:2010:IBQ

[Akl10] Selim G. Akl. On the importance of being quantum. Parallel Processing Letters, 20(3): 275–286, September 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Akl:2011:EN

[Akl11a] Selim G. Akl. Editorial note. Parallel Processing Letters, 21 (4):377, December 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Akl:2011:TAI

[Akl11b] Selim G. Akl. Twentieth anniversary issue: Editorial note. Parallel Processing Letters, 21 (3):273, September 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Akl:2012:ENa

[Akl12a] Selim G. Akl. Editorial note. Parallel Processing Letters, 22 (1):1201001, March 2012. CO-DEN PPLTEE. ISSN 0129-6264.

Akl:2012:ENb

[Akl12b] Selim G. Akl. Editorial note. Parallel Processing Letters, 22 (2):1201002, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2012:ENc

[Akl12c] Selim G. Akl. Editorial note. Parallel Processing Letters, 22 (3):1201003, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2012:EN

[Akl12d] Selim G. Akl. Editorial note. Parallel Processing Letters, 22 (4):1201004, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2013:ENa

[Akl13a] Selim G. Akl. Editorial note.

Parallel Processing Letters, 23 (1):1301001, March 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2013:ENb

[Akl13b] Selim G. Akl. Editorial note.

Parallel Processing Letters, 23 (2):1301002, June 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2013:ENc

[Akl13c] Selim G. Akl. Editorial note.

Parallel Processing Letters, 23 (3):1301003, September 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2013:ENd

[Akl13d] Selim G. Akl. Editorial note.

Parallel Processing Letters, 23 (4):1301004, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2014:DIE

[Akl14a] Selim G. Akl. December issue 2014 editorial note. Parallel Processing Letters, 24(4): 1401004, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2014:ENa

[Akl14b] Selim G. Akl. Editorial note. Parallel Processing Letters, 24 (1):1401001, March 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2014:ENb

[Akl14c] Selim G. Akl. Editorial note. Parallel Processing Letters, 24 (2):1401002, June 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2014:EN

[Akl14d] Selim G. Akl. Editorial note. Parallel Processing Letters, 24 (3):1401003, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2015:EN

[Akl15a] Selim G. Akl. Editorial note. Parallel Processing Letters, 25 (1):1501001, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2015:ENa

[Akl15b] Selim G. Akl. Editorial note. Parallel Processing Letters, 25 (2):1501002, June 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2015:ENb

[Akl15c] Selim G. Akl. Editorial note. Parallel Processing Letters, 25 (3):1501003, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2015:ENc

[Akl15d] Selim G. Akl. Editorial note.

Parallel Processing Letters, 25 (4):1501004, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2016:ENa

[Akl16a] Selim G. Akl. Editorial note.

Parallel Processing Letters, 26 (1):1601001, March 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2016:ENb

[Akl16b] Selim G. Akl. Editorial note.

Parallel Processing Letters, 26 (2):1601002, June 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2016:ENc

[Akl16c] Selim G. Akl. Editorial note.

Parallel Processing Letters, 26 (3):1601003, September 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2016:ENd

[Akl16d] Selim G. Akl. Editorial note.

Parallel Processing Letters, 26 (4):1601004, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2017:ENa

[Akl17a] Selim G. Akl. Editorial note. Parallel Processing Letters, 27 (1):1701001, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Akl:2017:ENb

[Akl17b] Selim G. Akl. Editorial note. Parallel Processing Letters, 27 (2):1701002, June 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Azagury:1999:NIR

[AKP99] Alain Azagury, Elliot K. Kolodner, and Erez Petrank. A note on the implementation of replication-based garbage collection for multithreaded applications and multiprocessor environments. Parallel Processing Letters, 9(3):391–??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Applebaum:2016:CHD

[AKPSR16] Benny Applebaum, Dariusz R. Kowalski, Boaz Patt-Shamir, and Adi Rosén. Clique here: On the distributed complexity in fully-connected networks. Parallel Processing Letters, 26(1): 1650004, March 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Aguilar:2005:PLS

[AL05] Jose Aguilar and Ernst Leiss. Parallel loop scheduling approaches for distributed and shared memory systems. Parallel Processing Letters, 15(1/2):

131–152, March/June 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Albacea:1991:DMP

[Alb91] E. A. Albacea. Determining the mode in parallel. Parallel Processing Letters, 1(1):3– 10, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Albacea:1992:CLE

[Alb92] E. A. Albacea. Computation list evaluation and its applications. Parallel Processing Letters, 2(4): 321–329, December 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Aldinucci:2003:ESS

[Ald03] Marco Aldinucci. eskimo: Experimenting with skeletons in the shared address model. Parallel Processing Letters, 13(3): 449–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Alnuweiri:1994:FRN

[Aln94] Hussein M. Alnuweiri. A fast reconfigurable network for graph connectivity and transitive closure. Parallel Processing Letters, 4(1-2):105–115, June 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Anantapantula:2018:MPS

[AMC18] Sai Anantapantula, Christopher Melekian, and Eddie Cheng. Matching preclusion for the shuffle-cubes. Parallel Processing Letters, 28(03): ??, September 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500123.

Andreka:2012:DUG

[AMN12] Hajnal Andréka, Judit Madarász, and István Németi. Decidability, undecidability, and Gödel’s incompleteness in relativity theories. Parallel Processing Letters, 22(3):1240011, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:1994:PTU

[Ano94]  Anonymous.  Parallelization

techniques for uniform algorithms Workshop. Parallel Processing Letters, 4(3):??, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1997:ENa

[Ano97a] Anonymous. Editorial note. Parallel Processing Letters, 7(1): 1–??, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1997:ENb

[Ano97b] Anonymous. Editorial note. Parallel Processing Letters, 7(2): 113–??, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1997:ENd

[Ano97c] Anonymous. Editorial note. Parallel Processing Letters, 7(4): 357–??, December 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:AIV

[Ano98a] Anonymous. Author index volume 8 (1998). Parallel Processing Letters, 8(2):599–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ENa

[Ano98b] Anonymous. Editorial note.

Parallel Processing Letters, 8(1): 1–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ENb

[Ano98c] Anonymous. Editorial note.

Parallel Processing Letters, 8(2): 135–??, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ENc

[Ano98d] Anonymous. Editorial note.

Parallel Processing Letters, 8(3): 269–??, September 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ENd

[Ano98e] Anonymous. Editorial note.

Parallel Processing Letters, 8(3): 270–??, September 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ENe

[Ano98f] Anonymous. Editorial note.

Parallel Processing Letters, 8(4): 419–??, December 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:1998:ITC

[Ano98g]  Anonymous.  Instructions

for typesetting camera-ready manuscripts using computer software. Parallel Processing Letters, 8(2):269–273, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1999:AI

[Ano99a] Anonymous. Author index. Parallel Processing Letters, 9(4):

587–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

6264.  URL http://www.
wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Anonymous:1999:ENa

[Ano99b] Anonymous. Editorial note. Parallel Processing Letters, 9(1): 1–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1999:ENb

[Ano99c] Anonymous. Editorial note. Parallel Processing Letters, 9(2): 171–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:1999:ENc

[Ano99d] Anonymous. Editorial note. Parallel Processing Letters, 9(3): 311–??, September 1999. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2000:AI

[Ano00a] Anonymous. Author index. Parallel Processing Letters, 10(4): 395–??, December 2000. CO-DEN PPLTEE. ISSN 0129-6264.

URL http://ejournals.wspc.
com.sg/ppl/10/1004/.

Anonymous:2000:ENa

[Ano00b] Anonymous. Editorial note. Parallel Processing Letters, 10 (2/3):151–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/.

Anonymous:2000:ENb

[Ano00c] Anonymous. Editorial note.

Parallel Processing Letters, 10 (4):251–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S012962640000024X.
html.

Anonymous:2001:AI

[Ano01a] Anonymous. Author index. Parallel Processing Letters, 11(4): 503–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2001:EN

[Ano01b] Anonymous. Editorial note.

Parallel Processing Letters, 11 (1):3–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2001:NEC

[Ano01c] Anonymous. Note from the Editor-in-Chief. Parallel Processing Letters, 11(1):1–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2002:AIV

[Ano02a] Anonymous. Author index volume 12 (2002). Parallel Processing Letters, 12(3–4):385–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2002:ENa

[Ano02b] Anonymous. Editorial note.

Parallel Processing Letters, 12 (1):1–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2002:ENb

[Ano02c] Anonymous. Editorial note. Parallel Processing Letters, 12(3–4):285–??, September– December 2002. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2003:AIV

[Ano03a]  Anonymous.  Author index

volume 13 (2003). Parallel Processing Letters, 13(4):735– ??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2003:EN

[Ano03b] Anonymous. Editorial note. Parallel Processing Letters, 13 (1):1–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2004:AIV

[Ano04a] Anonymous. Author index volume 14 (2004). Parallel Processing Letters, 14(3/4):423–??, September/December 2004. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2004:EN

[Ano04b] Anonymous. Editorial note. Parallel Processing Letters, 14(3/4):325–??, September/ December 2004. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2004:ENPa

[Ano04c] Anonymous. Editorial note: “Part I: Editorial from Alberto Apostolico”. Parallel Processing Letters, 14(2):131–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2004:ENPb

[Ano04d] Anonymous. Editorial note:

“Part II: Regular Papers”. Parallel Processing Letters, 14(2): 134–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2005:AIV

[Ano05a]  Anonymous.  Author index

volume 15 (2005). Parallel Processing Letters, 15(4):499– 502, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2005:ENa

[Ano05b] Anonymous. Editorial note.

Parallel Processing Letters, 15 (3):237, September 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2005:ENb

[Ano05c] Anonymous. Editorial note.

Parallel Processing Letters, 15 (4):353–356, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2005:P

[Ano05d] Anonymous. Preface. Parallel Processing Letters, 15(1/2): 1–2, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:AIV

[Ano06a] Anonymous. Author index volume 16. Parallel Processing Letters, 16(4):513–515, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:ENa

[Ano06b] Anonymous. Editorial note.

Parallel Processing Letters, 16 (1):1–2, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:ENb

[Ano06c] Anonymous. Editorial note. Parallel Processing Letters, 16 (2):149, June 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:ENc

[Ano06d] Anonymous. Editorial note. Parallel Processing Letters, 16 (4):405–406, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:ENS

[Ano06e] Anonymous. Editorial note: September special issue 2006. Parallel Processing Letters, 16 (3):281–282, September 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:GENb

[Ano06f] Anonymous. Guest editorial note. Parallel Processing Letters, 16(3):283–284, September 2006. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2006:GENa

[Ano06g] Anonymous. Guest Editor’s note. Parallel Processing Letters, 16(2):151, June 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index volume 17. Parallel Processing Letters, 17(4):469–471, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:ENa

[Ano07b] Anonymous. Editorial note.

Parallel Processing Letters, 17 (1):1, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:ENb

[Ano07c] Anonymous. Editorial note.

Parallel Processing Letters, 17 (2):125, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:ENc

[Ano07d] Anonymous. Editorial note.

Parallel Processing Letters, 17 (3):229, September 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:ENd

[Ano07e] Anonymous. Editorial note.

Parallel Processing Letters, 17 (4):337–338, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:GENa

[Ano07f] Anonymous. Guest editorial note. Parallel Processing Letters, 17(1):3, March 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2007:GENb

[Ano07g] Anonymous. Guest editorial note. Parallel Processing Letters, 17(3):231–232, September 2007. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:AIV

[Ano08a] Anonymous. Author index – volume 18. Parallel Processing Letters, 18(4):609–612, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:ENa

[Ano08b] Anonymous. Editorial note. Parallel Processing Letters, 18 (1):1, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:ENb

[Ano08c] Anonymous. Editorial note. Parallel Processing Letters, 18 (2):201, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:ENc

[Ano08d] Anonymous. Editorial note. Parallel Processing Letters, 18 (3):323, September 2008. CO-DEN PPLTEE. ISSN 0129-6264. Anonymous:2008:ENd

[Ano08e] Anonymous. Editorial note. Parallel Processing Letters, 18 (4):447, December 2008. CO-DEN PPLTEE. ISSN 0129-6264. Anonymous:2008:GENa

[Ano08f] Anonymous. Guest editorial note. Parallel Processing Letters, 18(1):3–5, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:GENb

[Ano08g] Anonymous. Guest editorial note. Parallel Processing Letters, 18(3):325–327, September 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2008:GEI

[Ano08h] Anonymous. Guest Editor’s introduction. Parallel Processing Letters, 18(2):203–204, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:AIV

[Ano09a] Anonymous. Author index volume 19. Parallel Processing Letters, 19(4):641–644, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:ENa

[Ano09b] Anonymous. Editorial note.

Parallel Processing Letters, 19 (1):1–2, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:ENb

[Ano09c] Anonymous. Editorial note.

Parallel Processing Letters, 19 (2):185, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:ENc

[Ano09d] Anonymous. Editorial note.

Parallel Processing Letters, 19 (3):377, September 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:ENd

[Ano09e] Anonymous. Editorial note.

Parallel Processing Letters, 19 (4):485, December 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2009:GEN

[Ano09f] Anonymous. Guest Editor’s note. Parallel Processing Letters, 19(3):379, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2010:AIV

[Ano10a] Anonymous. Author index volume 20. Parallel Processing Letters, 20(4):415–417, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2010:ENa

[Ano10b] Anonymous. Editorial note. Parallel Processing Letters, 20 (1):1, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2010:ENb

[Ano10c] Anonymous. Editorial note. Parallel Processing Letters, 20 (2):101, June 2010. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2010:ENc

[Ano10d] Anonymous. Editorial note. Parallel Processing Letters, 20 (3):209, September 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2010:ENd

[Ano10e] Anonymous. Editorial note. Parallel Processing Letters, 20 (4):287, December 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Anonymous:2011:AIV

[Ano11a] Anonymous. Author index volume 21. Parallel Processing Letters, 21(4):479–481, December 2011. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2011:ENa

[Ano11b] Anonymous. Editorial note. Parallel Processing Letters, 21 (1):1, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2011:ENb

[Ano11c] Anonymous. Editorial note. Parallel Processing Letters, 21 (2):107, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 22. Parallel Processing Letters, 22(4):1299001, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2013:AIV

[Ano13] Anonymous. Author index volume 23. Parallel Processing Letters, 23(4):1399001, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2014:AIV

[Ano14] Anonymous. Author index volume 24. Parallel Processing Letters, 24(4):1499001, December 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2015:AIV

[Ano15] Anonymous. Author index volume 25. Parallel Processing Letters, 25(4):1599001, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2016:AIV

[Ano16] Anonymous. Author index: Volume 26. Parallel Processing Letters, 26(4):1699001, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2017:AIV

[Ano17] Anonymous. Author index volume 27. Parallel Processing Letters, 27(3–4):1799001, 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2018:AIV

[Ano18] Anonymous. Author index volume 28. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418990013.

Anonymous:2019:AIV

[Ano19] Anonymous. Author index volume 29. Parallel Processing Letters, 29(04):??, December 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419990019.

Anonymous:2020:AIV

[Ano20] Anonymous. Author index volume 30. Parallel Processing Letters, 30(04):??, December 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420990017.

Anonymous:2021:AIV

[Ano21] Anonymous. Author index volume 31. Parallel Processing Letters, 31(04):??, December 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421990012.

Andreka:2012:CTC

[ANS12] Hajnal Andréka, István Németi, and Gergely Székely. Closed timelike curves in relativistic computation. Parallel Processing Letters, 22(3):1240010, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Arber:2005:IMB

[AP05] Leon Arber and Scott Pakin.

The impact of message-buffer alignment on communication performance. Parallel Processing Letters, 15(1/2):49–65, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Arantes:2001:BLC

[APSF01] L. Arantes, D. Poitrenaud, P. Sens, and B. Folliot. The barrier-lock clock: A scalable synchronization-oriented logical clock. Parallel Processing Letters, 11(1):65–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

AlNaMneh:2006:TPD

[APY06] Rami Al Na’Mneh, W. David Pan, and Seong-Moo Yoo. Two parallel 1-D FFT algorithms without all-to-all communication. Parallel Processing Letters, 16(2):153–164, June 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Andonov:1995:SRS

[AQRW95] R. Andonov, P. Quinton, S. Rajopadhye, and D. Wilde. A shift register-based systolic array for the unbounded knapsack problem. Parallel Processing Letters, 5(2):251–262, June 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Ahuja:1995:IGF

[AR95] Mohan Ahuja and Michel Raynal. An implementation of global flush primitives using counters. Parallel Processing Letters, 5(2):171–178, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Ambroise:1996:MTA

[AR96] D. Ambroise and B. Rozoy. Marrella: a tool to analyse the graph of states. Parallel Processing Letters, 6(4):583–594, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Algosaibi:2020:PBT

[ARA20] Abdulelah Algosaibi, Khaled Ragab, and Saleh Albahli. Parallel-based techniques for managing and analyzing the performance on semantic graph. Parallel Processing Letters, 30 (02):??, June 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500073.

Auletta:1993:FTR

[ARS93] V. Auletta, A. A. Rescigno, and V. Scarano. Fault tolerant routing in the Supercube. Parallel Processing Letters, 3(4):393– 405, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Akl:1992:SOS

[AS92] S. G. Akl and I. Stojmenovic.

A simple optimal systolic algorithm for generating permutations. Parallel Processing Letters, 2(2-3):231–239, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Anderson:1994:FHL

[AS94] Richard J. Anderson and Barbara B. Simons. A fast heuristic for loop parallelization. Parallel Processing Letters, 4(3):281– 299, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Ashir:1997:ECA

[AS97] Yaagoub A. Ashir and Iain A.

Stewart. On embedding cycles in k-ary n-cubes. Parallel Processing Letters, 7(1):49–55, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Akrida:2019:VMC

[AS19] Eleni C. Akrida and Paul G. Spirakis. On verifying and maintaining connectivity of interval temporal networks. Parallel Processing Letters, 29 (02):??, June 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500099.

Alonso-sanz:2019:QFG

[AsS19] Ramón Alonso-sanz and Haozhen Situ. Quantum fuzzy game simulation. Parallel Processing Letters, 29(03):??, September 2019. CODEN PPLTEE.

ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500105.

Amamiya:2001:AFC

[ATM01] M. Amamiya, H. Taniguchi, and T. Matsuzaki. An architecture of fusing communication and execution for global distributed processing. Parallel Processing Letters, 11(1):7–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Arnold:2001:CCD

[AVD01] D. C. Arnold, S. S. Vahdiyar, and J. J. Dongarra. On the convergence of computational and data grids. Parallel Processing Letters, 11(2–3):187–202, 2001. CODEN PPLTEE. ISSN 0129-

6264. URL http://ejournals.
wspc.com.sg/ppl/11/sample/
S012962640100052X.html;
http://www.netlib.org/utk/
people/JackDongarra/PAPERS/
convergence-data-grids.pdf.

Ben-Asher:1991:RRN

[BA91] Y. Ben-Asher. Ranking on reconfigurable networks. Parallel Processing Letters, 1(2):149– 156, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Ben-Asher:1995:CPP

[BA95] Y. Ben-Asher. The Cartesian product problem and implementing production systems on reconfigurable meshes. Parallel Processing Letters, 5(1):49–61, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Bruda:2001:NFM

[BA01] S. D. Bruda and S. G. Akl. On the necessity of formal models for real-time parallel computations. Parallel Processing Letters, 11(2–3):353–??, 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Barth:1993:EMD

[Bar93] D. Barth. Embedding meshes of d-ary trees into de Bruijn graphs. Parallel Processing Letters, 3(2):115–127, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Ben-Asher:1996:TTR

[BAS96a] Y. Ben-Asher and A. Schuster.

Time-size tradeoffs for reconfigurable meshes. Parallel Processing Letters, 6(2):231–245, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Ben-Asher:1996:LCA

[BAS96b] Yosi Ben-Asher and Assaf Schuster. Low crosstalk address encodings for optical message switching systems. Parallel Processing Letters, 6(1):87–100, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Baddar:2009:SSN

[BB09] Sherenaz W. Al-Haj Baddar and Kenneth E. Batcher. An 11step sorting network for 18 elements. Parallel Processing Letters, 19(1):97–103, March 2009. CODEN PPLTEE. ISSN 0129-

Bein:2004:OEH

[BBBL04] Doina Bein, Wolfgang W. Bein, Natasa Brajkovska, and Shahram Latifi. Optimal embedding of honeycomb networks into hypercubes. Parallel Processing Letters, 14(3/4):367–??, September/December 2004. CO-DEN PPLTEE. ISSN 0129-6264.

Barth:2005:CMR

[BBF05] Dominique Barth, Pascal Berthome, and Paraskevi Fragopoulou. The complexity of the maximal requests satisfaction problem in multipoint communication. Parallel Processing Letters, 15(1/2): 209–222, March/June 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Balaji:2011:MMC

[BBG^{+}11] Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, Torsten Hoefler, Sameer Kumar, Ewing Lusk, Rajeev Thakur, and Jesper Larsson Träff. MPI on millions of cores. Parallel Processing Letters, 21(1):45–60, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Baille:2004:NBS

[BBL04] Fabien Baille, Evripidis Bampis, and Christian Laforest. A note on bicriteria schedules with optimal approximations ratios. Parallel Processing Letters, 14(2): 315–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Barth:2009:DLE

[BCBB09] Dominique Barth, Johanne Cohen, Olivier Bournez, and Octave Boussaton. Distributed learning of equilibria in a routing game. Parallel Processing Letters, 19(2):189–204, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Bongiovanni:1995:MSM

[BCD95] Giancarlo Bongiovanni, Pierluigi Crescenzi, and Sergio De Agostino. MAX SAT and MIN SET cover approximation algorithms are P-complete. Parallel Processing Letters, 5(2): 293–298, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Boari:1997:RSO

[BCL97] Maurelio Boari, Antonioo Corradi, and Letizia Leonardi. A routing strategy for objectoriented applications in massively parallel architectures. Parallel Processing Letters, 7(3): 237–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Bourgoin:2012:SGP

[BCL12] Mathias Bourgoin, Emmanuel Chailloux, and Jean-Luc Lamotte. SPOC: GPGPU programming through stream processing with OCaml. Parallel Processing Letters, 22(2):1240007, June 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Bilbao-Castro:2004:PGC

[BCMC^{+}04] J. R. Bilbao-Castro, R. Marabini, J. M. Carazo, I. Garcia, and J. J. Fernandez. The potential of grid computing in three-dimensional electron microscopy. Parallel Processing Letters, 14(2):151–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Benner:1999:PDS

[BCQO99] Peter Benner, José M. Claver, and Enrique S. Quintana-Orti. Parallel distributed solvers for large stable generalized Lyapunov equations. Parallel Processing Letters, 9(1):147–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Byna:2005:ICS

[BCS05] Surendra Byna, Kirk W. Cameron, and Xian-He Sun. Isolating costs in shared memory communication buffering. Parallel Processing Letters, 15(4): 357–365, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Beauquier:2002:ASW

[BD02] B. Beauquier and E. Darrot. On arbitrary size Waksman networks and their vulnerability. Parallel Processing Letters, 12(3–4):287–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Bermond:1998:HCD

[BDDP98] J.-C. Bermond, E. Darrot, O. Delmas, and S. Perennes. Hamilton cycle decomposition of the butterfly network. Parallel Processing Letters, 8(3):371– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Brecht:1997:CDM

[BDG97] Timothy Brecht, Xiaotie Deng, and Nian Gu. Competitive dynamic multiprocessor allocation for parallel applications. Parallel Processing Letters, 7(1):89–100, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Barker:2008:PEN

[BDH^{+}08] Kevin J. Barker, Kei Davis, Adolfy Hoisie, Darren J. Kerbyson, Mike Lang, Scott Pakin, and Jose Carlos Sancho. A performance evaluation of the Nehalem quad-core processor for scientific computing. Parallel Processing Letters, 18(4):453– 469, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Boulet:1999:AIH

[BDR^{+}99] Pierre Boulet, Jack Dongarra, Fabrice Rastello, Yves Robert, and Frédéric Vivien. Algorithmic issues on heterogeneous computing platforms. Parallel Processing Letters, 9(2):197–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Bekakos:1996:CST

[BE96] M. P. Bekakos and O. B.

Efremides. A class of systolic tridiagonal linear solvers. Parallel Processing Letters, 6(3):355– 364, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Bamha:2003:PSI

[BE03] M. Bamha and M. Exbrayat.

Pipelining a skew-insensitive parallel join algorithm. Parallel Processing Letters, 13(3): 317–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Boros:2000:EIA

[BEGK00] E. Boros, K. Elbassioni, V. Gurvich, and L. Khachiyan. An efficient incremental algorithm for generating all maximal independent sets in hypergraphs of bounded dimension. Parallel Processing Letters, 10 (4):253–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1004/S0129626400000251.
                                        html.

Berthome:1992:EKH

[Ber92]  P. Berthome.  Efficient K-

selection in hypercube multiprocessors. Parallel Processing Letters, 2(2-3):221–230, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Blazewicz:2004:ENS

[BET04] Jacek Błażewicz, Klaus Ecker, and Denis Trystram. Editorial note: “Part I: the Best Papers from CIRM Workshop (October 2001)”. Parallel Processing Letters, 14(1):1–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Barlos:1994:JWP

[BF94] Fotios Barlos and Ophir Frieder. Join workload partitioning under uniform and skewed input relations. Parallel Processing Letters, 4(1-2):95–104, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Bodeveix:1997:TAV

[BF97] J.-P. Bodeveix and M. Filali.

Towards the automatic verification of atomic memory protocols. Parallel Processing Letters, 7(1):101–112, March 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Barth:2002:CMA

[BF02] D. Barth and P. Fragopoulou.

Compact multicast algorithms on grids and tori without intermediate buffering. Parallel Processing Letters, 12(1):31–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Bastoul:2005:APT

[BF05] Cédric Bastoul and Paul Feautrier.

Adjusting a program transformation for legality. Parallel Processing Letters, 15(1/2):3– 17, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Balasa:1994:TNL

[BFCD94] Florin Balasa, Frank H. M.

Franssen, Francky V. M. Catthoor, and Hugo J. De Man. Transformation of nested loops with modulo indexing to affine recurrences. Parallel Processing Letters, 4(3):271–280, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Balaton:2008:ECB

[BFG^{+}08] Zoltan Balaton, Zoltan Farkas, Gabor Gombas, Peter Kacsuk, Roberto Lovas, Attila Csaba Marosi, Ad Emmen, Gabor Terstyanszky, Tamas Kiss, Ian Kelley, Ian Taylor, Oleg Lodygensky, Miguel Cardenas-Montes, Gilles Fedak, and Filipe Araujo. EDGeS: The common boundary between service and desktop Grids. Parallel Processing Letters, 18(3):433–445, September 2008. CODEN PPLTEE. ISSN 0129-6264.

Bernard:2010:CLG

[BGJ10] Thomas A. M. Bernard, Clemens Grelck, and Chris R. Jesshope. On the compilation of a language for general concurrent target architectures. Parallel Processing Letters, 20(1):51–69, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Bampis:1998:PAF

[BGK^{+}98] E. Bampis, A. Giannakos, A. Karzanov, Y. Manoussakis, and I. Milis. A parallel algorithm for finding a perfect matching in a planar graph. Parallel Processing Letters, 8(3):399– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Bampis:2000:UDM

[BGK00] E. Bampis, R. Giroudeau, and J.-C. König. Using duplication for the multiprocessor scheduling problem with hierarchical communications. Parallel Processing Letters, 10(1):133–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000135.
html.

Bischof:2003:COP

[BGK03] Holger Bischof, Sergei Gorlatch, and Emanuel Kitzelmann.

Cost optimality and predictability of parallel programming with skeletons. Parallel Processing Letters, 13(4):575–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Bermond:2006:HAG

[BGK^{+}06] Jean-Claude Bermond, Jerome Galtier, Ralf Klasing, Nelson Morales, and Stephane Perennes. Hardness and approximation of gathering in static radio networks. Parallel Processing Letters, 16(2):165–183, June 2006. CODEN PPLTEE. ISSN 0129-6264.

Bischof:2003:DDP

[BGL03] Holger Bischof, Sergei Gorlatch, and Roman Leshchinskiy. Dat-TeL: A data-parallel C++ template library. Parallel Processing Letters, 13(3):461–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Bischof:2005:DPC

[BGLM05] Holger Bischof, Sergei Gorlatch, Roman Leshchinskiy, and Jens Müller. Data parallelism in C++ template programs: a Barnes–Hut case study. Parallel Processing Letters, 15(3): 257–272, September 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Blazewicz:2000:SCT

[BGPT00] J. Blazewicz, F. Guinand, B. Penz, and D. Trystram. Scheduling complete trees on two uniform processors with integer speed ratios and communication delays. Parallel Processing Letters, 10 (4):267–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1004/S0129626400000263.
                                        html.

Bruck:1993:EGC

[BH93] J. Bruck and Ching-Tien Ho. Efficient global combine operations in multi-port message-passing systems. Parallel Processing Letters, 3(4):335–346, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Breslauer:1996:OPC

[BH96] Dany Breslauer and Ramesh Hariharan. Optimal parallel construction of minimal suffix and factor automata. Parallel Processing Letters, 6(1):35–44, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Bollig:1995:ACC

[BHPS95] Beate Bollig, Martin Huhne, Stefan Polt, and Petr Savicky. On the average case circuit delay of disjunction. Parallel Processing Letters, 5(2):275–280, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Bermond:1992:CLP

[BHQ92] J.-C. Bermond, P. Hell, and J.-J. Quisquater. Construction of large packet radio networks. Parallel Processing Letters, 2(1): 3–12, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Bernaschi:2005:ERA

[BIC05] Massimo Bernaschi, Giulio Iannello, and Saverio Crea. Experimental results about MPI collective communication operations. Parallel Processing Letters, 15(1/2):223–236, March/ June 2005. CODEN PPLTEE. ISSN 0129-6264.

Bossomaier:1992:DPC

[BIL92] T. Bossomaier, N. Isidoro, and A. Loeff. Data parallel computation of Euclidean distance transforms. Parallel Processing Letters, 2(4):331–339, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Banu:2011:ADE

[BIW11] Nazreen Banu, Taisuke Izumi, and Koichi Wada. Adaptive and doubly-expedited one-step consensus in Byzantine asynchronous systems. Parallel Processing Letters, 21(4):461– 477, December 2011. CODEN PPLTEE. ISSN 0129-6264.

Bouabdallah:1992:IMM

[BK92] A. Bouabdallah and J.-C.

Konig. An improvement of Maekawa’s mutual exclusion algorithm to make it faulttolerant. Parallel Processing Letters, 2(2-3):283–290, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Bhatele:2008:BTA

[BK08] Abhinav Bhatelé and Laxmikant V.

Kalé. Benefits of topology aware mapping for mesh interconnects.

Parallel Processing Letters, 18 (4):549–566, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Bhatele:2009:QNC

[BK09] Abhinav Bhatelé and Laxmikant V. Kalé. Quantifying network contention on large parallel machines. Parallel Processing Letters, 19(4):553–572, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Beame:1994:IBE

[BKK94] Paul Beame, Miroslaw Kutylowski, and Marcin Kik. Information broadcasting by exclusiveread PRAMs. Parallel Processing Letters, 4(1-2):159–169, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Bampis:1992:LOS

[BKT92] E. Bampis, J.-C. Konig, and D. Trystram. A low overhead schedule for a 3D-grid graph. Parallel Processing Letters, 2(4): 363–372, December 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Bhattacharya:1995:QLI

[BKT95] S. Bhattacharya, S. Kirani, and W. T. Tsai. Quadtree layouts and I/O bandwidth. Parallel Processing Letters, 5(2): 231–240, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Brunie:1998:MPR

[BKW98] Lionel Brunie, Harald Kosch, and Wolfgang Wohner. From the modeling of parallel relational query processing to query optimization and simulation. Parallel Processing Letters, 8(1): 51–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Barnett:1992:UPL

[BL92] M. Barnett and C. Lengauer.

Unimodularity and the parallelization of loops. Parallel Processing Letters, 2(2-3):273– 281, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Berthold:2008:IDC

[BL08] Jost Berthold and Rita Loogen.

The impact of dynamic channels on functional topology skeletons. Parallel Processing Letters, 18(1):101–115, March 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Brahim:2012:DCS

[BLL12] Abdelhamid Salah Brahim, Bénédicte Le Grand, and Matthieu Latapy. Diffusion cascades: Spreading phenomena in blog network communities. Parallel Processing Letters, 22(1): 1240002, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Beaumont:2003:SSM

[BLMR03] O. Beaumont, A. Legrand, L. Marchal, and Y. Robert. Scheduling strategies for mixed data and task parallelism on heterogeneous clusters. Parallel Processing Letters, 13(2): 225–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Bahn:2008:DAM

[BLY^{+}08] Jun Ho Bahn, Seung Eun Lee, Yoon Seok Yang, Jungsook Yang, and Nader Bagherzadeh. On design and application mapping of a network-on-chip (NOC) architecture. Parallel Processing Letters, 18(2): 239–255, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Berkman:1995:FPA

[BM95] Omer Berkman and Yossi Matias. Fast parallel algorithms for minimum and related problems with small integer inputs. Parallel Processing Letters, 5(2): 223–230, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Badache:1997:GDC

[BM97] Nadjib Badache and Aomar Maddi. Gradual design of a causal broadcast protocol. Parallel Processing Letters, 7(3): 309–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Barriere:2018:OMR

[BMFU18] Lali Barrière, Xavier Muñoz, Janosch Fuchs, and Walter Unger. Online matching in regular bipartite graphs. Parallel Processing Letters, 28 (02):??, June 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500081.

Benoit:2013:APE

[BMRGR13] Anne Benoit, Rami Melhem, Paul Renaud-Goud, and Yves Robert. Assessing the performance of energy-aware mappings. Parallel Processing Letters, 23(2):1340003, June 2013.

CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Blazewicz:2004:PTS

[BMSW04] Jacek Błażewicz, Adrian Moret-Salvador, and Rafał Walkowiak. Parallel Tabu search approaches for two-dimensional cutting. Parallel Processing Letters, 14 (1):23–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Bar-Noy:1993:OAC

[BNKS93] A. Bar-Noy, S. Kipnis, and B. Schieber. An optimal algorithm for computing census functions in message-passing systems. Parallel Processing Letters, 3(1):19–23, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Brevik:2007:UMB

[BNW07] John Brevik, Daniel Nurmi, and Rich Wolski. Using model-based clustering to improve predictions for queueing delay on parallel machines. Parallel Processing Letters, 17(1):21–46, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Bhagavathi:1994:ULS

[BOSW94] D. Bhagavathi, S. Olariu, W. Shen, and L. Wilson. A unifying look at semigroup computations on meshes with multiple broadcasting. Parallel Processing Letters, 4(1-2):73–82, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Bhagavathi:1992:CPP

[BOSZ92] D. Bhagavathi, S. Olariu, J. L. Schwing, and J. Zhang. Convex polygon problems on meshes with multiple broadcasting. Parallel Processing Letters, 2 (2-3):249–256, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Becka:2015:NDO

[BOV15] Martin Bečka, Gabriel Okša, and Marián Vajteršic. New dynamic orderings for the parallel one–sided block-Jacobi SVD algorithm. Parallel Processing Letters, 25(2):1550003, June 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Boxer:2009:ECG

[Box09] Laurence Boxer. Efficient coarse grained permutation exchanges and matrix multiplication. Parallel Processing Letters, 19(3): 477–484, September 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Boxer:2021:CGP

[Box21]  Laurence Boxer.  Coarse

grained parallel selection. Parallel Processing Letters, 31 (01):??, March 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500031.

Banerjee:2019:SNT

[BR19] Avah Banerjee and Dana Richards. A sorting network on trees. Parallel Processing Letters, 29(04):??, December 2019. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500154.

Brandes:2000:HLC

[Bra00] T. Brandes. HPF library and compiler support for halos in data parallel irregular computations. Parallel Processing Letters, 10(2/3):189–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000196.
html.

Benoit:2009:CML

[BRT09] Anne Benoit, Yves Robert, and Eric Thierry. On the complexity of mapping linear chain applications onto heterogeneous platforms. Parallel Processing Letters, 19(3):383–397, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Baker:2001:ERJ

[BS01] M. Baker and G. Smith. Establishing a reliable Jini infrastructure for parallel applications. Parallel Processing Letters, 11 (2–3):203–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Bandyopadhyay:2016:PCA

[BSM^{+}16] Soumyadip Bandyopadhyay, Dipankar Sarkar, Chittaranjan Mandal, Kunal Banerjee, and Krishnam Raju Duddu. A path construction algorithm for translation validation using PRES+ models. Parallel Processing Letters, 26(2):1650010, June 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Burkhardt:2021:GCL

[Bur21] Paul Burkhardt. Graph connectivity in log steps using label propagation. Parallel Processing Letters, 31(04): ??, December 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500213.

Barrett:2013:RBB

[BVHR13] R. F. Barrett, C. T. Vaughan, S. D. Hammond, and D. Roweth. Reducing the bulk in the bulk synchronous parallel model. Parallel Processing Letters, 23 (4):1340010, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Bruda:2016:LCT

[BW16] Stefan D. Bruda and Mary Sarah Ruth Wilkin. Limitations of coverability trees for context-free parallel communicating grammar systems and why these grammar systems are not linear space. Parallel Processing Letters, 26(3): 1650012, September 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chong:1999:SMV

[CA99] Frederic T. Chong and Anant Agarwal. Shared memory versus message passing for iterative solution of sparse, irregular problems. Parallel Processing Letters, 9(1):159–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Campa:2008:ECC

[Cam08] Sonia Campa. Evaluating computational costs while handling data and control parallelism. Parallel Processing Letters, 18 (1):165–174, March 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Carbajal:2007:PTD

[Car07] Santiago Garcia Carbajal. Parallelizing three dimensional cellular automata with OpenMP. Parallel Processing Letters, 17 (4):349–361, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Charron-Bost:2003:NLG

[CBC03] Bernadette Charron-Bost and Robert Cori. A note on linearizability and the global time axiom. Parallel Processing Letters, 13(1):19–??, March 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Carino:2005:PAQ

[CBV^{+}05] Ricolindo L. Cariño, Ioana Banicescu, Ravi K. Vadapalli, Charles A. Weatherford, and Jianping Zhu. Parallel adaptive quantum trajectory method for wavepacket simulations. Parallel Processing Letters, 15(4):

415–422, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Carlsson:1992:PCH

[CC92] S. Carlsson and J. Chen. Parallel constructions of heaps and minmax heaps. Parallel Processing Letters, 2(4):311–320, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Chung:1993:NPP

[CC93] Kuo-Liang Chung and Hsun-Wen Chang. Novel pipelining and processor allocation strategy for monoid computations on unshuffle-exchange networks. Parallel Processing Letters, 3(2): 189–193, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Chan:1995:OSD

[CC95] I. W. Chan and F. Choi. An optimal systolic dictionary. Parallel Processing Letters, 5(3):451– 460, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Ceres:2011:FTS

[CC11] Pietro Ceres and Raffaele Ceres. On the fault-tolerance and size of WDM optical networks. Parallel Processing Letters, 21(1): 3–12, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Coccimiglio:2017:PLS

[CC17] Gaetano Coccimiglio and Salimur Choudhury. A parallel local search algorithm for clustering large biological networks. Parallel Processing Letters, 27 (3–4):1750007, 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chiola:2001:UCF

[CCA01] G. Chiola, G. Ciaccio, and C. Anglano. On the use of coscheduling in fast communication systems. Parallel Processing Letters, 11(2–3):223–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Cannataro:2004:PBP

[CCCV04] Mario Cannataro, Carmela Comito, Antonio Congiusta, and Pierangelo Veltri. PROTEUS: a bioinformatics problem solving environment on grids. Parallel Processing Letters, 14(2): 217–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Calude:2013:ICP

[CCQ13] Cristian S. Calude, Elena Calude, and Melissa S. Queen. Inductive complexity of the P versus NP problem. Parallel Processing Letters, 23(1): 1350007, March 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Cappello:2011:PMV

[CCR11] Franck Cappello, Henri Casanova, and Yves Robert. Preventive migration vs. preventive checkpointing for extreme scale supercomputers. Parallel Processing Letters, 21(2):111–132, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Chan:1999:NCG

[CD99] A. Chan and F. Dehne. A note on coarse grained parallel integer sorting. Parallel Processing Letters, 9(4):533–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Chatzigiannakis:2003:CSP

[CDM^{+}03] Ioannis Chatzigiannakis, Tassos Dimitriou, Marios Mavronicolas, Sotiris Nikoletseas, and Paul Spirakis. A comparative study of protocols for efficient data propagation in smart dust networks. Parallel Processing Letters, 13 (4):615–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Coudert:2007:SRR

[CDP^{+}07] D. Coudert, P. Datta, S. Perennes, H. Rivano, and M.-E. Voge. Shared risk resource group complexity and approximability issues. Parallel Processing Letters, 17(2):169–184, June 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Caron:2010:SSP

[CDPT10] Eddy Caron, Frédéric Desprez, Franck Petit, and Cédric Tedeschi. Snap-stabilizing prefix tree for peer-to-peer systems. Parallel Processing Letters, 20 (1):15–30, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Caceres:1993:FET

[CDSS93] E. N. Caceres, N. Deo, S. Sastry, and J. L. Szwarcfiter. On finding Euler Tours in parallel. Parallel Processing Letters, 3(3):223– 231, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Chaudhuri:1996:PCA

[CDZ96] S. Chaudhuri, Y. Dimopoulos, and C. D. Zaroliagis. On the parallel complexity of acyclic logic programs. Parallel Processing Letters, 6(2):223–230, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Coffman:1998:NLP

[CE98] E. G. Coffman, Jr. and S. Even.

A note on limited preemption. Parallel Processing Letters, 8(1): 3–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:1991:RPL

[CF91] M. Cosnard and A. Ferreira. On the real power of loosely coupled parallel architectures. Parallel Processing Letters, 1(2):103– 111, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Charles:1993:SSS

[CF93] H.-P. Charles and P. Fraigniaud. Scheduling a scatteringgathering sequence on hypercubes. Parallel Processing Letters, 3(1):29–42, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Correa:1995:ESP

[CF95] R. Correa and A. Ferreira. On the effectiveness of synchronous parallel branch-and-bound algorithms. Parallel Processing Letters, 5(3):375–386, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Chailloux:2003:PIO

[CF03] Emmanuel Chailloux and Christian Foisy. A portable implementation for Objective Caml Flight. Parallel Processing Letters, 13(3):425–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Comellas:1994:LVS

[CFG94] F. Comellas, M. A. Fiol, and J. Gomez. On large vertex symmetric 2-reachable digraphs. Parallel Processing Letters, 4(4):379–384, December 1994. CODEN PPLTEE. ISSN 0129-6264. URL http:/

/www-mat.upc.es/~comellas/
2reach/2reach.html.

Comin:2004:GDB

[CFG04] Matteo Comin, Carlo Ferrari, and Concettina Guerra. Grid deployment of bioinformatics applications: a case study in protein similarity determination. Parallel Processing Letters, 14 (2):163–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Collins:2012:ATP

[CFL12] Alexander Collins, Christian Fensch, and Hugh Leather. Auto-tuning parallel skeletons. Parallel Processing Letters, 22 (2):1240005, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Charpentier:1999:OAC

[CFM^{+}99] Michel Charpentier, Mamoun Filali, Philippe Mauran, GéRard Padiou, and Philippe QuéInnec.

The observation: An abstract communication mechanism. Parallel Processing Letters, 9(3):437–??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Calheiros:2009:SMV

[CFR09] Rodrigo N. Calheiros, Tiago Ferreto, and César A. F. De Rose. Scheduling and management of virtual resources in Grid sites: the site resource scheduler. Parallel Processing Letters, 19(1): 3–18, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Collard:1997:ADA

[CG97] Jean-Fançois Collard and Martin Griebl. Array dataflow analysis for explicitly parallel programs. Parallel Processing Letters, 7(2):117–131, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Chen:1998:IPB

[CG98a] Yung-Syau Chen and Jean-Luc Gaudiot. Implementing parallel branch-and-bound with extended Sisal 2.0. Parallel Processing Letters, 8(1):41–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Comellas:1998:GPD

[CG98b] F. Comellas and G. Giménez.

Genetic programming to design communication algorithms for parallel architectures. Parallel Processing Letters, 8(2): 549–??, December 1998. CO-DEN PPLTEE. ISSN 0129-6264.

                                        URL http://www-mat.upc.es/
                                        ~comellas/genprog/genprog.
                                        html.

Cerin:2001:BCW

[CG01] C. Cérin and J.-L. Gaudiot. Benchmarking clusters of workstations through parallel sorting and BSP libraries. Parallel Processing Letters, 11(1):25–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Cerin:2006:MSS

[CGK06] Christophe Cérin, Jean-Luc Gaudiot, and Michel Koskas. A multithreaded SQL service. Parallel Processing Letters, 16(2): 245–259, June 2006. CODEN PPLTEE. ISSN 0129-6264.

Crain:2016:FCF

[CGR16] Tyler Crain, Vincent Gramoli, and Michel Raynal. A fast contention-friendly binary search tree. Parallel Processing Letters, 26(3):1650015, September 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Casanova:2015:TMS

[CGS15] Henri Casanova, Anshul Gupta, and Frédéric Suter. Toward more scalable off-line simulations of MPI applications. Parallel Processing Letters, 25(3): 1541002, September 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chang:2005:SIS

[CGW05] Weng-Long Chang, Minyi Guo, and Jesse Wu. Solving the independent-set problem in a DNA-based supercomputer model. Parallel Processing Letters, 15(4):469–479, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Chou:1994:EFT

[CH94] Ray-Shyng Chou and Lih-Hsing Hsu. 1-edge fault-tolerant designs for meshes. Parallel Processing Letters, 4(4):385– 389, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Chen:1996:FEO

[CH96] D. Z. Chen and Xiaobo Hu.

Fast and efficient operations on parallel priority queues. Parallel Processing Letters, 6(4):451– 467, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Chung:2003:RSA

[CH03] I-Hsin Chung and Jeffrey K.

Hollingsworth. Runtime selection among different API implementations. Parallel Processing Letters, 13(2):123–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Chiao:2003:SSP

[Chi03] Yiwei Chiao. A self-stabilizing phase synchronization protocol. Parallel Processing Letters, 13 (1):25–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Chlebus:1992:TSA

[Chl92] B. Chlebus. Two selection algorithms on a mesh-connected computer. Parallel Processing Letters, 2(4):341–346, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Cholvi:1998:SBM

[Cho98] Vicent Cholvi. Specification of the behavior of memory operations in distributed systems. Parallel Processing Letters, 8(2): 589–??, December 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Claver:1996:SDL

[CHQ96] J. M. Claver, V. Hernandez, and E. S. Quintana. Solving discretetime Lyapunov equations for the Cholesky factor on a shared memory multiprocessor. Parallel Processing Letters, 6(3):365– 376, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Chung:1996:ITM

[Chu96] Kuo-Liang Chung. Image template matching on reconfigurable meshes. Parallel Processing Letters, 6(3):345–353, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:2001:APT

[CJ01] M. Cosnard and E. Jeannot. Automatic parallelization techniques based on compact DAG extraction and symbolic scheduling. Parallel Processing Letters, 11(1):151–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Chen:1999:PFP

[CJN99] Lin Chen, Julius Y. Jiang, and Maung T. Nyeu. Provably fastest parallel algorithms for bipartite permutation graphs.

Parallel Processing Letters, 9(3): 385–??, September 1999. CO-DEN PPLTEE. ISSN 0129-6264.

Cai:1993:DBS

[CK93] Leizhen Cai and J. M. Keil.

Degree-bounded spanners. Parallel Processing Letters, 3(4): 457–468, December 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Caragiannis:2000:SCA

[CKP00] I. Caragiannis, C. Kaklamanis, and P. Persiano. Symmetric communication in alloptical tree networks. Parallel Processing Letters, 10 (4):305–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000299.
html.

Cheung:1993:TLB

[CL93] S. Cheung and F. C. M. Lau.

Time lower bounds for permutation routing on multidimensional bused meshes. Parallel Processing Letters, 3(2): 129–138, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:1995:ATG

[CL95] M. Cosnard and M. Loi. Automatic task graph generation techniques. Parallel Processing Letters, 5(4):527–538, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Cheng:2013:CMP

[CL13] Eddie Cheng and László Lipták.

Conditional matching preclusion for (n, k)-star graphs. Parallel Processing Letters, 23(1): 1350004, March 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chong:2013:ASG

[CLH13] Song-Kong Chong, Cheng-Chi Lee, and Min-Shiang Hwang. An authentication scheme for the global mobility network. Parallel Processing Letters, 23(3): 1350009, September 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chande:2016:NSC

[CLL16] Manoj Kumar Chande, Cheng-Chi Lee, and Chun-Ta Li. A new self-certified convertible authenticated encryption scheme based on discrete logarithm problem. Parallel Processing Letters, 26 (4):1650018, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chandrasekharan:1993:EPA

[CLM93] N. Chandrasekharan, V. S. Lakshmanan, and M. Medidi. Efficient parallel algorithms for finding chordless cycles in graphs. Parallel Processing Letters, 3(2): 165–170, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Cohen:2016:SSA

[CLM^{+}16] Johanne Cohen, Jonas Lefèvre, Khaled Maâmra, Laurence Pilard, and Devan Sohier. A selfstabilizing algorithm for maximal matching in anonymous networks. Parallel Processing Letters, 26(4):1650016, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Clarke:2011:DLB

[CLR11] David Clarke, Alexey Lastovetsky, and Vladimir Rychkov. Dynamic load balancing of parallel computational iterative routines on highly heterogeneous HPC platforms. Parallel Processing Letters, 21(2):195–217, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Carrington:2013:CLS

[CLT13] Laura Carrington, Michael Laurenzano, and Ananta Tiwari. Characterizing large-scale HPC applications through trace extrapolation. Parallel Processing Letters, 23(4):1340008, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Corradi:1992:LBS

[CLZ92] A. Corradi, L. Leonardi, and F. Zambonelli. Load balancing strategies for massively parallel architectures. Parallel Processing Letters, 2(2-3):139– 148, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Chung:2011:HMH

[CLZC11] I-Hsin Chung, Che-Rung Lee, Jiazheng Zhou, and Yeh-Ching Chung. Hierarchical mapping for HPC applications. Parallel Processing Letters, 21(3):

279–299, September 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Chandra:1995:RFM

[CM95] A. Chandra and R. Melhem. Reconfiguration in fault-tolerant 3D meshes. Parallel Processing Letters, 5(3):387–399, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Carruth:1996:PRM

[CM96] J. Allen Carruth and J. Misra. Proof of a real-time mutualexclusion algorithm. Parallel Processing Letters, 6(2):251– 257, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Chang:1997:ASB

[CM97] Chihming Chang and Rami Melhem. Arbitrary size Benes networks. Parallel Processing Letters, 7(3):279–??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Carroll:2019:ACM

[CmL19a] Shane Carroll and Wei ming Lin. Applied on-chip machine learning for dynamic resource control in multithreaded processors. Parallel Processing Letters, 29(03):??, September 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626419500130.

Carroll:2019:RRT

[CmL19b] Shane Carroll and Wei ming Lin. Round robin thread selection optimization in multithreaded processors. Parallel Processing Letters, 29(01):??, March 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500038.

Carroll:2021:ELT

[CmL21] Shane Carroll and Wei ming Lin. Exploiting long-term temporal cache access patterns for LRU insertion prioritization. Parallel Processing Letters, 31(02):??, June 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500109.

Carle:1999:AAB

[CMS99] J. Carle, J.-F. Myoupo, and D. Seme. All-to-all broadcasting algorithms on honeycomb networks and applications. Parallel Processing Letters, 9(4):539–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Cole:1995:PPL

[Col95] Murray I. Cole. Parallel programming with list homomorphisms. Parallel Processing Letters, 5(2):191–203, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:1993:EN

[Cos93] M. Cosnard. Editorial note.

Parallel Processing Letters, 3(1):

??, March 1993. URL http:/

                                        /www.ens-lyon.fr/PPL/vol3_
                                        1.html.

Cosnard:1995:EN

[Cos95] M. Cosnard. Editorial note. Parallel Processing Letters, 5(2): 125–??, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:1996:ENa

[Cos96a] M. Cosnard. Editorial note. Parallel Processing Letters, 6(1): ??, March 1996. URL http:/

                                        /www.ens-lyon.fr/PPL/vol6_
                                        1.html.

Cosnard:1996:ENb

[Cos96b] M. Cosnard. Editorial note. Parallel Processing Letters, 6(2): ??, June 1996. URL http:/

                                        /www.ens-lyon.fr/PPL/vol6_
                                        2.html.

Cosnard:1997:EN

[Cos97] M. Cosnard. Editorial note. Parallel Processing Letters, 7(3): 217–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Cosnard:1999:EN

[Cos99] M. Cosnard. Editorial note. Parallel Processing Letters, 9(4): 451–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

                                        6264.  URL http://www.
                                        wspc.com/journals/ppl/09/
                                        0904/0002.html;   http://
                                        www.wspc.com/journals/ppl/
                                        09/0904/1190904.

Cosnard:2000:EN

[Cos00] M. Cosnard. Editorial note. Parallel Processing Letters, 10 (1):1–??, March 2000. CO-DEN PPLTEE. ISSN 0129-6264.

URL http://ejournals.wspc.
com.sg/ppl/10/1001/.

Cosnard:2001:EN

[Cos01] M. Cosnard. Editorial note.

Parallel Processing Letters, 11 (2–3):311–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Cosnard:2003:ENS

[Cos03] Michel Cosnard. Editorial note:

“Part II: Regular Papers”. Parallel Processing Letters, 13(4): 511–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Cosnard:2004:ENS

[Cos04] Michel Cosnard. Editorial note:

“Part II: Regular Papers”. Parallel Processing Letters, 14(1): 2–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Costa:2015:UT

[Cos15a] José Félix Costa. Uncertainty in time. Parallel Processing Letters, 25(1):1540007, March 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Costello:2015:CVD

[Cos15b] Ben De Lacy Costello. Calculating Voronoi diagrams using simple chemical reactions. Parallel Processing Letters, 25 (1):1540003, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Charpentier:1998:SVAa

[CP98a] Michel Charpentier and Gérard Padiou. Specification and verification of the ATMR protocol using unity part 1: Modeling and specification of the protocal. Parallel Processing Letters, 8(4): 421–??, December 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Charpentier:1998:SVAb

[CP98b] Michel Charpentier and Gérard Padiou. Specification and verification of the ATMR protocol using unity part 2: Correctness proof. Parallel Processing Letters, 8(4):433–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Choi:1998:SDI

[CP98c] Jong Hyuk Choi and Kyu Ho Park. Segment directory: An improvement to the pointer in directory cache coherence schemes. Parallel Processing Letters, 8(2):577–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Cheng:2012:MPC

[CP12] Eddie Cheng and Sachin Padmanabhan. Matching preclusion and conditional matching preclusion for crossed cubes. Parallel Processing Letters, 22 (2):1250005, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Cheng:2009:SNS

[CQS09] Eddie Cheng, Ke Qiu, and Zhizhang Shen. A short note on the surface area of star graphs. Parallel Processing Letters, 19 (1):19–22, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Cheng:2011:SAA

[CQS11] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the surface areas and average distances of meshes and tori. Parallel Processing Letters, 21(1):61–75, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Cheng:2016:CDH

[CQS16] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the conditional diagnosability of hyperbuttery graphs and related networks. Parallel Processing Letters, 26(1):1650005, March 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Cheng:2020:BAD

[CQSY20a] Eddie Cheng, Ke Qiu, Zhizhang Shen, and Weihua Yang. A brief account on the development and future research directions of connectivity properties of interconnection networks. Parallel Processing Letters, 30(03):??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400095.

Cheng:2020:P

[CQSY20b] Eddie Cheng, Ke Qiu, Zhizhang Shen, and Weihua Yang. Preface. Parallel Processing Letters, 30(03):??, September 2020.

CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420020016.

Chu:1995:ETD

[CR95] Yung-Kang Chu and D. T. Rover. An effective twodimensional mesh partitioning strategy. Parallel Processing Letters, 5(4):623–634, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Collard:1995:CDL

[CRF95] J. Collard, T. Risset, and P. Feautrier. Construction of DO loops from systems of affine constraints. Parallel Processing Letters, 5(3):421–436, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Chailloux:2008:HFT

[CRV08] Emmanuel Chailloux, Vivien Ravet, and Julien Verlaguet. HirondML: Fair threads migrations for Objective Caml. Parallel Processing Letters, 18(1): 55–69, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Cores:2003:IBE

[CRY^{+}03] F. Cores, A. Ripoll, X. Y. Yang, B. Qazzaz, R. Suppi, P. Hernandez, and E. Luque. Improving bandwidth efficiency in distributed video-on-demand architectures. Parallel Processing Letters, 13(4):589–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Cai:1995:CRU

[CS95] Wentong Cai and David B. Skillicorn. Calculating recurrences using the Bird-Meertens formalism. Parallel Processing Letters, 5(2):179–190, June 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Chen:2001:WSB

[CS01] Y. Chen and J. W. Sanders.

Weakest specifunctions for BSP. Parallel Processing Letters, 11 (4):439–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Chen:2003:TDB

[CS03] Yifeng Chen and J. W. Sanders.

Top-down design of bulksynchronous parallel programs. Parallel Processing Letters, 13 (3):389–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Cheng:2013:GCG

[CS13] Eddie Cheng and Nart Shawash.

The Q_{n,k,m} graph: a common generalization of various popular interconnection networks. Parallel Processing Letters, 23(3): 1350011, September 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chong:1995:MRS

[CSBS95] F. T. Chong, S. D. Sharma, E. A. Brewer, and J. Saltz. Multiprocessor runtime support for fine-grained, irregular DAGs. Parallel Processing Letters, 5(4): 671–683, December 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Costa:2008:ODD

[CSFK08] Fernando Costa, Luis Silva, Gilles Fedak, and Ian Kelley. Optimizing data distribution in desktop Grid platforms. Parallel Processing Letters, 18(3): 391–410, September 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Chung:2000:ECM

[CSK00] C.-M. Chung, P.-S. Shin, and S.-D. Kim. An effective configuration method for Java– Internet computing environment. Parallel Processing Letters, 10(1):73–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000093.
html.

Clausen:1994:DIS

[CT94] Jens Clausen and Jesper Larsson Traeff. Do inherently sequential branch-and-bound algorithms exist? Parallel Processing Letters, 4(1-2):3–13, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Cachera:1996:PDP

[CU96] D. Cachera and G. Utard. Proving data-parallel programs: a unifying approach. Parallel Processing Letters, 6(4):491– 505, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Cao:2021:IAL

[CWR21] Yang Cao, Fei Wu, and Thomas Robertazzi. Integrating Amdahl-like laws and divisible load theory. Parallel Processing Letters, 31(02):??, June 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500080.

Chartrand:2020:UCG

[CZ20] Gary Chartrand and Ping Zhang. Uniformly connected graphs — a survey. Parallel Processing Letters, 30(03): ??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400022.

Daniel:1993:SPM

[DA93] F. Daniel and G. Authie. Shortest paths multiplicity in generalized de Bruijn and Kautz networks. Parallel Processing Letters, 3(4):363–374, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Duran:2011:OPP

[DAB^{+}11] Alejandro Duran, Eduard Ayguadé, Rosa M. Badia, Jesús Labarta, Luis Martinell, Xavier Martorell, and Judit Planas. OmpSs: a proposal for programming heterogeneous multi-core architectures. Parallel Processing Letters, 21(2):173–193, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Danelutto:2001:ESS

[Dan01] M. Danelutto. Efficient support for skeletons on workstation clusters. Parallel Processing Letters, 11(1):41–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Davis:2009:PPM

[DBK09a] Kei Davis, Kevin J. Barker, and Darren J. Kerbyson. Performance prediction via modeling: a case study of the ORNL Cray XT4 upgrade. Parallel Processing Letters, 19(4):619– 639, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Dursun:2009:MPM

[DBK^{+}09b] Hikmet Dursun, Kevin J. Barker, Darren J. Kerbyson, Scott Pakin, Richard Seymour, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. An MPI performance monitoring interface for cell based compute nodes. Parallel Processing Letters, 19(4):535–552, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Dimakopoulos:1999:BT

[DD99] Vassilios V. Dimakopoulos and Nikitas J. Dimopoulos. On broadcasting time. Parallel Processing Letters, 9(1):3–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Datta:2017:SSW

[DDLV17] Ajoy K. Datta, Stephane Devismes, Lawrence L. Larmore, and Vincent Villain. Selfstabilizing weak leader election in anonymous trees using constant memory per edge. Parallel Processing Letters, 27(2): 1750002, June 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Darte:1996:COM

[DDR96] Alain Darte, Michele Dion, and Yves Robert. A characterization of one-to-one modular mappings. Parallel Processing Letters, 6(1):145–157, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Datta:2017:FPA

[DDS17] Amit Datta, Mallika De, and Bhabani P. Sinha. Fast parallel algorithm for prefix computation in multi-mesh architecture. Parallel Processing Letters, 27(3–4):1750009, 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Desprez:1995:PSL

[DDT95] F. Desprez, J. J. Dongarra, and B. Tourancheau. Performance study of LU factorization with low communication overhead on multiprocessors. Parallel Processing Letters, 5(2): 157–169, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Das:1999:SSA

[DDT99] S. K. Das, A. K. Datta, and S. Tixeuil. Self-stabilizing algorithms in DAG structured networks. Parallel Processing Letters, 9(4):563–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Danelutto:2017:PAS

[DDT17] Marco Danelutto, Daniele De Sensi, and Massimo Torquati. A power-aware, self-adaptive macro data flow framework. Parallel Processing Letters, 27 (1):1740004, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

DeFalco:1992:SGA

[DDTV92] I. De Falco, R. Del Balio, E. Tarantino, and R. Vaccaro. Simulation of genetic algorithms on MIMD multicomputers. Parallel Processing Letters, 2(4): 381–389, December 1992. CO-DEN PPLTEE. ISSN 0129-6264.

DeAgostino:2004:AWO

[De 04] Sergio De Agostino. Almost work-optimal PRAM EREW decoders of LZ compressed text. Parallel Processing Letters, 14(3/4):351–??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264.

Debbabi:1997:MBC

[Deb97] Maurad Debbabi. A modelbased concurrent specification language over CML: Semantic foundations. Parallel Processing Letters, 7(3):329–??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Dec:2021:LCI

[Dec21] Grzegorz Rafał Dec. LSTM cell implementation on FP-GAs. Parallel Processing Letters, 31(02):??, June 2021.

ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500110.

Drews:2004:SWB

[DEKS04] F. Drews, K. Ecker, O. Kao, and S. Schomann. Strategies for workload balancing in clusterbased image databases. Parallel Processing Letters, 14(1):33–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Devismes:2005:SSS

[Dev05] Stéphane Devismes. A silent self-stabilizing algorithm for finding cut-nodes and bridges. Parallel Processing Letters, 15(1/2):183–198, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Das:1999:EPA

[DF99] Sajal K. Das and Paolo Ferragina. An EREW PARM algorithm for updating minimum spanning trees. Parallel Processing Letters, 9(1):111–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Diallo:2001:NCE

[DFRC01] M. Diallo, A. Ferreira, and A. Rau-Chaplin. A note on communication-efficient deterministic parallel algorithms for planar point location and 2D Voronoï diagram. Parallel Processing Letters, 11(2–3):327–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Dash:2010:DMS

[DGB10] Dinesh Dash, Arobinda Gupta, and Arijit Bishnu. Dynamic maintenance of support coverage in sensor networks. Parallel Processing Letters, 20(2): 155–172, June 2010. CODEN PPLTEE. ISSN 0129-6264.

Drach:1995:ACA

[DGJS95] N. Drach, A. Gefflaut, P. Joubert, and A. Seznec. About cache associativity in lowcost shared memory multimicroprocessors. Parallel Processing Letters, 5(3):475–487, September 1995. CODEN PPLTEE. ISSN 0129-6264.

delaTorre:1992:OTR

[dGP92] P. de la Torre, R. Greenlaw, and T. M. Przytycka. Optimal tree ranking is in NC. Parallel Processing Letters, 2(1):31–41, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Dorta:2003:LPS

[DGRD03] Antonio J. Dorta, Jesús A. González, Casiano Rodríguez, and Francisco De Sande. LLC: a parallel skeletal language. Parallel Processing Letters, 13(3): 437–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Datta:1998:SSA

[DGT98] Ajoy K. Datta, Teofilo F. Gonzalez, and Visalakshi Thiagarajan. Self-stabilizing algorithms for tree metrics. Parallel Processing Letters, 8(1):121–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

DiazDeCerio:1996:CPH

[DGVG96] L. Diaz De Cerio, A. Gonzalez, and M. Valero-Garcia. Communication pipelining in hypercubes. Parallel Processing Letters, 6(4):507–523, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Das:1994:SDR

[DH94] Sajal K. Das and Ranette H.

Halverson. Simple deterministic and randomized algorithms for linked list ranking on the EREW PRAM model. Parallel Processing Letters, 4(1-2):15–27, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Doerner:2006:PCS

[DHBL06] Karl F. Doerner, Richard F.

Hartl, Siegfried Benkner, and Maria Lucka. Parallel cooperative savings based ant colony optimization — multiple search and decomposition approaches. Parallel Processing Letters, 16 (3):351–369, September 2006. CODEN PPLTEE. ISSN 0129-6264.

DiIanni:2000:WDP

[Di 00] M. Di Ianni. Wormhole deadlock prediction. Parallel Processing Letters, 10(4):295–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000287.
html.

Datta:1991:MMT

[DJM91] A. Datta, S. V. Joshi, and R. N. Mahapatra. Modelling a morphological thinning algorithm for shared memory SIMD computers. Parallel Processing Letters, 1(1):59–65, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Deelman:2013:HSM

[DJMN13] Ewa Deelman, Gideon Juve, Maciej Malawski, and Jarek Nabrzyski. Hosted science: Managing computational workflows in the cloud. Parallel Processing Letters, 23(2):1340004, June 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Diks:1998:BUT

[DKP98] Krzysztof Diks, Evangelos Kranakis, and Andrzej Pelc. Broadcasting in unlabeled tori. Parallel Processing Letters, 8(2): 177–188, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Darte:1991:LSN

[DKR91] A. Darte, L. Khachiyan, and Y. Robert. Linear scheduling is nearly optimal. Parallel Processing Letters, 1(2):73–81, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Dessmark:2003:TOB

[DLP03] Anders Dessmark, Andrzej Lingas, and Andrzej Pelc. Tradeoffs between load and degree in virtual path layouts. Parallel Processing Letters, 13(3): 485–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

DiCosmo:2008:SPP

[DLPW08] Roberto Di Cosmo, Zheng Li, Susanna Pelagatti, and Pierre Weis. Skeletal parallel programming with OcamlP3L 2.0. Parallel Processing Letters, 18(1): 149–164, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Destri:1996:BLA

[DM96] G. Destri and P. Marenzoni.

Benchmarking lattice-based applications on parallel architectures. Parallel Processing Letters, 6(3):309–320, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Dumrauf:2012:CNE

[DM12] Dominic Dumrauf and Burkhard Monien. Computing Nash equilibria for two-player restricted network congestion games is PLS-complete. Parallel Processing Letters, 22(4):1250014, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Diakite:2011:PSS

[DMNP11] Sékou Diakité, Loris Marchal, Jean-Marc Nicod, and Laurent Philippe. Practical steady-state scheduling for tree-shaped task graphs. Parallel Processing Letters, 21(4):397–412, December 2011. CODEN PPLTEE. ISSN 0129-6264.

Diks:1994:RTD

[DMP94] Krzysztof Diks, Adam Malinowski, and Andrzej Pelc. Reliable token dispersal with random faults. Parallel Processing Letters, 4(4):417–427, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Delacourt:1999:CTP

[DMS99] Emmanuel Delacourt, Jean-Frederic Myoupo, and David Seme. A constant time parallel detection of repetitions. Parallel Processing Letters, 9(1):81–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Dediu:2014:GEN

[DMVT14] Adrian-Horia Dediu, Carlos Martín-Vide, and Bianca Truthe. Guest Editors’ note — special issue: First International Conference on the Theory and Practice of Natural Computing, TPNC 2012. Parallel Processing Letters, 24(2):1402001, June 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dolev:1998:OTS

[Dol98] Shlomi Dolev. Optimal time self-stabilization in uniform dynamic systems. Parallel Processing Letters, 8(1):7–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Dong:2007:TTS

[Don07] Fangpeng Dong. A taxonomy of task scheduling algorithms in the Grid. Parallel Processing Letters, 17(4):439–454, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Doroshenko:1992:AAD

[Dor92] A. E. Doroshenko. On asynchronous avoidance of deadlocks in parallel programs. Parallel Processing Letters, 2(2-3):291– 297, September 1992. CODEN PPLTEE. ISSN 0129-6264.

DiCosmo:2003:CDA

[DP03] Roberto Di Cosmo and Susanna Pelagatti. A calculus for dense array distributions. Parallel Processing Letters, 13(3):377– ??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Dieudonne:2009:SR

[DP09] Yoann Dieudonné and Franck Petit. Scatter of robots. Parallel Processing Letters, 19(1): 175–184, March 2009. CODEN PPLTEE. ISSN 0129-6264.

DaRosaRighi:2010:OIM

[DPM^{+}10] Rodrigo Da Rosa Righi, Laércio Lima Pilla, Nicolas Maillard, Alexandre Carissimi, and Philippe Olivier Alexandre Navaux. Observing the impact of multiple metrics and runtime adaptations on BSP process rescheduling. Parallel Processing Letters, 20 (2):123–144, June 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Diaz:2000:FRG

[DPS00] J. Diaz, J. Petit, and M. Serna.

Faulty random geometric networks. Parallel Processing Letters, 10(4):343–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/.

Darte:1994:AP

[DR94] Alain Darte and Yves Robert. On the alignment problem. Parallel Processing Letters, 4(3): 259–270, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

DeMarco:2000:TTB

[DR00] G. De Marco and A. A. Rescigno. Tighter time bounds on broadcasting in torus networks in presence of dynamic faults. Parallel Processing Letters, 10(1):39–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1001/S0129626400000068.
                                        html.

Dikaiakos:1996:FAS

[DRS96] Marios D. Dikaiakos, Anne Rogers, and Kenneth Steiglitz. Functional algorithm simulation of the fast multipole method: Architectural implications. Parallel Processing Letters, 6(1): 55–66, March 1996. CODEN PPLTEE. ISSN 0129-6264.

De:1997:FPM

[DS97] Mallika De and Bhabani P. Sinha. Fast parallel multiplication using redundant quarternary number system. Parallel Processing Letters, 7(1):13–23, March 1997. CODEN PPLTEE. ISSN 0129-6264.

daSilva:2015:OTR

[dSJR^{+}15] Rafael Ferreira da Silva, Gideon Juve, Mats Rynge, Ewa Deelman, and Miron Livny. Online task resource consumption prediction for scientific workflows. Parallel Processing Letters, 25 (3):1541003, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Darte:1997:CRS

[DSV97] Alain Darte, Georges-André Silber, and Frédéric Vivien. Combining retiming and scheduling techniques for loop parallelization and loop tiling. Parallel Processing Letters, 7(4):379– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Djamegni:1996:SDA

[DT96] Clementin Tayou Djamegni and Maurice Tchuente. Scheduling of the DAG associated with pipeline inversion of triangular matrices. Parallel Processing Letters, 6(1):13–26, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Djamegni:2000:NAD

[DT00] C. Tayou Djamegni and M. Tchuente.

A new algorithm for dynamic programming on regular arrays. Parallel Processing Letters, 10(1):15–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000044.
html.

Dongarra:2001:PCC

[DT01] J. Dongarra and B. Tourancheau.

Preface: Clusters and computational grids for scientific computing. Parallel Processing Letters, 11(2–3):185–??, 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Dongarra:2003:P

[DT03] Jack Dongarra and Bernard Tourancheau. Preface. Parallel Processing Letters, 13(2):93–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Dongarra:2011:GEN

[DT11] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note. Parallel Processing Letters, 21(2):109, June 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Dongarra:2013:GEN

[DT13] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note: Special issue on clusters, clouds, and data for scientific computing. Parallel Processing Letters, 23(2):1302001, June 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2015:GEN

[DT15] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note: Special issue on clusters, clouds and data for scientific computing. Parallel Processing Letters, 25(3):1502002, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

DuBois:2005:TMS

[DTL05] André Rauber Du Bois, Phil Trinder, and Hans-Wolfgang Loidl. Towards mobility skeletons. Parallel Processing Letters, 15(3):273–288, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Das:1993:DFD

[DTLA93] A. Das, K. Thulasiraman, K. B.

Lakshmanan, and V. K. Agarwal. Distributed fault diagnosis of a ring of processors. Parallel Processing Letters, 3(2): 195–204, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Dourvas:2015:HAC

[DTST15] Nikolaos Dourvas, Michail-Antisthenis Tsompanas, Georgios Ch. Sirakoulis, and Philippos Tsalides. Hardware acceleration of cellular automata Physarum polycephalum model. Parallel Processing Letters, 25 (1):1540006, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Duato:1992:CCN

[Dua92] J. Duato. Channel classes: a new concept for deadlock avoidance in wormhole networks. Parallel Processing Letters, 2(4): 347–354, December 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Duato:1993:DDA

[Dua93]  J. Duato.  On the design

of deadlock-free adaptive multicast routing algorithms. Parallel Processing Letters, 3(4):321– 333, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Duato:1994:TIE

[Dua94] Jose Duato. A theory to increase the effective redundancy in wormhole networks. Parallel Processing Letters, 4(1-2): 125–138, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Darte:1995:RDK

[DV95] A. Darte and F. Vivien. Revisiting the decomposition of Karp, Miller and Winograd. Parallel Processing Letters, 5(4):551– 562, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Darte:1997:PNL

[DV97] Alain Darte and Frédéric Vivien. Parallelizing nested loops with approximation of distance vectors: a survey. Parallel Processing Letters, 7(2):133–144, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Dobrev:2002:OBT

[DV02] S. Dobrev and I. Vrto. Optimal broadcasting in tori with dynamic faults. Parallel Processing Letters, 12(1):17–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Dinneen:1999:CTR

[DVWZ99] Michael J. Dinneen, Jose A. Ventura, Mark C. Wilson, and Golbon Zakeri. Construction of time-relaxed minimal broadcast networks. Parallel Processing Letters, 9(1):53–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Drzadzewski:2007:CMN

[DW07a] Grzegorz Drzadzewski and Mark Wineberg. Comparing minimum neighborhood evaluation schemes for finding spatially robust solutions. Parallel Processing Letters, 17(3):299–309, September 2007. CODEN PPLTEE. ISSN 0129-6264.

Drzadzewski:2007:HSD

[DW07b] Grzegorz Drzadzewski and Mark Wineberg. How solution density affects the finding of spatially robust solutions. Parallel Processing Letters, 17(3): 311–321, September 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Deng:2006:OCS

[DWH^{+}06] Yuhui Deng, Frank Wang, Na Helian, Dan Feng, and Ke Zhou. Optimal clustering size of small file access in network attached storage device. Parallel Processing Letters, 16 (4):501–512, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Ding:2015:SSA

[DWS15] Yihua Ding, James Z. Wang, and Pradip K. Srimani. Selfstabilizing algorithm for minimal dominating set with safe convergence in an arbitrary graph. Parallel Processing Letters, 25(4):1550011, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dehne:2012:DSS

[DZ12] Frank Dehne and Hamidreza Zaboli. Deterministic sample sort for GPUs. Parallel Processing Letters, 22(3): 1250008, September 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dehne:2013:PCD

[DZ13] Frank Dehne and Hamidreza Zaboli. Parallel construction of data cubes on multi-core multidisk platforms. Parallel Processing Letters, 23(1):1350002, March 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Elhadef:2006:PPS

[EADN06] Mourad Elhadef, Kaouther Abrougui, Shantanu Das, and Amiya Nayak. A parallel probabilistic system-level fault diagnosis approach for large multiprocessor systems. Parallel Processing Letters, 16(1):63–79, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Evripidou:1998:MFP

[EB98] Paraskevas Evripidou and Robert Barry. Mapping Fortran programs to single assignment semantics for efficient parallelization. Parallel Processing Letters, 8(3):407–??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Eilam:1998:CCP

[EFZ98] Tamar Eilam, Michele Flammini, and Shmuel Zaks. A complete characterization of the path layout construction problem for ATM networks with given hop count and load. Parallel Processing Letters, 8(2): 207–220, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Ellis:1998:EGG

[EM98] John Ellis and Minko Markov.

Embedding Grids into Grids: Dilation four suffices. Parallel Processing Letters, 8(2): 243–250, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Elsasser:2004:ODS

[EMPF04] Robert Elsässer, Burkhard Monien, Robert Preis, and Andreas Frommer. Optimal diffusion schemes and load balancing on product graphs. Parallel Processing Letters, 14(1):61–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Elgindy:1999:SSL

[ER99] Hossam Elgindy and Sanguthevar Rajasekaran. Sorting and selection on a linear array with optical bus system. Parallel Processing Letters, 9(3):373– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Elhage:1992:SGC

[ES92] H. Elhage and I. Stojmenovic.

Systolic generation of combinations from arbitrary elements. Parallel Processing Letters, 2 (2-3):241–248, September 1992. CODEN PPLTEE. ISSN 0129-

Evripidou:2001:NCW

[ESSP01] P. Evripidou, G. Samaras, C. Spyrou, and A. Papagapiou. Net-console: Web-based development environment for parallel programs. Parallel Processing Letters, 11(1):139–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Emmart:2010:HPI

[EW10] Niall Emmart and Charles Weems. High precision integer addition, subtraction and multiplication with a graphics processing unit. Parallel Processing Letters, 20(4):293–306, December 2010. CODEN PPLTEE. ISSN 0129-6264. See later improvements [EW11].

Emmart:2011:HPI

[EW11] Niall Emmart and Charles C. Weems. High precision integer multiplication with a GPU using Strassen’s algorithm with multiple FFT sizes. Parallel Processing Letters, 21(3):359– 375, September 2011. CODEN PPLTEE. ISSN 0129-6264. See earlier work [EW10].

Emmart:2013:SBA

[EW13] Niall Emmart and Charles Weems. Search-based automatic code generation for multiprecision modular exponentiation on multiple generations of GPU. Parallel Processing Letters, 23 (4):1340009, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Eshaghian-Wilner:2004:SRM

[EWM04] Mary M. Eshaghian-Wilner and Russ Miller. The systolic reconfigurable mesh. Parallel Processing Letters, 14(3/4):337–??, September/December 2004. CO-DEN PPLTEE. ISSN 0129-6264.

Futamura:2003:PSA

[FAH03] Natsuhiko Futamura, Srinivas Aluru, and Xiaoqiu Huang. Parallel syntenic alignments. Parallel Processing Letters, 13(4): 689–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Fragopoulou:2012:GEN

[FAMP12] Paraskevi Fragopoulou, Artur Andrzejak, Carlo Mastroianni, and Costas Panagiotakis. Guest editors’ note: Special issue on community structures in networks: Methods and applications. Parallel Processing Letters, 22(1):1202001, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Farley:1998:ORN

[Far98] Arthur M. Farley. Optimally routable networks. Parallel Processing Letters, 8(2):567– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Floros:2005:TGS

[FC05] Evangelos Floros and Yiannis Cotronis. Towards a Grid services based framework for the virtualization, execution and composition of MPI applications. Parallel Processing Letters, 15(1/2):85–98, March/ June 2005. CODEN PPLTEE. ISSN 0129-6264.

Flatebo:1992:DDD

[FD92] M. Flatebo and A. K. Datta. Distributed deadlock detection algorithms. Parallel Processing Letters, 2(1):21–30, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Fon-Der-Flaass:1993:CIG

[FDFR93] D. Fon-Der-Flaass and I. Rival. Collecting information in graded ordered sets. Parallel Processing Letters, 3(3):253–260, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Fraguela:1999:MHP

[FDZ99] Basilio B. Fraguela, Ramón Doallo, and Emilio L. Zapata. Memory hierarchy performance prediction for blocked sparse algorithms. Parallel Processing Letters, 9(3):347–??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Feautrier:1994:TAD

[Fea94] Paul Feautrier. Toward automatic distribution. Parallel Processing Letters, 4(3):233– 244, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Fiol:1993:UAD

[FFSY93] M. A. Fiol, J. Fabrega, O. Serra, and J. L. A. Yebra. A unified approach to the design and control of dynamic memory networks. Parallel Processing Letters, 3(4): 445–456, December 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Fischer:2002:TUR

[FG02] J. Fischer and S. Gorlatch. Turing universality of recursive patterns for parallel programming. Parallel Processing Letters, 12 (2):229–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Fratto:2015:BCR

[FGK15] Brian E. Fratto, Nataliia Guz, and Evgeny Katz. Biomolecular computing realized in parallel flow systems: Enzyme-based double Feynman logic gate. Parallel Processing Letters, 25(1): 1540001, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Fiorini:1993:XOP

[Fio93] P. Fiorini. The XOmega and the Omega pyramids, a new family of competitive interconnection networks. Parallel Processing Letters, 3(4):495–506, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Fernandes:1996:RIN

[FK96] R. Fernandes and A. Kanevsky.

On recursive interconnection networks and their extensions. Parallel Processing Letters, 6(3): 377–388, September 1996. CO-DEN PPLTEE. ISSN 0129-6264.

Ferreira:1999:ORQ

[FKB^{+}99] Renato Ferreira, Tahsin Kurc, Michael Beynon, Chialin Chang, Alan Sussman, and Joel Saltz. Object-relational queries into multidimensional databases with the active data repository. Parallel Processing Letters, 9(2): 173–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Ferragina:1996:TTP

[FL96] P. Ferragina and F. Luccio. Three techniques for parallel maintenance of a minimum spanning tree under batch of updates. Parallel Processing Letters, 6(2):213–222, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Flammini:1997:HDI

[Fla97] Michele Flammini. On the hardness of devising interval routing schemes. Parallel Processing Letters, 7(1):39–47, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Friedman:2005:BCA

[FMR05] Roy Friedman, Achour Mostefaoui, and Michel Raynal. ♢P_{mute}-based consensus for asynchronous Byzantine systems. Parallel Processing Letters, 15(1/2):169–182, March/ June 2005. CODEN PPLTEE. ISSN 0129-6264.

Fatourou:2017:LWF

[FNP17] Panagiota Fatourou, Yiannis Nikolakopoulos, and Marina Papatriantafilou. Linearizable wait-free iteration operations in shared double-ended queues. Parallel Processing Letters, 27 (2):1750001, June 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Filip:2015:BEA

[FOCK15] Alexandra Vintila Filip, Ana-Maria Oprescu, Stefania Costache, and Thilo Kielmann. E-BaTS: Energy-aware scheduling for bag-of-task applications in HPC clusters. Parallel Processing Letters, 25(3):1541005, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Fox:2013:RSV

[Fox13] Geoffrey Fox. Robust scalable visualized clustering in vector and non vector semi-metric spaces. Parallel Processing Letters, 23(2):1340006, June 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Farley:1993:SN

[FP93] A. M. Farley and A. Proskurowski.

Self-repairing networks. Parallel Processing Letters, 3(4):381– 391, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Fujita:1998:NGH

[FPP98] Satoshi Fujita, Stephane Perennes, and Joseph G. Peters. Neighbourhood gossiping in hypercubes. Parallel Processing Letters, 8(2):189–195, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Furer:1996:PEC

[FR96] M. Furer and B. Raghavachari.

Parallel edge coloring approximation. Parallel Processing Letters, 6(3):321–329, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Fragopoulou:1993:CPC

[Fra93a] P. Fragopoulou. On the comparative powers of the 2D-PARBS and the CRCW-PRAM models. Parallel Processing Letters, 3(3): 301–304, September 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Fragopoulou:1993:ESN

[Fra93b] P. Fragopoulou. On the efficient summation of N numbers on an N-processor reconfigurable mesh. Parallel Processing Letters, 3(1):71–78, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Ferreira:1999:RBC

[FS99] Afonso Ferreira and Nicolas Schabanel. A randomized BSP/ CGM algorithm for the maximal independent set problem. Parallel Processing Letters, 9(3):411– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Feldmann:1992:CCN

[FU92] R. Feldmann and W. Unger. The cube-connected cycles network is a subgraph of the Butterfly network. Parallel Processing Letters, 2(1):13–19, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Fujimoto:2008:DMV

[Fuj08]  Noriyuki Fujimoto.  Dense

matrix-vector multiplication on the CUDA architecture. Parallel Processing Letters, 18(4): 511–530, December 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Feil:1999:CCW

[FUV99] M. Feil, A. Uhl, and M. Vajtersic. Computation of the continuous wavelet transform on massively parallel SIMD arrays. Parallel Processing Letters, 9(4):453–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Fraigniaud:1999:CHO

[FV99] Pierre Fraigniaud and Sandrine Vial. Comparison of heuristics for one-to-all and all-to-all communications in partial meshes. Parallel Processing Letters, 9(1): 9–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Flatz:2016:PNM

[FV16] Markus Flatz and Marián Vajteršic. Parallel nonnegative matrix factorization via Newton iteration. Parallel Processing Letters, 26(3):1650014, September 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Fernandez:2020:SPC

[FVR20] Carlos Fernandez, Ioannis Vourkas, and Antonio Rubio. Shortest path computing in directed graphs with weighted edges mapped on random networks of memristors. Parallel Processing Letters, 30(01):??, March 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420500024.

Fujita:1993:OGM

[FY93] S. Fujita and M. Yamashita. Optimal gossiping in mesh-bus computers. Parallel Processing Letters, 3(4):357–361, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Fang:2016:EMS

[FZL^{+}16] Jianbin Fang, Peng Zhang, Zhaokui Li, Tao Tang, Xuhao Chen, Cheng Chen, and Canqun Yang. Evaluating multiple streams on heterogeneous platforms. Parallel Processing Letters, 26(4):1640002, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Galley:1996:OLN

[Gal96a] C. N. Galley. An Ω(log n − k log k) time linear cost lower bound for the k functions coarsest partition problem. Parallel Processing Letters, 6(2): 195–202, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Garcia:1996:UIP

[GAL96b] Jordi Garcia, Eduard Ayguadé, and Jesus Labarta. Using a 0-1 integer programming model for automatic static data distribution. Parallel Processing Letters, 6(1):159–171, March 1996. CO-DEN PPLTEE. ISSN 0129-6264.

Gamess:2003:ESP

[Gam03] Eric Gamess. Execution of sequential and parallel Java bytecode in a metacomputing system. Parallel Processing Letters, 13(1):53–64, March 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Gamess:2007:AAB

[Gam07] Eric Gamess. An architectural approach to building grids using legacy code. Parallel Processing Letters, 17(4):363–378, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Garzon:2014:DCD

[Gar14] Max H. Garzon. DNA codeword design: Theory and applications. Parallel Processing Letters, 24(2):1440001, June 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gava:2003:FPF

[Gav03] Frédéric Gava. Formal proofs of functional BSP programs. Parallel Processing Letters, 13(3): 365–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Gava:2008:MID

[Gav08] Frédéric Gava. A modular implementation of data structures in bulk-synchronous parallel ML. Parallel Processing Letters, 18(1):39–53, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Gorlatch:1998:GMI

[GB98] Sergei Gorlatch and Holger Bischof. A generic MPI implementation for a data-parallel skeleton: Formal derivation and application to FFT. Parallel Processing Letters, 8(4):447– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Govindarajulu:2012:MMH

[GB12] Naveen Sundar Govindarajulu and Selmer Bringsjord. The myth of ‘the myth of hypercomputation’. Parallel Processing Letters, 22(3):1240012, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Graham:2007:OMH

[GBS^{+}07] Richard L. Graham, Brian W. Barrett, Galen M. Shipman, Timothy S. Woodall, and George Bosilca. Open MPI: A high performance, flexible implementation of MPI point-to-point communications. Parallel Processing Letters, 17(1):79–88, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Germain:1995:SNH

[GC95] C. Germain and F. Cappello. The static network: a high performance reconfigurable communication network. Parallel Processing Letters, 5(1):97–109, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Grosu:2006:TLB

[GC06] Daniel Grosu and Anthony T. Chronopoulos. A truthful load balancing mechanism with verification. Parallel Processing Letters, 16(1):3–17, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Gonzalez:1999:PIL

[GCP99] P. Gonzalez, J. C. Cabaleiro, and T. F. Pena. Parallel incomplete LU factorization as a preconditioner for Krylov subspace methods. Parallel Processing Letters, 9(4):467–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Germain:1994:SEM

[GDC94] C. Germain, F. Delaplace, and R. Carlier. A static execution model for data parallelism. Parallel Processing Letters, 4(4): 367–378, December 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Ghose:2017:FOT

[GDDM17] Anirban Ghose, Lokesh Dokara, Soumyajit Dey, and Pabitra Mitra. A framework for OpenCL task scheduling on heterogeneous multicores. Parallel Processing Letters, 27(3– 4):1750008, 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grant-Duff:1996:PH

[GDH96] Z. N. Grant-Duff and P. G. Harrison. Parallelism via homomorphisms. Parallel Processing Letters, 6(2):279–295, June 1996. CODEN PPLTEE. ISSN 0129-

Griebler:2017:SDH

[GDTF17] Dalvan Griebler, Marco Danelutto, Massimo Torquati, and Luiz Gustavo Fernandes. SPar: A DSL for high-level and productive stream parallelism. Parallel Processing Letters, 27(1): 1740005, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Garces-Erice:2003:HPP

[GEBR^{+}03] L. Garcés-Erice, E. W. Biersack, K. W. Ross, P. A. Felber, and G. Urvoy-Keller. Hierarchical peer-to-peer systems. Parallel Processing Letters, 13(4): 643–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Gelas:2003:TDH

[GEL03] Jean-Patrick Gelas, Saad El Hadri, and Laurent Lefèvre. Towards the design of a high performance active node. Parallel Processing Letters, 13(2): 149–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Geoffray:2001:OPI

[Geo01]  P. Geoffray.  OPIOM: Off-

processor I/O with Myrinet. Parallel Processing Letters, 11 (2–3):237–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Gerlach:2002:GPP

[Ger02] J. Gerlach. Generic programming of parallel applications with Janus. Parallel Processing Letters, 12(2):175–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Gerbessiotis:2018:SIS

[Ger18] Alexandros V. Gerbessiotis. A study of integer sorting on multicores. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500147.

Gupta:1996:AES

[GG96] A. K. Gupta and G. W. Greenwood. Applications of evolutionary strategies to finegrained task scheduling. Parallel Processing Letters, 6(4):551– 561, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Gairing:2004:DTI

[GGH^{+}04] Martin Gairing, Wayne Goddard, Stephen T. Hedetniemi, Petter Kristiansen, and Alice A. Mcrae. Distance-two information in self-stabilizing algorithms. Parallel Processing Letters, 14(3/4):387–??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264.

Gairing:2004:SSM

[GGHJ04] Martin Gairing, Wayne Goddard, Stephen T. Hedetniemi, and David P. Jacobs. Selfstabilizing maximal k-dependent sets in linear time. Parallel Processing Letters, 14(1):75–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Grosser:2012:PPP

[GGL12] Tobias Grosser, Armin Groesslinger, and Christian Lengauer. Polly — performing polyhedral optimizations on a low-level intermediate representation. Parallel Processing Letters, 22(4): 1250010, December 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Georgakoudis:2015:IQS

[GGS^{+}15] Giorgis Georgakoudis, Charles Gillan, Ahmed Sayed, Ivor Spence, Richard Faloon, and Dimitrios S. Nikolopoulos. Isoquality of service: Fairly ranking servers for real-time data analytics. Parallel Processing Letters, 25(3):1541004, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Garba:2012:APU

[GGV12] Michael T. Garba and Horacio González-Vélez. Asymptotic peak utilisation in heterogeneous parallel CPU/GPU pipelines: a decentralised queue monitoring strategy. Parallel Processing Letters, 22(2): 1240008, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gramoli:2010:IAT

[GHF10] Vincent Gramoli, Derin Harmanci, and Pascal Felber. On the input acceptance of transactional memory. Parallel Processing Letters, 20(1):31–50, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Goddard:2008:SSG

[GHJ^{+}08] Wayne Goddard, Stephen T.

Hedetniemi, David P. Jacobs, Pradip K. Srimani, and Zhenyu Xu. Self-stabilizing graph protocols. Parallel Processing Letters, 18(1):189–199, March 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Greenberg:1997:PAS

[GHS97] Ronald I. Greenberg, Shih-Chuan Hung, and Jau-Der Shih. Parallel algorithms for singlelayer channel routing. Parallel Processing Letters, 7(3):267– ??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Gupta:1994:IFF

[GHSJ94] S. K. S. Gupta, C.-H. Huang, P. Sadayappan, and R. W. Johnson. Implementing fast Fourier transforms on distributed-memory multiprocessors using data redistributions. Parallel Processing Letters, 4(4):477–488, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Galley:1994:SPA

[GI94] Clive N. Galley and Costas S.

Iliopoulos. A simple parallel algorithm for the single function coarsest partition problem. Parallel Processing Letters, 4(4): 437–445, December 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Gustedt:2009:EML

[GJQ09] Jens Gustedt, Emmanuel Jeannot, and Martin Quinson. Experimental methodologies for large-scale systems: a survey.

Parallel Processing Letters, 19 (3):399–418, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Gupta:2003:PAV

[GK03] Arvind Gupta and Ramesh Krishnamurti. Parallel algorithms for vehicle routing problems. Parallel Processing Letters, 13 (4):673–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Goraya:2013:FTT

[GK13] Major Singh Goraya and Lakhwinder Kaur. Fault tolerance task execution through cooperative computing in grid. Parallel Processing Letters, 23(1):1350003, March 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grosslinger:2014:GEN

[GK14] Armin Größlinger and Harald Köstler. Guest Editors’ note: Special issue on highperformance stencil computations. Parallel Processing Letters, 24(3):1402002, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gorlatch:2017:GEN

[GK17] Sergei Gorlatch and Herbert Kuchen. Guest Editors’ note. Parallel Processing Letters, 27 (1):1702001, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gioiosa:2014:OMS

[GKK14] Roberto Gioiosa, Gokcen Kestor, and Darren J. Kerbyson. Online monitoring systems for performance fault detection. Parallel Processing Letters, 24(4): 1442003, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Galdi:2005:SPN

[GKMP05] Clemente Galdi, Christos Kaklamanis, Manuela Montangero, and Giuseppe Persiano. Station placement in networks. Parallel Processing Letters, 15(1/2):117– 129, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Grelck:2008:HSM

[GKS08] Clemens Grelck, Steffen Kuthe, and Sven-Bodo Scholz. A hybrid shared memory execution model for a data parallel language with I/O. Parallel Processing Letters, 18(1):23–37, March 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Grebhahn:2014:EOP

[GKS^{+}14] Alexander Grebhahn, Sebastian Kuckuk, Christian Schmitt, Harald Köstler, Norbert Siegmund, Sven Apel, Frank Hannig, and Jürgen Teich. Experiments on optimizing the performance of stencil codes with SPL conqueror. Parallel Processing Letters, 24(3):1441001, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Geist:1999:HAV

[GKSP99] G. A. Geist II, James Arthur Kohl, Stephen Scott, and Philip M. Papadopoulos. Harness: Adaptable virtual machine environment for heterogeneous cluster. Parallel Processing Letters, 9(2):253–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Griebl:1994:SMW

[GL94] Martin Griebl and Christian Lengauer. On the space-time mapping of WHILE-loops. Parallel Processing Letters, 4(3): 221–232, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Gorlatch:2002:F

[GL02] S. Gorlatch and C. Lengauer. Foreword. Parallel Processing Letters, 12(2):137–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Gava:2005:FLD

[GL05] Frédéric Gava and Frédéric Loulergue. A functional language for departmental metacomputing. Parallel Processing Letters, 15(3):289–304, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Glasa:1993:BSA

[Gla93] J. Glasa. Bit-level systolic arrays for digital contour smoothing by Abel-Poisson kernel. Parallel Processing Letters, 3(1):43–51, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Geijer:2004:GCA

[GLMM^{+}04] Johan Geijer, Boris Lenhard, Roxana Merino-Martinez, Gunnar Norstedt, and Amilcar Flores-Morales. Grid computing for the analysis of regulatory elements in co-regulated sets of genes. Parallel Processing Letters, 14(2):137–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Gairing:2006:PAR

[GLMM06] Martin Gairing, Thomas Lücking, Marios Mavronicolas, and Burkhard Monien. The price of anarchy for restricted parallel links. Parallel Processing Letters, 16(1): 117–131, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Galbiati:1992:CEP

[GM92] G. Galbiati and F. Maffioli. Constructing an exact parity base is in RCN^{2}. Parallel Processing Letters, 2(4):301–309, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Gallard:2003:DSE

[GM03] Pascal Gallard and Christine Morin. Dynamic streams for efficient communications between migrating processes in a cluster. Parallel Processing Letters, 13 (4):601–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Green:2004:CSP

[GM04] Mark L. Green and Russ Miller.

A client-server prototype for grid-enabling application template design. Parallel Processing Letters, 14(2):241–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Gusatto:2005:EPA

[GMCC05] Éder Gusatto, José C. M. Mombach, Fernando P. Cercato, and Gerson H. Cavalheiro. An efficient parallel algorithm to evolve simulations of the cellular Potts model. Parallel Processing Letters, 15(1/2):199–208, March/ June 2005. CODEN PPLTEE. ISSN 0129-6264.

Grimshaw:2013:GXG

[GMK13] Andrew Grimshaw, Mark Morgan, and Avinash Kalyanaraman. GFFS — the XSEDE Global Federated File System. Parallel Processing Letters, 23 (2):1340005, June 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gastaldo:1992:TCP

[GMR92] M. Gastaldo, M. Morvan, and J. M. Robson. Transitive closure in parallel on a linear network of processors. Parallel Processing Letters, 2(2-3):195–203, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Gonzalez:2009:ICS

[Gon09] Teofilo F. Gonzalez. Improved communication schedules with buffers. Parallel Processing Letters, 19(1):129–139, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Gu:1996:FTR

[GP96] Qian-Ping Gu and Shietung Peng. Fault tolerant routing in hypercubes and star graphs. Parallel Processing Letters, 6(1): 127–136, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Gascard:2003:FPA

[GP03] Eric Gascard and Laurence Pierre. Formal proof of applications distributed in symmetric interconnection networks. Parallel Processing Letters, 13(1): 3–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Georgiou:2009:SRP

[GPP09] Chryssis Georgiou, Theophanis Pavlides, and Anna Philippou. Selfish routing in the presence of network uncertainty. Parallel Processing Letters, 19(1): 141–157, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Gontmakher:2003:CVJ

[GPS03] Alex Gontmakher, Sergey Polyakov, and Assaf Schuster. Complexity of verifying Java shared memory execution. Parallel Processing Letters, 13(4):721–734, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Garg:1999:NCC

[GR99] Vijay K. Garg and Michel Raynal. Normality: A consistency condition for concurrent objects. Parallel Processing Letters, 9(1): 123–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Gamboa:2011:SPB

[GR11] Carlos Fernando Gamboa and Thomas Robertazzi. Simple performance bounds for multicore and parallel channel systems. Parallel Processing Letters, 21(4):439–460, December 2011. CODEN PPLTEE. ISSN 0129-6264.

Greenlaw:1993:BSP

[Gre93] R. Greenlaw. Breadth-depth search is P-complete. Parallel Processing Letters, 3(3):209– 222, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Griebel:1992:CTS

[Gri92] M. Griebel. The combination technique for the sparse grid solution of PDEs on multiprocessor machines. Parallel Processing Letters, 2(1):61–70, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Georgiou:2007:FSA

[GRS07] Chryssis Georgiou, Alexander Russell, and Alexander A. Shvartsman. Failure-sensitive analysis of parallel algorithms with controlled memory access concurrency. Parallel Processing Letters, 17(2):153–168, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Gasperoni:1994:GCO

[GS94a] Franco Gasperoni and Uwe Schwiegelshohn. Generating close to optimum loop schedules on parallel processors. Parallel Processing Letters, 4(4):391– 403, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Gewali:1994:CEW

[GS94b] Laxmi P. Gewali and Ivan Stojmenovic. Computing external watchman routes on PRAM, BSR, and interconnection network models of parallel computation. Parallel Processing Letters, 4(1-2):83–93, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Gupta:1995:TGP

[GS95] V. Gupta and E. Schenfeld. Task graph partitioning and mapping in a reconfigurable parallel architecture. Parallel Processing Letters, 5(4):563–574, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Gupta:1997:SAR

[GS97] Manish Gupta and Edith Schonberg. Static analysis to reduce synchronization costs dataparallel programs with remote memory copy. Parallel Processing Letters, 7(2):145–156, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Gerbessiotis:1999:EDS

[GS99] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Efficient deterministic sorting on the BSP model. Parallel Processing Letters, 9(1):69–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Grelck:2003:SHL

[GS03] Clemens Grelck and Sven-Bodo Scholz. SAC — from high-level programming with arrays to efficient parallel execution. Parallel Processing Letters, 13(3): 401–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Gopalan:2015:MDS

[GS15] N. P. Gopalan and S. Suresh. Modified delay scheduling: A heuristic approach for Hadoop scheduling to improve fairness and response time. Parallel Processing Letters, 25(4): 1550009, December 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grelck:2008:GIN

[GSS08] Clemens Grelck, Sven-Bodo Scholz, and Alex Shafarenko. A gentle introduction to S-Net: Typed stream processing and declarative coordination of asynchronous components. Parallel Processing Letters, 18(2): 221–237, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Gruau:2000:CG

[GT00a] F. Gruau and J. Tromp. Cellular gravity. Parallel Processing Letters, 10(4):383–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000354.
html.

Guyennet:2000:LBU

[GT00b] H. Guyennet and M. Trehel. Load balancing using processor groups. Parallel Processing Letters, 10(1):59–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000081.
html.

Guo:2001:DSH

[Guo01] M. Guo. Denotational semantics of an HPF-like data-parallel language model. Parallel Processing Letters, 11(2–3):363–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Gastin:1993:ECS

[GV93] P. Gastin and V. Villain. An efficient crash-tolerant sequential traversal. Parallel Processing Letters, 3(1):87–97, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Grayson:1996:HPP

[GV96] Brian Grayson and Robert Van De Geijn. A high performance parallel Strassen implementation. Parallel Processing Letters, 6(1):3–12, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Gautama:2003:SPE

[GV03] H. Gautama and A. J. C. Van Gemund. Symbolic performance estimation of speculative parallel programs. Parallel Processing Letters, 13(4):513–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Grosser:2014:RBD

[GVCS14] Tobias Grosser, Sven Verdoolaege, Albert Cohen, and P. Sadayappan. The relation between diamond tiling and hexagonal tiling. Parallel Processing Letters, 24(3):1441002, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Ganapathy:1994:OSP

[GW94] Kumar Ganapathy and Benjamin W. Wah. Optimal synthesis of processor arrays with pipelined arithmetic units. Parallel Processing Letters, 4(3): 339–350, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Gu:2020:SDS

[GZL^{+}20] Zhendong Gu, Shuming Zhou, Jiafei Liu, Qianru Zhou, and Dajin Wang. Shapley distance and Shapley index for some special graphs. Parallel Processing Letters, 30(04):??, December 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500127.

Gao:2016:NMG

[GZW16] Jiaquan Gao, Yuanshen Zhou, and Kesong Wu. A novel multi-GPU parallel optimization model for the sparse matrixvector multiplication. Parallel Processing Letters, 26(4): 1640001, December 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Hadjidoukas:2010:NOP

[HA10] Panagiotis E. Hadjidoukas and Laurent Amsaleg. Nested OpenMP parallelization of a hierarchical data clustering algorithm. Parallel Processing Letters, 20(2):187–208, June 2010.

CODEN PPLTEE. ISSN 0129-6264.

Haddix:2008:ODA

[Had08] F. Furman Haddix. An order degree alternator for arbitrary topologies. Parallel Processing Letters, 18(2):307–322, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Haglin:1992:AMS

[Hag92] D. J. Haglin. Approximating maximum 2-CNF satisfiability. Parallel Processing Letters, 2 (2-3):181–187, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Haglin:1995:BEM

[Hag95] D. J. Haglin. Bipartite expander matching is in NC. Parallel Processing Letters, 5(3):413– 420, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Hammond:2003:AST

[HBL03] Kevin Hammond, Jost Berthold, and Rita Loogen. Automatic skeletons in Template Haskell. Parallel Processing Letters, 13 (3):413–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Hayashi:2002:ACA

[HC02] Y. Hayashi and M. Cole. Automated cost analysis of a parallel maximum segment sum program derivation. Parallel Processing Letters, 12(1):95–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Hassan:2019:EAM

[HCD^{+}19] Ambra Abdullahi Hassan, Valeria Cardellini, Pasqua D’ambra, Daniela Di Serafino, and Salvatore Filippone. Efficient algebraic multigrid preconditioners on clusters of GPUs. Parallel Processing Letters, 29(01):??, March 2019. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500014.

Head:2007:PET

[Hea07] Tom Head. Photocomputing: Explorations with transparency and opacity. Parallel Processing Letters, 17(4):339–347, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Herrmann:2005:GMP

[Her05] Christoph A. Herrmann. Generating message-passing programs from abstract specifications by partial evaluation. Parallel Processing Letters, 15(3):305– 320, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Hains:2012:GEN

[HGH12] Gaétan Hains, Frédéric Gava, and Kevin Hammond. Guest Editors’ note: High-level programming for heterogeneous and hierarchical parallel systems. Parallel Processing Letters, 22 (2):1202002, June 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Hsieh:1999:EPA

[HHH^{+}99] Sun-Yuan Hsieh, Chin-Wen Ho, Tsan-Sheng Hsu, Ming-Tat Ko, and Gen-Huey Chen. Efficient parallel algorithms on distance hereditary graphs. Parallel Processing Letters, 9(1): 43–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Hedetniemi:2013:SSA

[HHKM13] Sandra M. Hedetniemi, Stephen T.

Hedetniemi, K. E. Kennedy, and Alice A. Mcrae. Self-stabilizing algorithms for unfriendly partitions into two disjoint dominating sets. Parallel Processing Letters, 23(1):1350001, March 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Heller:2007:LCL

[HHL^{+}07] Steve Heller, Maurice Herlihy, Victor Luchangco, Mark Moir, William N. Scherer III, and Nir Shavit. A lazy concurrent listbased set algorithm. Parallel Processing Letters, 17(4):411– 424, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Hidalgo-Herrero:2002:OSP

[HHOM02] M. Hidalgo-Herrero and Y. Ortega-Mallén. An operational semantics for the parallel language Eden. Parallel Processing Letters, 12(2):211–??, June 2002. CODEN PPLTEE. ISSN 0129-

Hinsen:2003:HLP

[Hin03] Konrad Hinsen. High-level parallel software development with Python and BSP. Parallel Processing Letters, 13(3):473– ??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Hagerup:1995:FPP

[HK95a] Torben Hagerup and Jörg Keller. Fast parallel permutation algorithms. Parallel Processing Letters, 5(2):139–148, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Ho:1995:EBH

[HK95b] Ching-Tien Ho and Ming-Yang Kao. Efficient broadcast on hypercubes with wormhole and ecube routings. Parallel Processing Letters, 5(2):213–222, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Heinzlreiter:2003:VSG

[HK03] Paul Heinzlreiter and Dieter Kranzlmüller. Visualization services on the Grid: the Grid Visualization Kernel. Parallel Processing Letters, 13(2): 135–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Higham:2016:PMC

[HK16] Lisa Higham and Jalal Kawash. Pitfalls in memory consistency modelling. Parallel Processing Letters, 26(2):1650008, June 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Hromkovic:1998:ESA

[HKP^{+}98] Juraj Hromkovič, Ralf Klasing, Dana Pardubská, Juraj Waczulík, and Hubert Wagener. Effective systolic algorithms for gossiping in cycles. Parallel Processing Letters, 8(2):197–205, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Habbas:2001:SMI

[HKS01] Z. Habbas, M. Krajecki, and D. Singer. Shared memory implementation of constraint satisfaction problem resolution. Parallel Processing Letters, 11(4): 487–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Herrmann:1996:SMC

[HL96] C. Herrmann and C. Lengauer.

On the space-time mapping of a class of divide-and-conquer recursions. Parallel Processing Letters, 6(4):525–537, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Harutyunyan:1998:MB

[HL98a] Hovhannes A. Harutyunyan and Arthur L. Liestman. Messy broadcasting. Parallel Processing Letters, 8(2):149–159, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Hong:1998:NOB

[HL98b] Tzung-Pei Hong and Jyh-Jong Lee. A nearly optimal backpropagation learning algorithm on a bus-based architecture. Parallel Processing Letters, 8(3): 297–??, September 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Hambrusch:1999:MDT

[HL99] Susanne E. Hambrusch and Hyeong-Seok Lim. Minimizing the diameter in tree networks under edge reductions. Parallel Processing Letters, 9(3):361– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Herrmann:2000:HOL

[HL00] C. A. Herrmann and C. Lengauer. A higher-order language for divide-and-conquer. Parallel Processing Letters, 10 (2/3):239–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1002_03/S0129626400000238.
                                        html.

Herrmann:2002:FFI

[HL02] C. A. Herrmann and C. Lengauer. Functional formalisms and implementations. Parallel Processing Letters, 12(2):193–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Hains:2003:PSI

[HL03] Gaétan Hains and Frédéric Loulergue. Preface: Special issue on high-level parallel programming and applications. Parallel Processing Letters, 13(3):313–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Hung:2017:SFH

[HLH^{+}17] Chun-Nan Hung, Cheng-Kuan Lin, Lih-Hsing Hsu, Eddie Cheng, and László Lipták.

Strong fault-Hamiltonicity for the crossed cube and its extensions. Parallel Processing Letters, 27(2):1750005, June 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Hwang:1997:PRA

[HLJ97] F. K. Hwang, Tzai-Shunne Lin, and Rong-Hong Jan. A permutation routing algorithm for double loop networks. Parallel Processing Letters, 7(3):259– ??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Hamdi-larbi:2021:MLD

[HlMD21] Olfa Hamdi-larbi, Ichrak Mehrez, and Thomas Dufaud. Machine learning to design an auto-tuning system for the best compressed format detection for parallel sparse computations. Parallel Processing Letters, 31(04):??, December 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500195.

Hains:2001:PHL

[HM01] G. Hains and Q. Miller. Preface:

High-level parallel programming and applications. Parallel Processing Letters, 11(4): 375–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Hollingsworth:1999:BVM

[HMA99] Jeffrey K. Hollingsworth, Ethan L.

Miller, and Kennedy Akala. Binary version management for Computational Grids. Parallel Processing Letters, 9(2):215–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Hernandez-orozco:2020:SPL

[HoHqZS20] Santiago Hernández-orozco, Francisco Hernández-quiroz, Hector Zenil, and Wilfried Sieg. Shortening of proof length is elusive for theorem provers. Parallel Processing Letters, 30(04): ??, December 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420500139.

Hsu:1992:EML

[HP92] Wen-Jing Hsu and C. V. Page. Embedding mesh in a large family of graphs. Parallel Processing Letters, 2(2-3):149–155, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Herley:1999:FDP

[HPP99] Kieran T. Herley, Andrea Pietracaprina, and Geppino Pucci. Fast deterministic parallel branch-and-bound. Parallel Processing Letters, 9(3):325– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Haddad:2006:NEP

[HPP06] Serge Haddad and Jean-François Pradat-Peyre. New efficient Petri nets reductions for parallel programs verification. Parallel Processing Letters, 16(1): 101–116, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Hamdi:1998:CPA

[HQP98] M. Hamdi, C. Qiao, and Y. Pan.

On the computing power of arrays of processors with optical pipelined buses. Parallel Processing Letters, 8(4):503– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Hofinger:2018:FAE

[HRH18] Siegfried Höfinger, Thomas Ruh, and Ernst Haunschmid. Fast approximate evaluation of parallel overhead from a minimal set of measured execution times. Parallel Processing Letters, 28(1):1850003, March 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Higham:1993:PMA

[HS93] L. Higham and E. Schenk.

PRAM memory allocation and initialization. Parallel Processing Letters, 3(3):291–299, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Hily:1996:GDM

[HS96] Andre Hily and Dominique Sotteau. Gossiping in d-dimensional mesh-bus networks. Parallel Processing Letters, 6(1):101– 113, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Hirschberg:1997:DCP

[HS97] Daniel S. Hirschberg and Lynn M. Stauffer. Dictionary compression on the PRAM. Parallel Processing Letters, 7(3):

297–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Hoefler:2009:ENN

[HSL09] Torsten Hoefler, Timo Schneider, and Andrew Lumsdaine. The effect of network noise on large-scale collective communications. Parallel Processing Letters, 19(4):573–593, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Hsu:1993:MPS

[Hsu93] Lih-Hsing Hsu. Matching polynomials of series-parallel graphs. Parallel Processing Letters, 3(1): 13–18, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Heckler:1994:CLD

[HT94] Christian Heckler and Lothar Thiele. Computing linear data dependencies in nested loop programs. Parallel Processing Letters, 4(3):193–204, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Hofri:1998:MTM

[HT98] Micha Hofri and David L. Thomson. Matrix transposition on a mesh with blocking transmissions. Parallel Processing Letters, 8(2):561–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Hu:1999:COH

[HT99] Zhenjiang Hu and Masato Takeichi. Calculating an optimal homomorphic algorithm for bracket matching. Parallel Processing Letters, 9(3):335– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Heriban:2021:MRU

[HT21] Adam Heriban and Sébastien Tixeuil. Mobile robots with uncertain visibility sensors: Possibility results and lower bounds. Parallel Processing Letters, 31 (01):??, March 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S012962642150002X.

Hayashi:2005:PCS

[HTHH05] Ryoko Hayashi, Kenji Tanaka, Susumu Horiguchi, and Yasuaki Hiwatari. A parallelization casestudy of MD simulation of a low density physical system. Parallel Processing Letters, 15(4): 481–489, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Huang:2006:MWR

[Hua06] Kuo-Chan Huang. Minimizing waiting ratio for dynamic workload on parallel computers. Parallel Processing Letters, 16(4): 441–453, December 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Hofbauer:2006:QMC

[HUZ06] Heinz Hofbauer, Andreas Uhl, and Peter Zinterhof. Quasi Monte Carlo integration in Grid environments: Further leaping effects. Parallel Processing Letters, 16(3):285–311, September 2006. CODEN PPLTEE. ISSN 0129-6264.

Huang:2006:PAT

[HW06] Kuo-Chan Huang and Feng-Jian Wang. Parallelizing an application with time-increasing workload using adaptive processor allocation. Parallel Processing Letters, 16(4):455–466, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Hamann:2007:EC

[HW07] Heiko Hamann and Heinz Wörn. Embodied computation. Parallel Processing Letters, 17(3): 287–298, September 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Hao:2018:RRC

[HY18] Chen Hao and Weihua Yang. A remark on rainbow 6-cycles in hypercubes. Parallel Processing Letters, 28(02):??, June 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S012962641850007X.

Hao:2019:GCD

[HY19] Chen Hao and Weihua Yang. The generalized connectivity of data center networks. Parallel Processing Letters, 29 (02):??, June 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626419500075.

Hager:2008:DAC

[HZW08] Georg Hager, Thomas Zeiser, and Gerhard Wellein. Data access characteristics and optimizations for Sun UltraSPARC T2 and T2+ systems. Parallel Processing Letters, 18(4): 471–490, December 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Ishii:2005:ISC

[IDS^{+}05] Renato P. Ishii, Rodrigo F. De Mello, Luciano J. Senger, Marcos J. Santana, Regina H. C. Santana, and Laurence Tianruo Yang. Improving scheduling of communication intensive parallel applications on heterogeneous computing environments. Parallel Processing Letters, 15(4):423–438, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Iliopoulos:1996:OPS

[IK96] C. S. Iliopoulos and M. Korda. Optimal parallel superprimitivity testing for square arrays. Parallel Processing Letters, 6(3): 299–308, September 1996. CO-DEN PPLTEE. ISSN 0129-6264.

Ino:2009:HPI

[IKMH09] Fumihiko Ino, Yuki Kotani, Yuma Munekawa, and Kenichi Hagihara. Harnessing the power of idle GPUs for acceleration of biological sequence alignment. Parallel Processing Letters, 19(4):513–533, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Igarashi:1999:SPM

[IN99] Y. Igarashi and Y. Nishitani.

Speedup of the n-process mutual exclusion algorithm. Parallel Processing Letters, 9(4):

475–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

6264.  URL http://www.
wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Izadi:1998:OSA

[IÖ98] Baback A. Izadi and Füsun Özgüner. Optimal subcube allocation in a circuit-switched faulty hypercube. Parallel Processing Letters, 8(1):77–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Ibarra:1995:TOE

[IR95] Louis Ibarra and Dana Richards. Tree open ear decomposition in parallel graph algorithms. Parallel Processing Letters, 5(2): 129–138, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Imbs:2016:TRE

[IR16] Damien Imbs and Michel Raynal. Trading off t-resilience for efficiency in asynchronous Byzantine reliable broadcast. Parallel Processing Letters, 26 (4):1650017, December 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Irony:2002:TRC

[IT02] D. Irony and S. Toledo. Trading replication for communication in parallel distributed-memory dense solvers. Parallel Processing Letters, 12(1):79–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Johnen:2002:OSS

[JADT02] Colette Johnen, Luc O. Alima, Ajoy K. Datta, and Sébastien Tixeuil. Optimal snap-stabilizing neighborhood synchronizer in tree networks. Parallel Processing Letters, 12(3–4):327–??, September– December 2002. CODEN PPLTEE. ISSN 0129-6264.

Jain:1999:PSM

[Jai99] A. Jain. Parallel search and multisearch in matrices with sorted columns. Parallel Processing Letters, 9(4):575–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Johnson:1992:PBM

[JD92] T. Johnson and T. A. Davis.

Parallel buddy memory management. Parallel Processing Letters, 2(4):391–398, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Jesshope:1993:LRV

[Jes93] C. Jesshope. Latency reduction in VLSI routers. Parallel Processing Letters, 3(4):485– 494, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Jesshope:2006:MMD

[Jes06] Chris Jesshope. Microthreading a model for distributed instruction-level concurrency. Parallel Processing Letters, 16 (2):209–228, June 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Jesshope:2008:OSS

[Jes08] Chris Jesshope. Operating systems in silicon and the dynamic management of resources in many-core chips. Parallel Processing Letters, 18(2): 257–274, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Jard:1996:ITD

[JJ96] C. Jard and G.-V. Jourdan. Incremental transitive dependency tracking in distributed computations. Parallel Processing Letters, 6(3):427–435, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Jones:2008:GEN

[JKRW08] Alex K. Jones, Darren J. Kerbyson, Ram Rajamony, and Charles Weems. Guest Editor’s note: Large-scale parallel processing. Parallel Processing Letters, 18(4):449–451, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Jones:2009:GEN

[JKRW09] Alex K. Jones, Darren J. Kerbyson, Ram Rajamony, and Charles Weems. Guest Editor’s note: Large scale parallel processing. Parallel Processing Letters, 19(4):487–490, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Joo:2005:ASM

[JL05] Kil Hong Joo and Won Suk Lee. Adaptive selection of materialized queries in a mediator. Parallel Processing Letters, 15(4): 451–458, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Jones:2015:MIP

[Jon15] Jeff Jones. Mechanisms inducing parallel computation in a model of Physarum polycephalum transport networks. Parallel Processing Letters, 25 (1):1540004, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

JaJa:1996:ORP

[JR96] J. JáJá and Kwan Woo Ryu.

An optimal randomized parallel algorithm for the single function coarsest partition problem. Parallel Processing Letters, 6(2): 187–193, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Jana:2006:IPP

[JS06] Prasanta K. Jana and Bhabani P. Sinha. An improved parallel prefix algorithm on OTIS-Mesh. Parallel Processing Letters, 16(4):429–440, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Jain:2015:SSP

[JS15] Surabhi Jain and N. Sadagopan.

Simpler sequential and parallel biconnectivity augmentation in trees. Parallel Processing Letters, 25(4):1550010, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Jain:2018:HCL

[JS18] Tripti Jain and Klaus Schneider. The half cleaner lemma: Constructing efficient interconnection networks from sorting networks. Parallel Processing Letters, 28(1):1850001, March 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Jeannot:2007:IRT

[JSYD07] Emmanuel Jeannot, Keith Seymour, Asym Yarkhan, and Jack J. Dongarra. Improved runtime and transfer time prediction mechanisms in a network enabled servers middleware. Parallel Processing Letters, 17(1):47–59, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Jones:2008:SEA

[JSZM08] Alex K. Jones, Shuyi Shao, Yu Zhang, and Rami Melhem. Symbolic expression analysis for compiled communication. Parallel Processing Letters, 18(4): 567–587, December 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Jafer:2007:PAC

[JW07] Shafagh Jafer and Gabriel A. Wainer. Parallel algorithms for cellular models simulation. Parallel Processing Letters, 17(3): 263–285, September 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Kelly:2005:GAM

[KB05] Paul H. J. Kelly and Olav Beckmann. Generative and adaptive methods in performance programming. Parallel Processing Letters, 15(3):239–255, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Kerbyson:2011:MPD

[KB11] Darren J. Kerbyson and Kevin J.

Barker. Modeling the performance of direct numerical simulation on parallel systems. Parallel Processing Letters, 21(3): 301–318, September 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Kehrer:2019:EPS

[KB19] Stefan Kehrer and Wolfgang Blochinger. Elastic parallel systems for high performance cloud computing: State-of-the-art and future directions. Parallel Processing Letters, 29 (02):??, June 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500063.

Kelly:2001:TCD

[KBFB01] P. H. J. Kelly, O. Beckmann, T. Field, and S. B. Baden. THEMIS: Component dependence metadata in adaptive parallel applications. Parallel Processing Letters, 11(4):455– ??, December 2001. CODEN PPLTEE. ISSN 0129-6264.

Kerbyson:2013:PAT

[KBG^{+}13] Darren J. Kerbyson, Kevin J.

Barker, Diego S. Gallo, Dong Chen, Jose R. Brunheroto, Kyung Dong Ryu, George L.

Chiu, and Adolfy Hoisie. A performance analysis of three generations of Blue Gene. Parallel Processing Letters, 23(4): 1340007, December 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Khachiyan:2007:CMM

[KBGE07] Leonid Khachiyan, Endre Boros, Vladimir Gurvich, and Khaled Elbassioni. Computing many maximal independent sets for hypergraphs in parallel. Parallel Processing Letters, 17(2): 141–152, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Kosch:2003:ENS

[KBH03] Harald Kosch, Laszlo Böszörményi, and Hermann Hellwagner. Editorial note: “Part I: the Best Papers from Euro-Par 2003”. Parallel Processing Letters, 13(4): 509–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Kuchen:2002:APS

[KC02] H. Kuchen and M. Cole. Application programming with skeletons. Parallel Processing Letters, 12(2):141–??, June 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Kim:1992:RMS

[KCH92] Do-Hyung Kim, Kwang-Moo Choe, and Taisook Han. Refined mark(s)-set-based backtrack literal selection for AND parallelism in logic programs. Parallel Processing Letters, 2(1):71–79, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Kim:2000:WRE

[KCPP00] B. W. Kim, J. H. Choi, K.- I. Park, and K. H. Park. A wormhole router with embedded broadcasting virtual bus for mesh computers. Parallel Processing Letters, 10(1):29–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000056.
html.

Kulkarni:1993:PPA

[KDK^{+}93] R. Kulkarni, S. Date, B. Kulkarni, U. Kulkarni, and A. S. Kolaskar. PRAS: parallel alignment of sequences algorithm. Parallel Processing Letters, 3(3): 243–252, September 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Kim:2000:WDS

[KE00] S. W. Kim and R. Eigenmann. Where does the speedup go: Quantitative modeling of performance losses in sharedmemory programs. Parallel Processing Letters, 10 (2/3):227–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000226.
html.

Kutil:2006:PWF

[KE06] Rade Kutil and Peter Eder. Parallelization of wavelet filters using SIMD extensions. Parallel Processing Letters, 16(3): 335–349, September 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Kumar:1995:CAP

[KESH95] B. Kumar, K. Eswar, P. Sadayappan, and C.-H. Huang. A clustering algorithm for parallel sparse Cholesky factorization. Parallel Processing Letters, 5(4):685–696, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Kyriacou:2006:CCO

[KET06] Costas Kyriacou, Paraskevas Evripidou, and Pedro Trancoso. CacheFlow: Cache optimizations for data driven multithreading. Parallel Processing Letters, 16(2):229–244, June 2006. CODEN PPLTEE. ISSN 0129-6264.

Kee:2002:EIB

[KH02] Y. Kee and S. Ha. An efficient implementation of the BSP programming library for VIA. Parallel Processing Letters, 12(1): 65–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Krings:2004:SIS

[KHAM04] Axel W. Krings, William S. Harrison, Azad Azadmanesh, and Miles Mcqueen. Scheduling issues in survivability applications using hybrid fault models. Parallel Processing Letters, 14(1): 5–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Kerbyson:2005:UPP

[KHW05] Darren J. Kerbyson, Adolfy Hoisie, and Harvey J. Wasserman. Use of predictive performance modeling during largescale system installation. Parallel Processing Letters, 15(4): 387–395, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Kim:1998:EBH

[KKC98] Sook-Yeon Kim, Oh-Heum Kwon, and Kyung-Yong Chwa. Embeddings of butterflies into hypermeshes. Parallel Processing Letters, 8(3):337–??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Kucevalovs:2014:ITL

[KKFZ14] Ilja Kucevalovs, Ojārs Krasts, Rūsiņš Freivalds, and Thomas Zeugmann. On the influence of technology on learning processes. Parallel Processing Letters, 24(2):1440003, June 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kronawitter:2018:ADL

[KKKL18] Stefan Kronawitter, Sebastian Kuckuk, Harald Köstler, and Christian Lengauer. Automatic data layout transformations in the ExaStencils code generator. Parallel Processing Letters, 28(03):??, September 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500093.

Kaklamanis:2000:HAS

[KKS00] C. Kaklamanis, C. Konstantopoulos, and A. Svolos. A hypercube algorithm for sliding window compression. Parallel Processing Letters, 10 (4):315–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1004/S0129626400000305.
                                        html.

Lin:2020:SSC

[kLCL20] Cheng kuan Lin, Eddie Cheng, and László Lipták. Structure and substructure connectivity of hypercube-like networks. Parallel Processing Letters, 30(03):??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420400071.

Koster:2014:CRS

[KLH^{+}14] Marcel Köster, Roland Leißa, Sebastian Hack, Richard Membarth, and Philipp Slusallek. Code refinement of stencil codes. Parallel Processing Letters, 24 (3):1441003, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kerbyson:2008:IRT

[KLJ08] Darren J. Kerbyson, Mike Lang, and Gregory Johnson. Infini-Band routing table optimizations for scientific applications. Parallel Processing Letters, 18 (4):589–608, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Karpinski:2004:WEA

[KLN04] Marek Karpinski, Lawrence Larmore, and Yakov Nekrich. Workefficient algorithms for the construction of length-limited Huffman codes. Parallel Processing Letters, 14(1):99–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Klappenecker:2010:SST

[KLW10] Andreas Klappenecker, Hyunyoung Lee, and Jennifer L. Welch. Scheduling sensors by tiling lattices. Parallel Processing Letters, 20(1):3–13, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Kavi:2002:MEO

[KM02] Krishna M. Kavi and Dinesh P.

Mehta. Mutual exclusion on optical buses. Parallel Processing Letters, 12(3–4):341–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Kutrib:2009:CDI

[KM09] Martin Kutrib and Andreas Malcher. Computations and decidability of iterative arrays with restricted communication. Parallel Processing Letters, 19(2): 247–264, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Kakulavarapu:2001:DLB

[KMAG01] P. Kakulavarapu, O. C. Maquelin, J. N. Amaral, and G. R. Gao. Dynamic load balancers for a multithreaded multiprocessor system. Parallel Processing Letters, 11(1):169–??, March 2001. CODEN PPLTEE. ISSN 0129-

Kmet:2014:NNS

[Kme14] Tibor Kmet. Neural networks solving free and fixed final time optimal control problems. Parallel Processing Letters, 24(2): 1440002, June 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kim:1997:REH

[KMY97] Jin S. Kim, Seung Ryoul Maeng, and H. Yoon. Ring embedding in hypercubes with faculty nodes. Parallel Processing Letters, 7(3): 285–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Krishnamurti:1992:OSA

[KN92] R. Krishnamurti and B. Narahari. Optimal subcube assignment for partitionable hypercubes. Parallel Processing Letters, 2(1):89–95, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Krizanc:1995:ZSM

[KN95] Danny Krizanc and Lata Narayanan. Zero-one sorting on the mesh. Parallel Processing Letters, 5(2): 149–155, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Karpinski:2006:ACO

[KN06] Marek Karpinski and Yakov Nekrich. Algorithms for construction of optimal and almostoptimal length-restricted codes. Parallel Processing Letters, 16 (1):81–92, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Kawai:2014:ADT

[KOKM14] Shinji Kawai, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Analysis of distributed token circulation algorithm with faulty random number generator. Parallel Processing Letters, 24(1):1450002, March 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Koppelman:1996:LBA

[Kop96] David M. Koppelman. A lower bound on the average physical length of edges in the physical realization of graphs. Parallel Processing Letters, 6(1): 137–143, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Kelly:1994:SAM

[KP94] Wayne Kelly and William Pugh.

Selecting affine mappings based on performance estimation. Parallel Processing Letters, 4(3): 205–219, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Kwai:1996:GHN

[KP96] Ding-Ming Kwai and B. Parhami.

A generalization of hypercubic networks based on their chordal ring structures. Parallel Processing Letters, 6(4):469– 477, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Kou:2021:CEC

[KQY21] Shuai Kou, Chengfu Qin, and Weihua Yang. Contractible edges in 3-connected cubic graphs. Parallel Processing Letters, 31(03):??, September 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500146.

Karpinski:1998:STP

[KR98] Marek Karpinski and Wojciech Rytter. On a sublinear time parallel construction of optimal binary search trees. Parallel Processing Letters, 8(3):387– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Kranzlmuller:2001:NAS

[Kra01] D. Kranzlmüller. Nondeterminism analysis on supercomputers and clusters. Parallel Processing Letters, 11(2–3):251–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Kremer:1997:ONO

[Kre97] Ulrich Kremer. Optimal and near–optimal solutions for hard compilation problems. Parallel Processing Letters, 7(4):371– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Krizanc:1991:NOP

[Kri91] D. Krizanc. A note on off-line permutation routing on a meshconnected processor array. Parallel Processing Letters, 1(1):67– 70, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Krizanc:1993:CMS

[KRS93] D. Krizanc, S. Rajasekaran, and S. S. Shende. A comparison of meshes with static buses and half-duplex wrap-arounds. Parallel Processing Letters, 3(2): 109–114, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Kari:2015:WCS

[KRS15] Chadi Kari, Alexander Russell, and Narasimha Shashidhar. Work-competitive scheduling on task dependency graphs. Parallel Processing Letters, 25 (2):1550001, June 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Krumme:1998:ODP

[Kru98] David W. Krumme. An optimal disjoint pair of additive spanners for the 3D Grid. Parallel Processing Letters, 8(2): 251–258, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Krusche:2008:EEB

[Kru08] Peter Krusche. Experimental evaluation of BSP programming libraries. Parallel Processing Letters, 18(1):7–21, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Kaplow:1996:POB

[KS96] Wesley K. Kaplow and Boleslaw K.

Szymanski. Program optimization based on compile-time cache performance prediction. Parallel Processing Letters, 6(1): 173–184, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Kaplow:1997:CTC

[KS97] Wesley K. Kaplow and Boleslaw K.

Szymanski. Compile-time cache performance prediction and its application to tiling. Parallel Processing Letters, 7(4):393– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Khonsari:2005:PMS

[KSAOK05] A. Khonsari, H. Sarbazi-Azad, and M. Ould-Khaoua. A performance model of software-based deadlock recovery routing algorithm in hypercubes. Parallel Processing Letters, 15(1/2):153– 168, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Kronawitter:2014:DSO

[KSHL14] Stefan Kronawitter, Holger Stengel, Georg Hager, and Christian Lengauer. Domainspecific optimization of two Jacobi smoother kernels and their evaluation in the ECM performance model. Parallel Processing Letters, 24(3): 1441004, September 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Khonsari:2007:PMD

[KSOK07] A. Khonsari, A. Shahrabi, and M. Ould-Khaoua. A performance model of Disha routing in k-ary n-cube networks. Parallel Processing Letters, 17(2): 213–228, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Kaufmann:1995:RSR

[KSS95] M. Kaufmann, H. Schroder, and J. F. Sibeyn. Routing and sorting on reconfigurable meshes. Parallel Processing Letters, 5(1): 81–95, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Kaya:2009:EAT

[KU09] Kamer Kaya and Bora Uçar.

Exact algorithms for a task assignment problem. Parallel Processing Letters, 19(3):451– 465, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Kulkarni:2007:TA

[Kul07] Sandeep S. Kulkarni. Terminating alternator. Parallel Processing Letters, 17(4):379– 390, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Kutylowski:1992:PST

[KW92] M. Kutylowski and R. Wanka.

Periodic sorting on two-dimensional meshes. Parallel Processing Letters, 2(2-3):213–220, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Klimke:2006:CDA

[KW06] Andreas Klimke and Barbara Wohlmuth. Constructing dimension-adaptive sparse grid interpolants using parallel function evaluations. Parallel Processing Letters, 16(4):407– 418, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Kumar:2019:IBM

[KW19] Saptaparni Kumar and Jennifer L. Welch. Implementing ⋄P with bounded messages on a network of ADD channels. Parallel Processing Letters, 29(01):??, March 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500026.

Kurzyniec:2003:TSO

[KWDS03] Dawid Kurzyniec, Tomasz Wrzosek, Dominik Drzewiecki, and Vaidy Sunderam. Towards selforganizing distributed computing frameworks: the H2O approach. Parallel Processing Letters, 13(2):273–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Kim:1996:FSL

[KY96] Geunmo Kim and Hyunsoo Yoon. Free submesh list strategy: a best fit submesh allocation in mesh connected multicomputers. Parallel Processing Letters, 6(1):75–86, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Kiniwa:2006:RLT

[KY06] Jun Kiniwa and Masafumi Yamashita. A randomized 1latent, time-adaptive and safe self-stabilizing mutual exclusion protocol. Parallel Processing Letters, 16(1):53–61, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Latifi:1991:SEF

[Lat91] S. Latifi. Subcube embeddability of folded hypercubes. Parallel Processing Letters, 1(1):43– 50, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Lavault:2002:EPI

[Lav02] Christian Lavault. Embeddings into the Pancake interconnection network. Parallel Processing Letters, 12(3–4):297–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Lin:2014:TES

[LBD^{+}14] Paul Lin, Matthew Bettencourt, Stefan Domino, Travis Fisher, Mark Hoemmen, Jonathan Hu, Eric Phipps, Andrey Prokopenko, Sivasankaran Rajamanickam, Christopher Siefert, and Stephen Kennon. Towards extreme-scale simulations for low Mach fluids with secondgeneration Trilinos. Parallel Processing Letters, 24(4): 1442005, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Leidel:2014:HSS

[LC14] John D. Leidel and Yong Chen.

HMC-SIM: A simulation framework for hybrid memory cube devices. Parallel Processing Letters, 24(4):1442002, December 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Lechtchinsky:2002:PFP

[LCK02] R. Lechtchinsky, M. M. T.

Chakravarty, and G. Keller. Parallel functional programming. Parallel Processing Letters, 12 (2):249–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Li:2014:EPP

[LCS^{+}14] Bo Li, Hung-Ching Chang, Shuaiwen Song, Chun-Yi Su, Timmy Meyer, John Mooring, and Kirk Cameron. Extending PowerPack for profiling and analysis of high-performance accelerator-based systems. Parallel Processing Letters, 24(4): 1442001, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Lemeire:2008:MPC

[LDC08] Jan Lemeire, Erik Dirkx, and Walter Colitti. Modeling the performance of communication schemes on network topologies. Parallel Processing Letters, 18 (2):205–220, June 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Lenders:1992:MAA

[Len92] P. M. Lenders. Multi-rate arrays and affine recurrence equations. Parallel Processing Letters, 2(4): 373–380, December 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Lee:1997:MMD

[LF97] Hyuk-Jae Lee and José A. B. Fortes. Modular mappings and data distribution independent computations. Parallel Processing Letters, 7(2):169–180, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Lapillonne:2014:UCD

[LF14] Xavier Lapillonne and Oliver Fuhrer. Using compiler directives to port large scientific applications to GPUs: An example from atmospheric science. Parallel Processing Letters, 24 (1):1450003, March 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Liu:2010:MPP

[LFC^{+}10] Ning Liu, Jing Fu, Christopher D. Carothers, Onkar Sahni, Kenneth E. Jansen, and Mark S. Shephard. Massively parallel I/O for partitioned solver systems. Parallel Processing Letters, 20(4):377–395, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Larsson:1999:SME

[LFJ99] O. Larsson, M. Feig, and L. Johnsson. Some metacomputing experiences for scientific applications. Parallel Processing Letters, 9(2):243–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Latifi:1997:REB

[LG97] Shahram Latifi and Ramesh Gajjala. Reliability evaluation of braided networks using a recursive method. Parallel Processing Letters, 7(1):77–88, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Lee:1999:PEJ

[LGCH99] Bu-Sung Lee, Yan Gu, Wentong Cai, and Alfred Heng. Performance evaluation of JPVM. Parallel Processing Letters, 9(3): 401–??, September 1999. CO-DEN PPLTEE. ISSN 0129-6264.

Lumsdaine:2007:CPG

[LGHB07] Andrew Lumsdaine, Douglas Gregor, Bruce Hendrickson, and Jonathan Berry. Challenges in parallel graph processing. Parallel Processing Letters, 17(1):

5–20, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Lee:1995:SSP

[LH95] B. Lee and A. R. Hurson. A strategy for scheduling partially ordered program graphs onto multicomputers. Parallel Processing Letters, 5(4):575– 586, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Loh:2002:EFT

[LH02] P. K. K. Loh and W. J. Hsu. Embedding of fault-tolerant trees in the Josephus cube. Parallel Processing Letters, 12(1):3–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Lee:1996:TSF

[LHCT96] Bu-Sung Lee, A. Heng, W. Cai, and Tai-Ann Tan. Task scheduling facility for PVM. Parallel Processing Letters, 6(4):563– 574, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Lindon:1992:DAA

[Lin92] L. F. Lindon. Discriminating analysis and its application to matrix by vector multiplication on the PRAM. Parallel Processing Letters, 2(1):43–50, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Lutz:1995:DFC

[LJ95] David R. Lutz and D. N. Jayasimha. Do fixed-processor communication-time tradeoffs exist? Parallel Processing Letters, 5(2):311–320, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Lucas:2010:SRO

[LJ10] Keny T. Lucas and Prasanta K.

Jana. Sorting and routing on OTIS-Mesh of trees. Parallel Processing Letters, 20(2): 145–154, June 2010. CODEN PPLTEE. ISSN 0129-6264.

Liu:2021:ADH

[LL21] Wenjun Liu and Wenjun Li.

Adaptive diagnosis of Hamiltonian networks under the comparison model. Parallel Processing Letters, 31(03): ??, September 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500158.

Lee:2011:TAT

[LLC11] Cheng-Chi Lee, Chun-Ta Li, and Shun-Der Chen. Two attacks on a two-factor user authentication in wireless sensor networks. Parallel Processing Letters, 21 (1):21–26, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Lawley:2003:ANB

[LLD^{+}03] Richard Lawley, Michael Luck, Keith Decker, Terry Payne, and Luc Moreau. Automated negotiation between publishers and consumers of grid notifications. Parallel Processing Letters, 13 (4):537–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Lingas:1997:SOP

[LM97] Andrzej Lingas and Anil Maheshwari. A simple optimal parallel algorithm for reporting paths in a tree. Parallel Processing Letters, 7(1):3–11, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Loechner:2000:MCD

[LM00] V. Loechner and C. Mongenet. Memory cost due to anticipated broadcast. Parallel Processing Letters, 10 (2/3):177–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000184.
html.

Lecroq:1998:OPP

[LMS98] Thierry Lecroq, Jean-Frederic Myoupo, and David Seme. A one-phase parallel algorithm for the sequence alignment problem. Parallel Processing Letters, 8(4): 515–??, December 1998. CO-DEN PPLTEE. ISSN 0129-6264.

Luong:2010:NSG

[LMT10] Thé Van Luong, Nouredine Melab, and El-Ghazali Talbi. Neighborhood structures for GPU-based local search algorithms. Parallel Processing Letters, 20(4):307–324, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Laure:1999:OHC

[LMZ99] Erwin Laure, Piyush Mehrotra, and Hans Zima. Opus: Heterogeneous computing with data parallel tasks. Parallel Processing Letters, 9(2):275–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Lin:2005:GAE

[LN05] Man Lin and Sai Man Ng. A genetic algorithm for energy aware task scheduling in heterogeneous systems. Parallel Processing Letters, 15(4):439–449, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Louca:2000:MFP

[LNLE00] S. Louca, N. Neophytou, A. Lachanas, and P. Evripidou. MPI-FT: Portable fault tolerance scheme for MPI. Parallel Processing Letters, 10 (4):371–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000342.
html.

Looges:1993:PPC

[LO93] P. J. Looges and S. Olariu. A practical platform for CREW emulation. Parallel Processing Letters, 3(2):139–145, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Lefevre:2009:TEA

[LO09] Laurent Lefèvre and Anne-Cecile Orgerie. Towards energy aware reservation infrastructure for large-scale experimental distributed systems. Parallel Processing Letters, 19(3): 419–433, September 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Lin:1993:SEM

[LOSZ93] R. Lin, S. Olariu, J. L. Schwing, and J. Zhang. Simulating enhanced meshes, with applications. Parallel Processing Letters, 3(1):59–70, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Lin:1994:CRB

[LOSZ94] Rong Lin, Stephan Olariu, James L. Schwing, and Jingyuan Zhang. Computing on reconfigurable buses — a new computational paradigm. Parallel Processing Letters, 4(4):465– 476, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Loulergue:2001:DEF

[Lou01] F. Loulergue. Distributed evaluation of functional BSP programs. Parallel Processing Letters, 11(4):423–??, December 2001. CODEN PPLTEE. ISSN 0129-6264.

Levcopoulos:1994:WTP

[LP94] Christos Levcopoulos and Teresa M. Przytycka. A work-time tradeoff in parallel computation of Huffman trees and concave least weight subsequence problem. Parallel Processing Letters, 4(1-2):37–43, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Liestman:1998:MAT

[LP98] Arthur L. Liestman and Nataša Pržulj. Minimum average time broadcast graphs. Parallel Processing Letters, 8(2):139–147, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Lin:2020:SEC

[LPL20] Shangwei Lin, Jianfeng Pei, and Chunfang Li. Super edgeconnected linear hypergraphs. Parallel Processing Letters, 30 (03):??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400034.

Lee:2001:IPA

[LPP01] S. Lee, C.-I. Park, and C.-M.

Park. An improved parallel algorithm for Delaunay triangulation on distributed memory parallel computers. Parallel Processing Letters, 11(2–3):341–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Liestman:1993:PG

[LR93] A. L. Liestman and D. Richards.

Perpetual gossiping. Parallel Processing Letters, 3(4):347– 355, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Liestman:1991:ASH

[LS91] A. L. Liestman and T. C. Shermer. Additive spanners for hypercubes. Parallel Processing Letters, 1(1):35–42, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Lucas:2014:PIF

[LSF14] Andrew Lucas, Mark Stalzer, and John Feo. Parallel implementation of fast randomized algorithms for low rank matrix decomposition. Parallel Processing Letters, 24(1):1450004, March 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Lam:1997:OCM

[LSW97] Chi-Chung Lam, P. Sadayappan, and Rephael Wenger. On optimizing a class of multidimensional loops with reduction for parallel execution. Parallel Processing Letters, 7(2): 157–168, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Loidl:2001:TTG

[LTB01] H.-W. Loidl, P. W. Trinder, and C. Butz. Tuning task granularity and data locality of data parallel G P H programs. Parallel Processing Letters, 11(4): 471–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Li:1998:SPA

[LTZ98] Yueming Li, Jianhui Tao, and S. Q. Zheng. A symmetric processor array with synchronous optical buses and switches. Parallel Processing Letters, 8(3): 283–??, September 1998. CO-DEN PPLTEE. ISSN 0129-6264.

LeGuyadec:1996:SPD

[LV96] Y. Le Guyadec and B. Virot. Sequential-like proofs of data-parallel programs. Parallel Processing Letters, 6(3):415– 426, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Lusk:2007:LHP

[LY07] Ewing Lusk and Katherine Yelick. Languages for highproductivity computing: the DARPA HPCS Language Project. Parallel Processing Letters, 17 (1):89–102, March 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Liu:2020:ECD

[LYW20] Aixia Liu, Jun Yuan, and Shiying Wang. The g-extra conditional diagnosability of graphs in terms of g-extra connectivity. Parallel Processing Letters, 30(03):??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S012962642040006X.

Liu:2021:REB

[LZC^{+}21] Jiafei Liu, Shuming Zhou, Eddie Cheng, Gaolin Chen, and Min Li. Reliability evaluation of bicube-based multiprocessor system under the g-good-neighbor restriction. Parallel Processing Letters, 31(04): ??, December 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500183.

Liu:2019:NIN

[LZG^{+}19] Jiafei Liu, Shuming Zhou, Zhendong Gu, Yihong Wang, and Qianru Zhou. A note of independent number and domination number of Q_{n,k,m} - graph. Parallel Processing Letters, 29(03):??, September 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500117.

Lv:2018:RED

[LZS^{+}18] Mengjie Lv, Shuming Zhou, Xueli Sun, Guanqin Lian, and Jiafei Liu. Reliability evaluation of data center network DCell. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500159.

Lu:2018:RCB

[LZY18] Miao Lu, Shurong Zhang, and Weihua Yang. Regular connected bipancyclic spanning subgraphs of torus networks. Parallel Processing Letters, 28(04):??, December 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500135.

Mackenzie:1993:LBO

[Mac93] P. D. Mackenzie. A lower bound for order-preserving broadcast in the postal model. Parallel Processing Letters, 3(4):313– 320, December 1993. CODEN PPLTEE. ISSN 0129-6264.

MacKenzie:1995:SBR

[Mac95] P. D. MacKenzie. A separation between reconfigurable mesh models. Parallel Processing Letters, 5(1):15–22, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Maclennan:2012:ECA

[Mac12] Bruce J. Maclennan. Embodied computation: Applying the physics of computation to artificial morphogenesis. Parallel Processing Letters, 22(3):

1240013, September 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Maksoud:2009:ECS

[Mak09] Ehab Y. Abdel Maksoud. Efficient combined scheduling of hard and soft real-time tasks in multiprocessor systems under a processing power-share strategy. Parallel Processing Letters, 19 (1):23–38, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Malinowski:1997:EBA

[Mal97] Adam Malinowski. Efficient Byzantine agreement in networks with random faults. Parallel Processing Letters, 7(1): 69–76, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Manzini:1991:LSR

[Man91] G. Manzini. Large sorting and routing problems on the Hypercube and related networks. Parallel Processing Letters, 1(2): 113–124, December 1991. CO-DEN PPLTEE. ISSN 0129-6264.

Mendez:2019:PAR

[MAP^{+}19] Diego Mendez, David Arevalo, Diego Patino, Eduardo Gerlein, and Ricardo Quintana. Parallel architecture of reconfigurable hardware for massive output active noise control. Parallel Processing Letters, 29(03):??, September 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500142.

Marsh:1998:NAH

[Mar98] Andy Marsh. A novel algorithm and its hardware support for the run-time detection of andparallelism. Parallel Processing Letters, 8(4):489–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Marowka:2003:EOT

[Mar03]  Ami Marowka.  Extending

OpenMP for task parallelism. Parallel Processing Letters, 13 (3):341–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Massingill:1999:EPP

[Mas99] B. L. Massingill. Experiments with program parallelization using archetypes and stepwise refinement. Parallel Processing Letters, 9(4):487–??, December 1999. CODEN PPLTEE. ISSN 0129-6264. URL http://www.

                                        wspc.com/journals/ppl/09/
                                        0904/;   http://www.wspc.
                                        com/journals/ppl/09/0904/
                                        1190904.html.

Merry:1995:CTS

[MB95] M. S. Merry and J. Baker. A constant time sorting algorithm for a three dimensional reconfigurable mesh and reconfigurable network. Parallel Processing Letters, 5(3):401–412, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Murshed:1998:CTA

[MB98] M. Manzur Murshed and Richard P. Brent. Constant time algorithms for computing the contour of maximal elements on a reconfigurable mesh. Parallel Processing Letters, 8(3):351– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Moles:2015:DLC

[MBT15] Josh Moles, Peter Banda, and Christof Teuscher. Delay line as a chemical reaction network. Parallel Processing Letters, 25 (1):1540002, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Medarametla:2015:LME

[MCCW15] Dhruv Medarametla, Eddie Cheng, Steven Cheng, and Lawrance Wu. Linearly many edge-faults in 2-bijective connection networks. Parallel Processing Letters, 25(4):1550006, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Madanayake:2012:BPS

[MCDB12] H. L. P. Arjuna Madanayake, R. J. Cintra, V. S. Dimitrov, and L. T. Bruton. Block-parallel systolic-array architecture for 2-D NTT-based fragile watermark embedding. Parallel Processing Letters, 22(3):1250009, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Miller:2001:PIB

[MCI^{+}01] B. P. Miller, M. Christodorescu, R. Iverson, T. Kosar, A. Mirgorodskii, and F. Popovici. Playing inside the black box: Using dynamic instrumentation to create security holes. Parallel Processing Letters, 11(2–3):267–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Merchant:2017:ABL

[MCR^{+}17] Farhad Merchant, Anupam Chattopadhyay, Soumyendu Raha, S. K. Nandy, and Ranjani Narayan. Accelerating BLAS and LAPACK via efficient floating point architecture design. Parallel Processing Letters, 27(3–4):1750006, 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Misale:2017:CBD

[MDAT17] Claudia Misale, Maurizio Drocco, Marco Aldinucci, and Guy Tremblay. A comparison of big data frameworks on a layered dataflow model. Parallel Processing Letters, 27(1):1740003, March 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Merry:2015:PCS

[Mer15] Bruce Merry. A performance comparison of sort and scan libraries for GPUs. Parallel Processing Letters, 25(4): 1550007, December 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Maignan:2009:CAM

[MG09] Luidnel Maignan and Frédéric Gruau. A 1D cellular automaton that moves particles until regular spatial placement. Parallel Processing Letters, 19(2): 315–331, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Merrill:2011:HPS

[MG11] Duane Merrill and Andrew Grimshaw. High performance and scalable radix sorting: a case study of implementing dynamic parallelism for GPU computing. Parallel Processing Letters, 21(2):245–272, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Marquez:2007:IPM

[MGBG07] A. Márquez, C. Gil, R. Baños, and J. Gómez. Improving the performance of multi-objective evolutionary algorithms using the island parallel model. Parallel Processing Letters, 17(2): 127–139, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Maigne:2004:PMC

[MHC^{+}04] Lydia Maigne, David Hill, Pascal Calvat, Vincent Breton, Romain Reuillon, Delphine Lazaro, Yannick Legre, and Denise Donnarieix. Parallelization of Monte Carlo simulations and submission to a grid environment. Parallel Processing Letters, 14(2): 177–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Matsuzaki:2005:SDT

[MHKT05] Kiminori Matsuzaki, Zhenjiang Hu, Kazuhiko Kakehi, and Masato Takeichi. Systematic derivation of tree contraction algorithms. Parallel Processing Letters, 15(3):321–336, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Michail:1998:OBS

[Mic98] Amir Michail. Optimal broadcast and summation on hierarchical ring architectures. Parallel Processing Letters, 8(1): 83–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Michalakes:2016:OWM

[MIJ16] John Michalakes, Michael J. Iacono, and Elizabeth R. Jessup. Optimizing weather model radiative transfer physics for Intel’s Many Integrated Core (MIC) architecture. Parallel Processing Letters, 26(4): 1650019, December 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Morin:2009:CNP

[MJGR09] Christine Morin, Yvon Jégou, Jérôme Gallard, and Pierre Riteau. Clouds: a new playground for the XtreemOS Grid operating system. Parallel Processing Letters, 19(3):435– 449, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Marsh:1998:UPP

[MKA98] Andy Marsh, Dimitra I. Kaklamani, and Konstantinos Adam. Using parallel processing as a computational tool to solve multi-plate electromagnetic scattering problems. Parallel Processing Letters, 8(4):535–??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Megson:1999:FTR

[MLY99] G. M. Megson, X. Liu, and X. Yang. Fault-tolerant ring embedding in a honeycomb torus with node failures. Parallel Processing Letters, 9(4): 551–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

6264.  URL http://www.
wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Mishra:2016:RSA

[MM16] Abhishek Mishra and Pramod Kumar Mishra. A randomized scheduling algorithm for multiprocessor environments using local search. Parallel Processing Letters, 26(1):1650002, March 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Mendiburu:2006:IPE

[MMAL06] A. Mendiburu, J. Miguel-Alonso, and J. A. Lozano. Implementation and performance evaluation of a parallelization of estimation of Bayesian network algorithms. Parallel Processing Letters, 16(1):133–148, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Ma:2018:FMP

[MMCW18] Tianlong Ma, Yaping Mao, Eddie Cheng, and Jinling Wang. Fractional matching preclusion for (n, k)-star graphs. Parallel Processing Letters, 28(04): ??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626418500172.

Morajko:2003:APA

[MMJ^{+}03] Anna Morajko, Oleg Morajko, Josep Jorba, Tomàs Margalef, and Emilio Luque. Automatic performance analysis and dynamic tuning of distributed applications. Parallel Processing Letters, 13(2):169–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Milovanovic:1994:STL

[MMMS94] E. I. Milovanovic, M. D. Mihajlovic, I. Z. Milovanovic, and M. K. Stojcev. Solving tridiagonal linear systems on MIMD computers. Parallel Processing Letters, 4(1-2):53–64, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Mishra:2012:RSA

[MMMT12] Pramod Kumar Mishra, Kamal Sheel Mishra, Abhishek Mishra, and Anil Kumar Tripathi. A randomized scheduling algorithm for multiprocessor environments. Parallel Processing Letters, 22(4):1250015, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Mostefaoui:2006:TFA

[MMRT06] Achour Mostefaoui, Eric Mourgaya, Michel Raynal, and Corentin Travers. A time-free assumption to implement eventual leadership. Parallel Processing Letters, 16(2):189–207, June 2006. CODEN PPLTEE. ISSN 0129-6264.

Mancini:2017:PCB

[MMT17] Toni Mancini, Annalisa Massini, and Enrico Tronci. Parallelization of cycle-based logic simulation. Parallel Processing Letters, 27(2):1750003, June 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Mizuno:1996:OPL

[MNR96] M. Mizuno, M. L. Neilsen, and M. Raynal. An optimistic protocol for a linearizable distributed shared memory system. Parallel Processing Letters, 6(2): 265–278, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Mongenet:1994:DCS

[Mon94] Catherine Mongenet. Data compiling for systems of uniform recurrence equations. Parallel Processing Letters, 4(3):245– 257, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Malard:1993:DRD

[MP93] J. Malard and C. C. Paige. Data replication in dense matrix factorization. Parallel Processing Letters, 3(4):419–430, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Mostefaoui:2018:SOS

[MPR18] Achour Mostefaoui, Matthieu Perrin, and Michel Raynal. A simple object that spans the whole consensus hierarchy. Parallel Processing Letters, 28(02):??, June 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500068.

Mostefaoui:2001:LBC

[MR01] A. Mostefaoui and M. Raynal. Leader-based consensus. Parallel Processing Letters, 11(1): 95–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Marchal:2007:SAD

[MRRV07] Loris Marchal, Veronika Rehn, Yves Robert, and Frédéric Vivien. Scheduling algorithms for data redistribution and loadbalancing on master-slave platforms. Parallel Processing Letters, 17(1):61–77, March 2007. CODEN PPLTEE. ISSN 0129-6264.

Medina:2004:ASH

[MRS04] Maira T. Medina, Celso C. Ribeiro, and Luiz F. G. Soares. Automatic scheduling of hypermedia documents with elastic times. Parallel Processing Letters, 14(1):45–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Maheshwari:1999:SOA

[MS99] Anil Maheshwari and Jörg-Rüdiger Sack. Simple optimal algorithms for rectilinear link path and polygon separation problems. Parallel Processing Letters, 9(1):31–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Migliardi:2001:EPP

[MS01a] M. Migliardi and V. Sunderam.

Emulating parallel programming environments in the harness metacomputing system. Parallel Processing Letters, 11 (2–3):281–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Mongelli:2001:PPM

[MS01b] H. Mongelli and S. W. Song.

Parallel pattern matching with scaling. Parallel Processing Letters, 11(1):125–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Malony:2005:OCP

[MS05] Allen D. Malony and Sameer S.

Shende. Overhead compensation in performance profiling. Parallel Processing Letters, 15(1/2): 19–35, March/June 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Margenstern:2009:NUC

[MS09] Maurice Margenstern and Yu Song.

A new universal cellular automaton on the Pentagrid. Parallel Processing Letters, 19(2): 227–246, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Mahesh:1995:PBS

[MSP95] S. Mahesh, C. Siva Ram Murthy, and C. Pandu Rangan. Performance bounds for static multiprocessor scheduling of multi-task jobs. Parallel Processing Letters, 5(3):343– 355, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Michail:2021:BRG

[MST21] Othon Michail, Paul G. Spirakis, and Michail Theofilatos. Beyond rings: Gathering in 1interval connected graphs. Parallel Processing Letters, 31(04): ??, December 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500201.

Maurer:2016:TRB

[MT16] Alexandre Maurer and Sebastien Tixeuil. Tolerating random Byzantine failures in an unbounded network. Parallel Processing Letters, 26(1):1650003, March 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Mattes:1998:PIP

[MTD98] J. Mattes, D. Trystram, and J. Demongeot. Parallel image processing using neural networks: Applications in contrast enhancement of medical images. Parallel Processing Letters, 8(1): 63–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Monien:2004:IBC

[MV04] Burkhard Monien and Imrich Vrťo. Improved bounds on cutwidths of shuffle-exchange and de Bruijn graphs. Parallel Processing Letters, 14(3/4):361– ??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264.

Mejias:2008:RRF

[MV08a] Boris Mejías and Peter Van Roy. The relaxed-ring: a fault-tolerant topology for structured overlay networks. Parallel Processing Letters, 18(3): 411–432, September 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Michalakes:2008:GAN

[MV08b] John Michalakes and Manish Vachharajani. GPU acceleration of numerical weather prediction. Parallel Processing Letters, 18(4):531–548, December 2008. CODEN PPLTEE. ISSN 0129-6264.

Ma:2019:REC

[MWZ19] Tianlong Ma, Jinling Wang, and Mingzu Zhang. The restricted edge-connectivity of Kronecker product graphs. Parallel Processing Letters, 29(03): ??, September 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500129.

Myoupo:1992:SLS

[Myo92] J. F. Myoupo. Synthesizing linear systolic arrays for dynamic programming problems. Parallel Processing Letters, 2(1):97–110, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Nethi:2005:AMS

[NA05] Murali K. Nethi and James H.

Aylor. Advances in modelling and simulation of large parallel/distributed systems. Parallel Processing Letters, 15(4):

397–405, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Nagy:2007:PQI

[NA07a] Marius Nagy and Selim G. Akl. Parallelism in quantum information processing defeats the universal computer. Parallel Processing Letters, 17(3):233– 262, September 2007. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2007:ABC

[NA07b] Naya Nagy and Selim G. Akl. Aspects of biomolecular computing. Parallel Processing Letters, 17(2):185–211, June 2007. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2007:AQK

[NA07c] Naya Nagy and Selim G. Akl. Authenticated quantum key distribution without classical communication. Parallel Processing Letters, 17(3):323–335, September 2007. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2010:CDP

[NA10a] Marius Nagy and Selim G. Akl. Coping with decoherence: Parallelizing the Quantum Fourier Transform. Parallel Processing Letters, 20(3):213–226, September 2010. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2010:EVA

[NA10b] Marius Nagy and Selim G. Akl. Entanglement verification with an application to quantum key distribution protocols. Parallel Processing Letters, 20(3):227– 237, September 2010. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2010:OTP

[NA10c] Naya Nagy and Selim G. Akl.

One-time pads without prior encounter. Parallel Processing Letters, 20(3):263–273, September 2010. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2010:QCS

[NA10d] Naya Nagy and Selim G. Akl.

A quantum cryptographic solution to the problem of access control in a hierarchy. Parallel Processing Letters, 20(3): 251–261, September 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Nagy:2009:LMT

[Nag09] Marius Nagy. Locating the median of a tree in real time. Parallel Processing Letters, 19(1): 39–55, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Nakano:1995:BPP

[Nak95a] K. Nakano. A bibliography of published papers on dynamically reconfigurable architectures. Parallel Processing Letters, 5(1):111–124, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Nakano:1995:PAR

[Nak95b] K. Nakano. Prefix-sums algorithms on reconfigurable meshes. Parallel Processing Letters, 5(1): 23–35, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Neocleous:2007:FMG

[NDFM07] Kyriakos Neocleous, Marios D. Dikaiakos, Paraskevi Fragopoulou, and Evangelos P. Markatos. Failure management in Grids: the case of the EGEE infrastructure. Parallel Processing Letters, 17(4):391–410, December 2007. CODEN PPLTEE. ISSN 0129-6264.

Nakashima:2006:COP

[NF06] Takaaki Nakashima and Akihiro Fujiwara. A cost optimal parallel algorithm for patience sorting. Parallel Processing Letters, 16(1):39–51, March 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Ning:2020:CEF

[Nin20] Wantao Ning. The connectivity of exchanged folded hypercube. Parallel Processing Letters, 30(01):??, March 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500036.

Nitsche:2002:LSF

[Nit02] T. Nitsche. Lifting sequential functions to parallel skeletons. Parallel Processing Letters, 12 (2):267–??, June 2002. CODEN PPLTEE. ISSN 0129-6264.

Nitsche:2005:DSC

[Nit05] Thomas Nitsche. Deriving and scheduling communication operations for generic skeleton implementations. Parallel Processing Letters, 15(3):337–352, September 2005. CODEN PPLTEE. ISSN 0129-6264.

Narayanan:2003:DSD

[NKCS03] Sivaramakrishnan Narayanan, Tahsin Kurc, Umit Catalyurek, and Joel Saltz. Database support for data-driven scientific applications in the Grid. Parallel Processing Letters, 13(2): 245–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Nicol:1996:BPK

[NM96] D. M. Nicol and Weizhan Mao.

On bottleneck partitioning of kary n-cubes. Parallel Processing Letters, 6(3):389–399, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Nagy:2010:KDV

[NNA10] Naya Nagy, Marius Nagy, and Selim G. Akl. Key distribution versus key enhancement in quantum cryptography. Parallel Processing Letters, 20(3): 239–250, September 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Novakovic:2020:BCS

[Nov20] Vedran Novaković. Batched computation of the singular value decompositions of order two by the AVX-512 vectorization. Parallel Processing Letters, 30(04):??, December 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500152.

Nikolopoulos:2004:PAR

[NP04] Stavros D. Nikolopoulos and Leonidas Palios. Parallel algorithms for recognizing p_{5}- free and p̄_{5}-free weakly chordal graphs. Parallel Processing Letters, 14(1):119–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Norman:1995:CSC

[NPT95] M. G. Norman, S. Pelagatti, and P. Thanisch. On the complexity of scheduling with communication delay and contention. Parallel Processing Letters, 5(3): 331–341, September 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Niedermeier:1995:OOA

[NR95] Rolf Niedermeier and Peter Rossmanith. On optimal OROW-PRAM algorithms for computing recursively defined functions. Parallel Processing Letters, 5(2):299–309, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Nivat:1994:EPS

[NRP^{+}94] M. Nivat, G. D. S. Ramkumar, C. Pandu Rangan, A. Saoudi, and R. Sundaram. Efficient parallel shuffle recognition. Parallel Processing Letters, 4(4):455– 463, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Nguyen-Tuong:1999:URI

[NTG99] Anh Nguyen-Tuong and Andrew S. Grimshaw. Using reflection for incorporating faulttolerance techniques into distributed applications. Parallel Processing Letters, 9(2):291–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Neggazi:2017:MSS

[NTHK17] Brahim Neggazi, Volker Turau, Mohammed Haddad, and Hamamache Kheddouci. A O(m) self-stabilizing algorithm for maximal triangle partition of general graphs. Parallel Processing Letters, 27(2):1750004, June 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Ngonmang:2012:LCI

[NTV12] Blaise Ngonmang, Maurice Tchuente, and Emmanuel Viennet. Local community identification in social networks. Parallel Processing Letters, 22(1): 1240004, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Ning:1995:ADC

[NvG95] Qi Ning, V. van Dongen, and G. R. Gao. Automatic data and computation decomposition for distributed-memory machines. Parallel Processing Letters, 5(4): 539–550, December 1995. CO-DEN PPLTEE. ISSN 0129-6264.

ODonnell:1994:CPP

[O’D94] John T. O’Donnell. A correctness proof of parallel scan. Parallel Processing Letters, 4(3): 329–338, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Ooshita:2012:PMO

[OII12] Fukuhito Ooshita, Tomoko Izumi, and Taisuke Izumi. The price of multi-organization constraint in unrelated parallel machine scheduling. Parallel Processing Letters, 22(2):1250006, June 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Okoubi:2020:PND

[OK20] Firmin Andzembe Okoubi and Jonas Koko. Parallel Nesterov domain decomposition method for elliptic partial differential equations. Parallel Processing Letters, 30(01):??, March 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500048.

Oskin:2002:OST

[OKH^{+}02] Mark Oskin, Diana Keen, Justin Hensley, Lucian-Vlad Lita, and Frederic T. Chong. Operating systems techniques for parallel computation in intelligent memory. Parallel Processing Letters, 12(3–4):311–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Oprescu:2011:BEC

[OKL11] Ana-Maria Oprescu, Thilo Kielmann, and Haralambie Leahu. Budget estimation and control for bag-of-tasks scheduling in clouds. Parallel Processing Letters, 21(2):219–243, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Oksa:1995:TLS

[Oks95] G. Oksa. Time-minimal linear systolic arrays for the Toeplitz system of linear equations. Parallel Processing Letters, 5(3):

461–474, September 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Orgerie:2011:EEE

[OL11] Anne-Cecile Orgerie and Laurent Lefevre. Eridis: Energyefficient reservation infrastructure for large-scale distributed systems. Parallel Processing Letters, 21(2):133–154, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Oskin:2000:ACP

[OLC^{+}00] M. Oskin, L.-V. Lita, F. T.

Chong, J. Hensley, and D. Keen. Algorithmic complexity with page-based intelligent memory. Parallel Processing Letters, 10(1):99–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000111.
html.

Olariu:1993:SAL

[OSW93] S. Olariu, W. Shen, and L. Wilson. Sub-logarithmic algorithms for the largest empty rectangle problem. Parallel Processing Letters, 3(1):79–85, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Olariu:1991:PTP

[OSZ91] S. Olariu, J. L. Schwing, and Jingyuan Zhang. On the power of two-dimensional processor arrays with reconfigurable bus systems. Parallel Processing Letters, 1(1):29–34, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Obrecht:2015:PEO

[OTK15] Christian Obrecht, Bernard Tourancheau, and Frédéric Kuznik. Performance evaluation of an OpenCL implementation of the Lattice Boltzmann Method on the Intel Xeon Phi. Parallel Processing Letters, 25 (3):1541001, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Oksa:2006:PPB

[OV06] Gabriel Okša and Marián Vajteršic. Preconditioned parallel block-Jacobi SVD algorithm. Parallel Processing Letters, 16 (3):371–379, September 2006. CODEN PPLTEE. ISSN 0129-6264.

Olariu:1992:PAF

[OW92] S. Olariu and Z. Wen. A parallel algorithm for forest reconstruction. Parallel Processing Letters, 2(2-3):157–160, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Pavel:1995:EAE

[PA95] Sandy Pavel and Selim G. Akl. Efficient algorithms for the Euclidean distance transform. Parallel Processing Letters, 5(2): 205–212, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Pham:1999:TMA

[PA99] C. Pham and C. Albrecht. Tuning message aggregation on high performance clusters for efficient parallel simulations. Parallel Processing Letters, 9(4): 521–??, December 1999. CO-DEN PPLTEE. ISSN 0129-

6264.  URL http://www.
wspc.com/journals/ppl/09/
0904/;   http://www.wspc.
com/journals/ppl/09/0904/
1190904.html.

Plank:2003:AHP

[PADB03] James S. Plank, Scott Atchley, Ying Ding, and Micah Beck. Algorithms for high performance, wide-area distributed file downloads. Parallel Processing Letters, 13(2):207–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Pelaez:2008:HLP

[PAG08] Ignacio Peláez, Francisco Almeida, and Daniel González. High level parallel skeletons for dynamic programming. Parallel Processing Letters, 18(1):133–147, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Page:2013:PAS

[Pag13] Daniel R. Page. Parallel algorithm for second-order restricted weak integer composition generation for shared memory machines. Parallel Processing Letters, 23(3):1350010, September 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Pionteck:2008:ACA

[PAKM08] Thilo Pionteck, Carsten Albrecht, Roman Koch, and Erik Maehle. Adaptive communication architectures for runtime reconfigurable system-on-chips. Parallel Processing Letters, 18 (2):275–289, June 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Pan:1994:MEC

[Pan94] Yi Pan. A more efficient constant time algorithm for computing the Hough transform. Parallel Processing Letters, 4(1-2):45–52, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Panaite:1997:RPG

[Pan97] Petrişor Panaite. Routing permutations on a 2D grid with oneway edges. Parallel Processing Letters, 7(3):225–??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Pancake:2003:UID

[Pan03] Cherri M. Pancake. Usability issues in developing tools for the Grid — and how visual representations can help. Parallel Processing Letters, 13(2): 189–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Parberry:1992:PSN

[Par92] I. Parberry. The pairwise sorting network. Parallel Processing Letters, 2(2-3):205–211, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Pal:1996:OPA

[PB96] M. Pal and G. P. Bhattacharjee. An optimal parallel algorithm to color an interval graph. Parallel Processing Letters, 6(4):439– 449, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Purushotham:1992:PEL

[PBKP92] B. V. Purushotham, A. Basu, P. S. Kumar, and L. M. Patnaik. Performance estimation of LU factorisation on message passing multiprocessors. Parallel Processing Letters, 2(1):51–60, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Perumalla:1995:PAM

[PD95] K. Perumalla and N. Deo. Parallel algorithms for maximum subsequence and maximum subarray. Parallel Processing Letters, 5(3):367–373, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Pelc:1992:BTS

[Pel92] A. Pelc. Broadcasting time in sparse networks with faulty transmissions. Parallel Processing Letters, 2(4):355–361, December 1992. CODEN PPLTEE. ISSN 0129-6264.

Pellegrini:1993:BBD

[Pel93]  F. Pellegrini.  Bounds for

the bandwidth of the d-ary de Bruijn graph. Parallel Processing Letters, 3(4):431– 443, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Pelc:1995:ABA

[Pel95] A. Pelc. Almost-safe Byzantine agreement in sparse networks with random faults. Parallel Processing Letters, 5(3): 489–497, September 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Pelc:2006:VMA

[Pel06] Andrzej Pelc. Voting mechanisms in asynchronous Byzantine environments. Parallel Processing Letters, 16(1):93–99, March 2006. CODEN PPLTEE. ISSN 0129-6264.

Persiano:1994:OAD

[Per94] Giuseppe Persiano. An optimal algorithm for the dining philosophers problem. Parallel Processing Letters, 4(1-2):181–187, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Perepu:2021:OIP

[Per21] Pavan Kumar Perepu. OpenMP implementation of parallel longest common subsequence algorithm for mathematical expression retrieval. Parallel Processing Letters, 31(02):??, June 2021. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500079.

Petit:2003:CSS

[Pet03] Jordi Petit. Combining spectral sequencing and parallel simulated annealing for the Minla problem. Parallel Processing Letters, 13(1):77–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Papadakis:2009:HBO

[PFM^{+}09] Harris Papadakis, Paraskevi Fragopoulou, Evangelos P. Markatos, Marios D. Dikaiakos, and Alexandros Labrinidis. Hash-based overlay partitioning in unstructured peer-to-peer systems. Parallel Processing Letters, 19(1):57–71, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Petrank:2004:PCG

[PK04] Erez Petrank and Elliot K.

Kolodner. Parallel copying garbage collection using delayed allocation. Parallel Processing Letters, 14(2):271–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Poldner:2008:IFS

[PK08] Michael Poldner and Herbert Kuchen. On implementing the farm skeleton. Parallel Processing Letters, 18(1):117–131, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Pohl:2003:OPC

[PKW^{+}03] Thomas Pohl, Markus Kowarschik, Jens Wilke, Klaus Iglberger, and Ulrich Rüde. Optimization and profiling of the cache performance of parallel lattice Boltzmann codes. Parallel Processing Letters, 13(4):549–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Pan:2011:TCR

[PL11] Tien-Tai Pan and Shun-Shii Lin.

The transitive closure and related algorithms of digraph on the reconfigurable architecture. Parallel Processing Letters, 21 (1):27–43, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Palis:1995:OSD

[PLR^{+}95] M. A. Palis, Jing-Chiou Liou, S. Rajasekaran, S. Shende, and D. S. L. Wei. Online scheduling of dynamic trees. Parallel Processing Letters, 5(4):635– 646, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Phoha:2012:DFD

[PMW^{+}12] Shashi Phoha, Goutham Mallapragada, Yicheng Wen, Doina Bein, and Asok Ray. Designing a fusion-driven sensor network to selectively track mobile targets. Parallel Processing Letters, 22(1):1250001, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Panda:2002:PAH

[PND02] B. S. Panda, V. Natarajan, and S. K. Das. Parallel algorithms for Hamiltonian 2separator chordal graphs. Parallel Processing Letters, 12(1): 51–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Paek:1997:CSM

[PP97] Yunheung Paek and David A. Padua. Compiling for scalable multiprocessors with Polaris. Parallel Processing Letters, 7(4): 425–??, December 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Phillips:2001:PDP

[PP01] C. Phillips and R. Perrott. Problems with data parallelism. Parallel Processing Letters, 11(1): 77–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Papadakis:2012:LCR

[PPF12] Harris Papadakis, Costas Panagiotakis, and Paraskevi Fragopoulou. Locating communities on real dataset graphs using synthetic coordinates. Parallel Processing Letters, 22(1):1240003, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Plata:1993:EPA

[PPRZ93] O. Plata, T. F. Pena, F. F.

Rivera, and E. L. Zapata. An efficient processor allocation for nested parallel loops on distributed memory hypercubes. Parallel Processing Letters, 3(2): 179–187, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Pollner:2012:PCC

[PPV12] Peter Pollner, Gergely Palla, and Tamas Vicsek. Parallel clustering with CFinder. Parallel Processing Letters, 22(1): 1240001, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Popov:2003:PAB

[PRH^{+}03] Konstantin Popov, Mahmoud Rafea, Fredrik Holmgren, Per Brand, Vladimir Vlassov, and Seif Haridi. Parallel agent-based simulation on a cluster of workstations. Parallel Processing Letters, 13(4):629–??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Picquendar:1995:CSD

[PRS95] M. Picquendar, A. L. Rosenberg, and V. Scarano. A cost-effective streamlining of the DIOGENES design methodology. Parallel Processing Letters, 5(3):513–524, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Pradeep:1993:CTA

[PS93] B. Pradeep and C. Siva Ram Murthy. A constant time algorithm for redundancy elimination in task graphs on processor arrays with reconfigurable bus systems. Parallel Processing Letters, 3(2):171–177, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Peters:1998:GCC

[PS98] Joseph G. Peters and Curtis C. Spencer. Global communication on circuit-switched toroidal meshes. Parallel Processing Letters, 8(2):161–175, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Paterson:2002:PCA

[PSSV02] M. S. Paterson, H. Schröder, O. Sýkora, and I. Vrto. Permutation communication in alloptical rings. Parallel Processing Letters, 12(1):23–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Papatriantafilou:1997:SSW

[PT97] Marina Papatriantafilou and Philippas Tsigas. On selfstabilizing wait-free clock synchronization. Parallel Processing Letters, 7(3):321–??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Prylli:1998:EDS

[PT98] L. Prylli and B. Tourancheau.

Execution-driven simulation of parallel applications. Parallel Processing Letters, 8(1):95–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Plank:2007:ENA

[PT07] James S. Plank and Michael G.

Thomason. An exploration of non-asymptotic low-density, parity check erasure codes for wide-area storage applications. Parallel Processing Letters, 17 (1):103–123, March 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Purcz:2006:CCS

[Pur06] Pavol Purcz. Communication complexity and speed-up in the explicit difference method. Parallel Processing Letters, 16(3): 313–321, September 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Petit:1999:OSS

[PV99] Franck Petit and Vincent Villain. Optimality and selfstabilization in rooted tree networks. Parallel Processing Letters, 9(3):313–??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Petit:2000:OSS

[PV00] F. Petit and V. Villain. Optimality and self-stabilization in rooted tree networks. Parallel Processing Letters, 10(1):3–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/

                                        10/1001/S0129626400000032.
                                        html.

Qiao:1995:DCA

[Qia95]  C. Qiao.  On designing

communication-intensive algorithms for a spanning optical bus based array. Parallel Processing Letters, 5(3):499–511, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Qiu:2007:UNB

[Qiu07] Ke Qiu. On a unified neighbourhood broadcasting scheme for interconnection networks. Parallel Processing Letters, 17(4): 425–437, December 2007. CO-DEN PPLTEE. ISSN 0129-6264.

Qiu:2010:GEN

[Qiu10] Ke Qiu. Guest Editor’s note: Advances in quantum computation. Parallel Processing Letters, 20(3):211–212, September 2010. CODEN PPLTEE. ISSN 0129-6264.

Qiao:1998:IEB

[QM98] Chunming Qiao and Yousong Mei. An improved embedding of binary trees in a square reconfigurable array with spanning optical buses. Parallel Processing Letters, 8(3):321–??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Quinton:1997:MPU

[QRR97] Patrice Quinton, Sanjay Rajopadhye, and Tanguy Risset. On manipulating Z-polyhedra using a canonical representation. Parallel Processing Letters, 7(2):

181–194, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Quinlan:2000:RCS

[Qui00]  D. Quinlan.  ROSE: Com-

piler support for object-oriented frameworks. Parallel Processing Letters, 10(2/3):215– ??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL

http://ejournals.wspc.com.
sg/ppl/10/1002_03/S0129626400000214.
html.

Qin:2020:SCC

[QY20] Chengfu Qin and Weihua Yang.

5-shredders of contractioncritical 5-connected graphs. Parallel Processing Letters, 30(03): ??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400083.

Raghavan:1998:EPS

[Rag98] Padma Raghavan. Efficient parallel sparse triangular solution using selective inversion. Parallel Processing Letters, 8(1): 29–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Rajasekaran:2002:EPA

[Raj02] Sanguthevar Rajasekaran. Efficient parallel algorithms for template matching. Parallel Processing Letters, 12(3–4):359–??, September–December 2002. CO-DEN PPLTEE. ISSN 0129-6264.

Ramesh:1994:RMR

[Ram94] R. Ramesh. R^{2}-M: A reconfigurable rewrite machine. Parallel Processing Letters, 4(1-2): 171–180, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Rantakokko:2005:DMO

[Ran05] Jarmo Rantakokko. A dynamic MPI–OpenMP model for structured adaptive mesh refinement. Parallel Processing Letters, 15 (1/2):37–47, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Raynal:2021:VBB

[Ray21] Michel Raynal. On the versatility of Bracha’s Byzantine reliable broadcast algorithm. Parallel Processing Letters, 31(03): ??, September 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500067.

Rowley:1993:NAH

[RB93] R. Rowley and B. Bose. On the number of arc-disjoint Hamiltonian circuits in the de Bruijn graph. Parallel Processing Letters, 3(4):375–380, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Ramaswamy:1995:SAS

[RB95] S. Ramaswamy and P. Banerjee. Simultaneous allocation and scheduling using convex programming techniques. Parallel Processing Letters, 5(4):587– 598, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Reza:2013:RMI

[RB13] Sajjid Reza and Gregory T.

Byrd. Reducing migrationinduced misses in an oversubscribed multiprocessor system. Parallel Processing Letters, 23(1):1350006, March 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rufai:2005:MPO

[RBAA05] Raimi Rufai, Muslim Bozyigit, Jaralla Alghamdi, and Moataz Ahmed. Multithreaded parallelism with OpenMP. Parallel Processing Letters, 15(4):367– 378, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Ribeiro:2015:FEE

[RBS15] Roberto Ribeiro, João Barbosa, and Luís Paulo Santos. A framework for efficient execution of data parallel irregular applications on heterogeneous systems. Parallel Processing Letters, 25 (2):1550004, June 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rijpkema:2000:DPN

[RDK00] E. Rijpkema, E. F. Deprettere, and B. Kienhuis. Deriving process networks from nested loop algorithms. Parallel Processing Letters, 10 (2/3):165–??, September 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000172.
html.

Rescigno:1997:FTP

[Res97] Adele Anna Rescigno. Faulttolerant parallel communication in the star network. Parallel Processing Letters, 7(1):57–68, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Ruby:2006:ODC

[RGM06] Catherine L. Ruby, Mark L. Green, and Russ Miller. The Operations Dashboard: a collaborative environment for monitoring virtual organization-specific compute element operational status. Parallel Processing Letters, 16(4):485–500, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Ramroach:2021:ADP

[RJ21] Sterling Ramroach and Ajay Joshi. Accelerating data-parallel neural network training with weighted-averaging reparameterisation. Parallel Processing Letters, 31(02):??, June 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500092.

Ranawake:1991:LBC

[RLG91] U. A. Ranawake, P. M. Lenders, and S. M. Goodnick. On lower bounds for the communication volume in distributed systems. Parallel Processing Letters, 1(2): 125–133, December 1991. CO-DEN PPLTEE. ISSN 0129-6264.

Robertazzi:2011:PFS

[Rob11] Thomas G. Robertazzi. A product form solution for tree networks with divisible loads. Parallel Processing Letters, 21(1): 13–20, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Rahmouni:1994:LSA

[ROJ94] Maher Rahmouni, Kevin O’Brien, and Ahmed A. Jerraya. A loopbased scheduling algorithm for hardware description languages. Parallel Processing Letters, 4(3): 351–364, September 1994. CO-DEN PPLTEE. ISSN 0129-6264.

Raman:1993:UAP

[RP93] A. Raman and C. Pandu Rangan. A unified approach to parallel algorithms for the domatic partition problem on special classes of perfect graphs. Parallel Processing Letters, 3(3): 233–241, September 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Risset:1991:SPA

[RR91] T. Risset and Y. Robert. Synthesis of processor arrays for the algebraic path problem: unifying old results and deriving new architectures. Parallel Processing Letters, 1(1):19–28, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Rodriguez:2008:IRI

[RR08] Noemi Rodriguez and Silvana Rossetto. Integrating remote invocations with asynchronism and cooperative multitasking. Parallel Processing Letters, 18 (1):71–85, March 2008. CODEN PPLTEE. ISSN 0129-6264.

Rajasingh:2015:LTA

[RRM15] Indra Rajasingh, R. Sundara Rajan, and Paul Manuel. A linear time algorithm for embedding Christmas trees into certain trees. Parallel Processing Letters, 25(4):1550008, December 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rodriguez:2012:EMH

[RRS12] Pedro Rodríguez, María Cecilia Rivara, and Isaac D. Scherson. Exploiting the memory hierarchy of multicore systems for parallel triangulation refinement. Parallel Processing Letters, 22(3):1250007, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Ravishankar:1996:GRR

[RS96] Krishnamurthi Ravishankar and Suresh Singh. Gossiping on a ring with radios. Parallel Processing Letters, 6(1):115– 126, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Rajasekaran:1997:DRA

[RS97] Sanguthevar Rajasekaran and Sartaj Sahni. Deterministic routing on the array with reconfigurable optical buses. Parallel Processing Letters, 7(3):219– ??, September 1997. CODEN PPLTEE. ISSN 0129-6264.

Ramesh:1998:ISP

[RS98a] S. Ramesh and Chandrashekhar M. Shetty. Impossibility of synchronization in the presence of preemption. Parallel Processing Letters, 8(1):111–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Rottger:1998:EDG

[RS98b] Markus Röttger and Ulf-Peter Schroeder. Embedding 2dimensional grids into optimal hypercubes with edgecongestion 1 or 2. Parallel Processing Letters, 8(2):231–242, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Rosenberg:1995:RRP

[RSS95] A. L. Rosenberg, V. Scarano, and R. K. Sitaraman. The reconfigurable ring of Processors: efficient algorithms via hypercube simulation. Parallel Processing Letters, 5(1):37–48, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Rozman:2006:CPL

[RšT06] Igor Rozman, Marjan šterk, and Roman Trobec. Communication performance of LAM/MPI and MPICH on a Linux cluster. Parallel Processing Letters, 16(3): 323–334, September 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Rai:1992:RTF

[RT92] S. Rai and J. L. Trahan. A reconfiguration technique for fault tolerance in a hypercube. Parallel Processing Letters, 2(2-3): 129–138, September 1992. CO-DEN PPLTEE. ISSN 0129-6264.

Roch:1996:FPC

[RV96] J.-L. Roch and G. Villard. Fast parallel computation of the Jordan normal form of matrices. Parallel Processing Letters, 6(2): 203–212, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Rescigno:1998:HFT

[RV98] A. A. Rescigno and U. Vaccaro. Highly fault-tolerant routing in the star and hypercube interconnection networks. Parallel Processing Letters, 8(2): 221–230, June 1998. CODEN PPLTEE. ISSN 0129-6264.

Rieksts:2009:TRF

[RV09] Brian Q. Rieksts and Jose A. Ventura. Time-relaxed 1-fault tolerant broadcast networks. Parallel Processing Letters, 19 (2):335–353, June 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Reijns:2001:PAM

[RvGG01] G. L. Reijns, A. J. C. van Gemund, and H. Gautama. Performance analysis of multi-stage interconnection networks with deterministic service times. Parallel Processing Letters, 11(1): 109–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Rus:1998:UMC

[RvW98] Teodor Rus and Eric van Wyk. Using model checking in a parallelizing compiler. Parallel Processing Letters, 8(4):459– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Sahni:1995:DMD

[Sah95] S. Sahni. Data manipulation on the distributed memory bus computer. Parallel Processing Letters, 5(1):3–14, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Soliman:2010:SIP

[SAJ10] Mostafa I. Soliman and Abdulmajid F. Al-Junaid. SystemC implementation and performance evaluation of a decoupled general-purpose matrix processor. Parallel Processing Letters, 20(2):103–121, June 2010. CODEN PPLTEE. ISSN 0129-6264.

Santos:2003:TSM

[San03] Eunice E. Santos. Tridiagonal solvers with multiple right hand sides on k-dimensional mesh and torus interconnection networks. Parallel Processing Letters, 13 (4):659–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Sarnath:1992:DLH

[Sar92] R. Sarnath. DTML is logspace hard under NC^{1} reductions. Parallel Processing Letters, 2 (2-3):189–193, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Saxena:2020:ANS

[Sax20] Sanjeev Saxena. All nearest smallers made simple. Parallel Processing Letters, 30 (02):??, June 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500085.

Senger:2012:BSB

[SB12] Hermes Senger and Fabrício Alves Barbosa Da Silva. Bounds on the scalability of bag-of-tasks applications running on masterslave platforms. Parallel Processing Letters, 22(2):1250004, June 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Sharma:2015:DQD

[SB15] Gokarna Sharma and Costas Busch. Distributed queuing in dynamic networks. Parallel Processing Letters, 25(2):1550005, June 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Saxena:1994:PPC

[SBP94] Sanjeev Saxena, P. C. P. Bhatt, and V. C. Prasad. On parallel prefix computation. Parallel Processing Letters, 4(4):429– 436, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Sapio:2017:PCA

[SBR^{+}17] Amedeo Sapio, Mario Baldi, Fulvio Risso, Narendra Anand, and Antonio Nucci. Packet capture and analysis on MED-INA, a massively distributed network data caching platform. Parallel Processing Letters, 27 (3–4):1750010, 2017. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Scheiman:1996:PDR

[SC96] C. Scheiman and P. Cappello.

A processor-time-minimal design for 3D rectilinear mesh algorithms. Parallel Processing Letters, 6(4):539–550, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Siegelin:2001:SCD

[SCF01] C. Siegelin, L. Castillo, and U. Finger. Smart Cards: Distributed computing with $5 devices. Parallel Processing Letters, 11(1):57–??, March 2001. CODEN PPLTEE. ISSN 0129-6264.

Schwabe:1993:OVD

[Sch93] E. J. Schwabe. Optimality of a VLSI decomposition scheme for the deBruijn graph. Parallel Processing Letters, 3(3):261– 265, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Schreiber:1997:HPF

[Sch97] Robert Schreiber. High Performance Fortran, Version 2. Parallel Processing Letters, 7(4):437– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Schikuta:2005:TXB

[Sch05] Erich Schikuta. Towards an XML based datagrid description language. Parallel Processing Letters, 15(4):379–386, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Schulz:2009:HFI

[Sch09] Matthias Schulz. How far is it to the next recurrent configuration? an NP-complete problem in the Sandpile Model. Parallel Processing Letters, 19(2): 265–281, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Schumann:2015:TCB

[Sch15] Andrew Schumann. Towards context-based concurrent formal theories. Parallel Processing Letters, 25(1):1540008, March 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Silveira:2010:IBS

[SCJ^{+}10] César L. B. Silveira, Gerson Geraldo H. Cavalheiro, Cláudio R. Jung, Julio C. S. Jacques, Jr., and Soraia R. Musse. An improved background subtraction algorithm and concurrent implementations. Parallel Processing Letters, 20(1):71–89, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Sudo:2018:CSS

[SDLM18] Yuichi Sudo, Ajoy K. Datta, Lawrence L. Larmore, and Toshimitsu Masuzawa. Constant space self-stabilizing center finding algorithms in chains and trees. Parallel Processing Letters, 28(1):1850002, March 2018. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:2001:TTD

[Ser01] J. Serot. Tagged-token data-flow for skeletons. Parallel Processing Letters, 11(4):377–??, December 2001. CODEN PPLTEE. ISSN 0129-6264.

Sun:1991:PAS

[SF91] K. T. Sun and H. C. Fu. An O(n) parallel algorithm for solving the traffic control problem on crossbar switch networks. Parallel Processing Letters, 1(1):51– 58, September 1991. CODEN PPLTEE. ISSN 0129-6264.

Schubert:2011:HPS

[SFHW11] Gerald Schubert, Holger Fehske, Georg Hager, and Gerhard Wellein. Hybrid-parallel sparse matrix-vector multiplication with explicit communication overlap on current multicorebased systems. Parallel Processing Letters, 21(3):339–358, September 2011. CODEN PPLTEE. ISSN 0129-6264.

Sheu:1999:FDS

[SH99] Jyh-Jian Sheu and Lih-Hsing Hsu. Fault diameter for supercubes. Parallel Processing Letters, 9(1):21–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Steuwer:2014:HLP

[SHBG14] Michel Steuwer, Michael Haidl, Stefan Breuer, and Sergei Gorlatch. High-level programming of stencil computations on multi-GPU systems using the SkelCL library. Parallel Processing Letters, 24(3): 1441005, September 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Sibeyn:2002:OLR

[Sib02] Jop F. Sibeyn. Optimal list ranking on one-dimensional arrays. Parallel Processing Letters, 12(3–4):375–??, September– December 2002. CODEN PPLTEE. ISSN 0129-6264.

Siegelmann:1996:NSC

[Sie96] H. T. Siegelmann. On NIL: the software constructor of neural networks. Parallel Processing Letters, 6(4):575–582, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Singh:1996:BTP

[Sin96] A. K. Singh. Bounded timestamps in process networks. Parallel Processing Letters, 6(2): 259–264, June 1996. CODEN PPLTEE. ISSN 0129-6264.

Swift:1994:PCL

[SJL94] L. K. Swift, T. Johnson, and P. E. Livadas. Parallel creation of linear octrees from quadtree slices. Parallel Processing Letters, 4(4):447–453, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Sancho:2010:PTI

[SKL10] Jose Carlos Sancho, Darren J. Kerbyson, and Michael Lang. On the performance and technological impact of adding memory controllers in multi-core processors. Parallel Processing Letters, 20(4):341–357, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Soman:2010:SGA

[SKN10] Jyothish Soman, Kishore Kothapalli, and P. J. Narayanan. Some GPU algorithms for graph connected components and spanning tree. Parallel Processing Letters, 20(4):325–339, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Shin:2004:NAD

[SLG04] Chulho Shin, Seong-Won Lee, and Jean-Luc Gaudiot. The need for adaptive dynamic thread scheduling in simultaneous multithreading. Parallel Processing Letters, 14(3/4):327– ??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264.

Sudo:2020:LER

[SM20] Yuichi Sudo and Toshimitsu Masuzawa. Leader election requires logarithmic time in population protocols. Parallel Processing Letters, 30(01):??, March 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S012962642050005X.

Scholz:2003:PPR

[SMH03] Bernhard Scholz, Eduard Mehofer, and Nigel Horspool. Predicated partial redundancy elimination using a cost analysis. Parallel Processing Letters, 13(4):525– ??, December 2003. CODEN PPLTEE. ISSN 0129-6264.

Sanders:1998:DAL

[SMK98] Beverly A. Sanders, Berna L.

Massingill, and Svetlana Kryukova.

Derivation of an algorithm for location management for mobile communication devices. Parallel Processing Letters, 8(4):473– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Szymanski:1995:SPR

[SMS95] B. K. Szymanski, W. Maniatty, and B. Sinharoy. Simultaneous parallel reduction on SIMD machines. Parallel Processing Letters, 5(3):437–449, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Soliman:2009:EIT

[Sol09a] Mostafa I. Soliman. Exploiting ILP, TLP, and DLP to improve multi-core performance of onesided Jacobi SVD. Parallel Processing Letters, 19(2):355–375, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Soliman:2009:PEM

[Sol09b] Mostafa I. Soliman. Performance evaluation of multi-core Intel Xeon processors on Basic Linear Algebra Subprograms. Parallel Processing Letters, 19 (1):159–174, March 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Stefanescu:2004:GEN

[SPA04] Radu Stefanescu, Xavier Pennec, and Nicholas Ayache. Gridenabled non-rigid registration of medical images. Parallel Processing Letters, 14(2):197–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Srimani:1996:SRI

[Sri96] P. K. Srimani. Super rotator: incrementally extensible directed network graph of sublogarithmic diameter. Parallel Processing Letters, 6(4):479–490, December 1996. CODEN PPLTEE. ISSN 0129-6264.

Sur:1992:SDA

[SS92] S. Sur and P. K. Srimani. A selfstabilizing distributed algorithm to construct BFS spanning trees of a symmetric graph. Parallel Processing Letters, 2(2-3):171– 179, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Sibeyn:2003:OGC

[SŠ03] Jop F. Sibeyn and Michal Šoch.

Optimal gossiping on CCCs of even dimension. Parallel Processing Letters, 13(1):35–??, March 2003. CODEN PPLTEE. ISSN 0129-6264.

Soliman:2005:PEB

[SS05] Mostafa I. Soliman and Stanislav G.

Sedukhin. Performance evaluation of BLAS on the Trident processor. Parallel Processing Letters, 15(4):407–414, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Slawinski:2013:TCU

[SS13] Jaroslaw Slawinski and Vaidy Sunderam. Towards computing as a utility via adaptive middleware: an experiment in cross-paradigm execution. Parallel Processing Letters, 23(2):

1340002, June 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Schmitt:2018:RHG

[SSK^{+}18] Christian Schmitt, Moritz Schmid, Sebastian Kuckuk, Harald Köstler, Jürgen Teich, and Frank Hannig. Reconfigurable hardware generation of multigrid solvers with conjugate gradient coarse-grid solution. Parallel Processing Letters, 28(04): ??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626418500160.

Schmeck:1993:PHT

[SSSM93] H. Schmeck, H. Schroder, J. Staunstrup, and M. Middendorf. Problem heaps and their evaluation. Parallel Processing Letters, 3(2):147–155, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Stacho:2004:RBC

[SŠZ04] Ladislav Stacho, Jozef Širáň, and Sanming Zhou. Routing balanced communications on Hamilton decomposable networks. Parallel Processing Letters, 14(3/4):377–??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264.

Spirakis:1997:PGH

[ST97] Paul Spirakis and Vassilis Triantafillou. Pure greedy hotpotato routing in the 2-D mesh with random destinations. Parallel Processing Letters, 7(3):

249–??, September 1997. CO-DEN PPLTEE. ISSN 0129-6264.

Salem:2012:CBP

[ST12] Fatima K. Abu Salem and Lama B. Tamim. Communication balancing in the parallel Göttfert algorithm. Parallel Processing Letters, 22(4): 1250011, December 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Siebert:2014:PLB

[ST14] Christian Siebert and Jesper Larsson Träff. Perfectly load-balanced, stable, synchronization-free parallel merge. Parallel Processing Letters, 24(1):1450005, March 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Stalzer:1995:PFM

[Sta95] Mark A. Stalzer. A parallel fast multipole method for the Helmholtz equation. Parallel Processing Letters, 5(2): 263–274, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Stannett:2012:GEN

[Sta12] Mike Stannett. Guest Editor’s note: Special issue on hypercomputation, physics and computation. Parallel Processing Letters, 22(3):1202003, September 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Stewart:2012:CCR

[Ste12] Iain A. Stewart. On the computational complexity of routing in faulty K-ary N-cubes and hypercubes. Parallel Processing Letters, 22(1):1250003, March 2012. CODEN PPLTEE. ISSN 0129-6264.

Stojmenovic:1996:GNR

[Sto96] Ivan Stojmenovic. Generating nary reflected Gray codes on a linear array of processors. Parallel Processing Letters, 6(1):27–34, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Suel:1995:PRS

[Sue95] T. Suel. Permutation routing and sorting on meshes with row and column buses. Parallel Processing Letters, 5(1):63–80, March 1995. CODEN PPLTEE. ISSN 0129-6264.

Suijlen:2017:MBT

[Sui17] Wijnand Suijlen. Mock BSPlib for testing and debugging bulk synchronous parallel software. Parallel Processing Letters, 27 (1):1740001, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Subramonian:1993:EMB

[SV93] R. Subramonian and N. Venkatasubramanyan. Efficient multipleitem broadcast in the logP model. Parallel Processing Letters, 3(4):407–417, December 1993. CODEN PPLTEE. ISSN 0129-6264.

Sengupta:2000:DFF

[SV00] A. Sengupta and S. Viswanathan.

Deadlock-free fault-tolerant multicast routing in hypercubes. Parallel Processing Letters, 10 (4):327–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1004/S0129626400000317.
html.

Sieling:1993:NOB

[SW93] D. Sieling and I. Wegener. NCalgorithms for operations on binary decision diagrams. Parallel Processing Letters, 3(1):3–12, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Schaffer:2008:UHM

[SW08] Kevin Schaffer and Robert A.

Walker. Using hardware multithreading to overcome broadcast/reduction latency in an associative SIMD processor. Parallel Processing Letters, 18(4): 491–509, December 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Shahzad:2013:SCR

[SWK^{+}13] Faisal Shahzad, Markus Wittmann, Moritz Kreutzer, Thomas Zeiser, Georg Hager, and Gerhard Wellein. A survey of checkpoint/restart techniques on distributed memory systems. Parallel Processing Letters, 23(4): 1340011, December 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Serna:2002:PAF

[SX02] M. Serna and F. Xhafa. The parallel approximability of the false and true gates problems for NOR-circuits. Parallel Processing Letters, 12(1):127–??, March 2002. CODEN PPLTEE. ISSN 0129-6264.

Shirazi:1995:ESS

[SYL95] B. Shirazi, Hee Yong Youn, and D. M. Lorts. Evaluation of static scheduling heuristics for realtime multiprocessing. Parallel Processing Letters, 5(4):599– 610, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Shantharam:2013:SAC

[SYR13a] Manu Shantharam, Youngtae Youn, and Padma Raghavan. Speedup-aware co-schedules for efficient workload management. Parallel Processing Letters, 23 (2):1340001, June 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Syropoulos:2013:HFR

[Syr13b] Apostolos Syropoulos. Hypercomputation: Fantasy or reality? a position paper. Parallel Processing Letters, 23(1): 1350005, March 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Szabo:2012:MFP

[Sza12] László E. Szabó. Mathematical facts in a physicalist ontology. Parallel Processing Letters, 22(3):1240009, September 2012.

CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Spirov:2014:EDG

[SZH14] A. V. Spirov, E. A. Zagriychuk, and D. M. Holloway. Evolutionary design of gene networks: Forced evolution by genomic parasites. Parallel Processing Letters, 24(2):1440004, June 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Sun:2011:PAP

[SZJK11] Yanhua Sun, Gengbin Zheng, Pritish Jetley, and Laxmikant V. Kale. Parssse: an adaptive parallel state space search engine. Parallel Processing Letters, 21(3):319–338, September 2011. CODEN PPLTEE. ISSN 0129-6264.

Thomsen:2009:GEN

[TA09a] Michael Kirkedal Thomsen and Holger Bock Axelsen. Guest Editor’s note. Parallel Processing Letters, 19(2):225, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Thomsen:2009:PRR

[TA09b] Michael Kirkedal Thomsen and Holger Bock Axelsen. Parallelization of reversible ripplecarry adders. Parallel Processing Letters, 19(2):205–222, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Tang:2009:SPA

[Tan09] Joseph Tang. A simple parallel adaptive mesh CFD method suitable for small engineering workstations. Parallel Processing Letters, 19(3):469–476, September 2009. CODEN PPLTEE. ISSN 0129-6264.

Trahan:1994:PRA

[TB94] Jerry L. Trahan and Hosangadi Bhanukumar. Parallel Random Access Machines without Boolean operations. Parallel Processing Letters, 4(1-2): 117–124, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Tassone:2018:IAS

[TB18] Joseph Tassone and Mike Biocchi. The importance of applying security practices in wireless communication: Bluetooth low energy and RFID. Parallel Processing Letters, 28(03): ??, September 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S012962641850010X.

Trahan:1998:TBC

[TBV98] Jerry L. Trahan, Anu G. Bourgeois, and Ramachandran Vaidyanathan. Tighter and broader complexity results for reconfigurable models. Parallel Processing Letters, 8(3):271– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Trefethen:1999:LIC

[TDS99] Anne E. Trefethen, Mishi Derakhshan, and Stefano Salvini.

Libraries and infrastructure for clustered computing. Parallel Processing Letters, 9(2):303–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Touati:2004:EPR

[TE04] Sid-Ahmed-Ali Touati and Christine Eisenbeis. Early periodic register allocation on ILP processors. Parallel Processing Letters, 14(2):287–??, June 2004. CODEN PPLTEE. ISSN 0129-6264.

Tel:1995:LEH

[Tel95] G. Tel. Linear election in hypercubes. Parallel Processing Letters, 5(3):357–366, September 1995. CODEN PPLTEE. ISSN 0129-6264.

Tabirca:2005:AFG

[TFTY05] Tatiana Tabirca, Len Freeman, Sabin Tabirca, and Laurence Tianruo Yang. Applying the feedback guided dynamic loop scheduling method for the shortest path problem. Parallel Processing Letters, 15(4):491– 497, December 2005. CODEN PPLTEE. ISSN 0129-6264.

Tsai:1994:SSA

[TH94] Ming-Shin Tsai and Shing-Tsaan Huang. A self-stabilizing algorithm for the shortest paths problem with a fully distributed demon. Parallel Processing Letters, 4(1-2):65–72, June 1994. CODEN PPLTEE. ISSN 0129-

Thabtah:2013:MAM

[TH13] Fadi Thabtah and Suhel Hammoud. MR-ARM: a Map-Reduce association rule mining framework. Parallel Processing Letters, 23(3):1350012, September 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Thabtah:2015:PAC

[THAJ15] Fadi Thabtah, Suhel Hammoud, and Hussein Abdel-Jaber. Parallel associative classification data mining frameworks based MapReduce. Parallel Processing Letters, 25(2):1550002, June 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Thornton:2003:PEP

[Tho03] Mitchell Thornton. Performance evaluation of a parallel decoupled data driven multiprocessor. Parallel Processing Letters, 13(3):497–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Tiskin:2001:NWD

[Tis01] A. Tiskin. A new way to divide and conquer. Parallel Processing Letters, 11(4):409–??, December 2001. CODEN PPLTEE. ISSN 0129-6264.

Tatas:2008:RPD

[TKE^{+}08] Konstantinos Tatas, Costas Kyriacou, Paraskevas Evripidou, Pedro Trancoso, and Stephan Wong. Rapid prototyping of the data-driven chipmultiprocessor (D^{2}-CMP) using FPGAs. Parallel Processing Letters, 18(2):291–306, June 2008. CODEN PPLTEE. ISSN 0129-6264.

Tan:1991:PMJ

[TL91] Kian-Lee Tan and Hongjun Lu. On processing multi-joins in parallel systems. Parallel Processing Letters, 1(2):157– 164, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Tzeng:2011:BVM

[TLH11] Shiang-Feng Tzeng, Cheng-Chi Lee, and Min-Shiang Hwang. A batch verification for multiple proxy signature. Parallel Processing Letters, 21(1):77–84, March 2011. CODEN PPLTEE. ISSN 0129-6264.

Tachon:2017:AGB

[TLHL17] Thibaut Tachon, Chong Li, Gaétan Hains, and Frédéric Loulergue. Automated generation of BSP automata. Parallel Processing Letters, 27(1): 1740002, March 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Thangavel:1993:PAG

[TM93] P. Thangavel and V. P. Muthuswamy.

A parallel algorithm to generate N-ary reflected Gray codes in a linear array with reconfigurable bus system. Parallel Processing Letters, 3(2):157–164, June 1993. CODEN PPLTEE. ISSN 0129-6264.

Trefftz:2016:ECE

[TMKS16] Christian Trefftz, Hugh McGuire, Zachary Kurmas, and Jerry Scripps. Exhaustive community enumeration in parallel. Parallel Processing Letters, 26 (2):1650006, June 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Tian:2020:CBG

[TMW20] Yingzhi Tian, Huaping Ma, and Liyun Wu. The connectivity of a bipartite graph and its bipartite complementary graph. Parallel Processing Letters, 30(03):??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420400058.

Torbey:2009:TFI

[Tor09] Sami Torbey. Towards a framework for intuitive programming of cellular automata. Parallel Processing Letters, 19(1):73–83, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Traff:2009:RBR

[Trä09] Jesper Larsson Träff. Relationships between regular and irregular collective communication operations on clustered multiprocessors. Parallel Processing Letters, 19(1):85–96, March 2009. CODEN PPLTEE. ISSN 0129-6264.

Tromp:1993:US

[Tro93]  J. Tromp.  On update-last

schemes. Parallel Processing Letters, 3(1):25–28, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Tsay:2004:DSA

[Tsa04] Jong-Chuang Tsay. Designing a systolic algorithm for generating well-formed parenthesis strings. Parallel Processing Letters, 14 (1):83–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Thomas:2006:AHL

[TSS^{+}06] Nathan Thomas, Steven Saunders, Tim Smith, Gabriel Tanase, and Lawrence Rauchwerger. ARMI: A high level communication library for STAPL. Parallel Processing Letters, 16 (2):261–280, June 2006. CO-DEN PPLTEE. ISSN 0129-6264.

Talia:2005:APD

[TT05] Domenico Talia and Paolo Trunfio. Adapting a pure decentralized peer-to-peer protocol for Grid services invocation. Parallel Processing Letters, 15(1/2): 67–84, March/June 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Tamassia:2000:PAP

[TTV00] R. Tamassia, I. G. Tollis, and J. S. Vitter. A parallel algorithm for planar orthogonal grid drawings. Parallel Processing Letters, 10(1):141–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S0129626400000147.
html.

Turau:2010:SSV

[Tur10] Volker Turau. Self-stabilizing vertex cover in anonymous networks with optimal approximation ratio. Parallel Processing Letters, 20(2):173–186, June 2010. CODEN PPLTEE. ISSN 0129-6264.

Tourancheau:2001:SMM

[TW01] B. Tourancheau and R. Westrelin. Study of the medium message performance of BIP/ Myrinet. Parallel Processing Letters, 11(2–3):297–??, 2001. CODEN PPLTEE. ISSN 0129-6264.

Trevisan:1998:PCP

[TX98] Luca Trevisan and Fatos Xhafa. The parallel complexity of positive linear programming. Parallel Processing Letters, 8(4):527– ??, December 1998. CODEN PPLTEE. ISSN 0129-6264.

Umeo:2009:FSS

[UKY09] Hiroshi Umeo, Naoki Kamikawa, and Jean-Baptiste Yunès. A family of smallest symmetrical four-state firing squad synchronization protocols for ring arrays. Parallel Processing Letters, 19(2):299–313, June 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Ullman:1992:MBP

[UN92] S. Ullman and B. Narahari. Mapping binary precedence trees to hypercubes. Parallel Processing Letters, 2(1):81–87, March 1992. CODEN PPLTEE. ISSN 0129-6264.

Uddin:2011:HLS

[UVJ11] M. Irfan Uddin, Michiel W. Van Tol, and Chris R. Jesshope. High level simulation of SVP many-core systems. Parallel Processing Letters, 21(4):413– 438, December 2011. CODEN PPLTEE. ISSN 0129-6264.

vanDongen:1994:CDL

[van94] Vincent van Dongen. Compiling distributed loops onto SPMD code. Parallel Processing Letters, 4(3):301–312, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Valero:1994:NSO

[VAP94] Mateo Valero, Eduard Ayguade, and Montse Peiron. Network synchronization and outof-order access to vectors. Parallel Processing Letters, 4(4):405– 415, December 1994. CODEN PPLTEE. ISSN 0129-6264.

Vasilev:2003:BVR

[Vas03] Vasil P. Vasilev. BSPGRID:

Variable resources parallel computation and multiprogrammed parallelism. Parallel Processing Letters, 13(3):329–??, September 2003. CODEN PPLTEE. ISSN 0129-6264.

Vadhiyar:2003:SFD

[VD03] Sathish S. Vadhiyar and Jack J.

Dongarra. SRS: a framework for developing malleable and migratable parallel applications for distributed systems. Parallel Processing Letters, 13(2): 291–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Villebonnet:2015:BML

[VDL^{+}15] Violaine Villebonnet, Georges Da Costa, Laurent Lefevre, Jean-Marc Pierson, and Patricia Stolf. “Big, medium, little”: Reaching energy proportionality with heterogeneous computing scheduler. Parallel Processing Letters, 25(3):1541006, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Vee:2000:LPL

[VH00] V.-Y. Vee and W.-J. Hsu. Locality-preserving load-balancing mechanisms for synchronous simulations on shared-memory multiprocessors. Parallel Processing Letters, 10(1):111–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1001/S0129626400000123.
                                        html.

Vee:2001:SES

[VH01] V.-Y. Vee and W.-J. Hsu. A scalable and efficient storage allocator on shared-memory multiprocessors. Parallel Processing Letters, 11(2–3):313–??, 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Vidyasankar:2006:HCG

[Vid06]  K. Vidyasankar.  A highly

concurrent group mutual Lexclusion algorithm. Parallel Processing Letters, 16(4):467– 483, December 2006. CODEN PPLTEE. ISSN 0129-6264.

Violard:1994:MTE

[Vio94] Eric Violard. A mathematical theory and its environment for parallel programming. Parallel Processing Letters, 4(3):313– 328, September 1994. CODEN PPLTEE. ISSN 0129-6264.

Virbitskaite:1993:SCN

[Vir93] I. B. Virbitskaite. Some characteristics of nondeterministic processes. Parallel Processing Letters, 3(1):99–106, March 1993. CODEN PPLTEE. ISSN 0129-6264.

Vaidehi:1999:AGA

[VKS99] V. Vaidehi, C. N. Krishnan, and P. Swaminathan. An aided genetic algorithm for multiprocessor scheduling. Parallel Processing Letters, 9(3):423– ??, September 1999. CODEN PPLTEE. ISSN 0129-6264.

Valero:1991:CSV

[VLL^{+}91] M. Valero, T. Lang, J. M.

Llaberia, M. Peiron, Navarro, J. J., and E. Ayguade. Conflictfree strides for vectors in matched memories. Parallel Processing Letters, 1(2):95– 102, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Vallee:2003:CSS

[VLR^{+}03] Geoffroy Vallée, Renaud Lottiaux, Louis Rilling, Jean-Yves Berthou, Ivan Dutka Malhen, and Christine Morin. A case for single system image cluster operating systems: the Kerrighed approach. Parallel Processing Letters, 13(2):95–??, June 2003. CODEN PPLTEE. ISSN 0129-6264.

Varghese:2011:CAI

[VMA11] Blesson Varghese, Gerard Mckee, and Vassil Alexandrov. Can agent intelligence be used to achieve fault tolerant parallel computing systems? Parallel Processing Letters, 21(4): 379–396, December 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Veglis:1996:MBA

[VP96] A. A. Veglis and A. S. Pombortsis. Memory bandwidth analysis of shared-memory multiprocessors using dynamic request rate. Parallel Processing Letters, 6(1): 67–74, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Vajnovszki:1998:PAL

[VP98] Vincent Vajnovszki and Jean Pallo. Parallel algorithms for listing well-formed parentheses strings. Parallel Processing Letters, 8(1):19–??, March 1998. CODEN PPLTEE. ISSN 0129-6264.

Vajnovszki:1999:SGK

[VP99] Vincent Vajnovszki and Chris Phillips. Systolic generation of k-ary trees. Parallel Processing Letters, 9(1):93–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Vargas:2020:EPF

[VRR20] Karla Vargas, Sergio Rajsbaum, and Michel Raynal. An eventually perfect failure detector for networks of arbitrary topology connected with ADD channels using time-to-live values. Parallel Processing Letters, 30(02):??, June 2020. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500061.

Venkatachar:1997:CGB

[VRT97] A. Venkatachar, J. Ramanujam, and A. Thirumalai. Communication generation for blockcyclic distributions. Parallel Processing Letters, 7(2):195– 202, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Venema:2000:NAS

[VSS00] S. Venema, H. Shen, and F. Suraweera. NC algorithms for the single most vital edge problem with respect to all pairs shortest paths. Parallel Processing Letters, 10(1):51–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

ejournals.wspc.com.sg/ppl/
10/1001/S012962640000007X.
html.

Wang:1993:POR

[Wan93] P. S. P. Wang. Parallel object representation and recognition. Parallel Processing Letters, 3(3): 279–290, September 1993. CO-DEN PPLTEE. ISSN 0129-6264.

Wang:1996:BAB

[Wan96] Bung-Feng Wang. A better analysis of Ben-Asher’s algorithm for the conditional Cartesian product problem. Parallel Processing Letters, 6(3):331–344, September 1996. CODEN PPLTEE. ISSN 0129-6264.

Weinzierl:2014:BFD

[WBUW14] Tobias Weinzierl, Michael Bader, Kristof Unterweger, and Roland Wittmann. Block fusion on dynamically adaptive spacetree grids for shallow water waves. Parallel Processing Letters, 24 (3):1441006, September 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Wu:2003:QRH

[WD03] Jie Wu and Fei Dai. QoS routing in hypercube multicomputers. Parallel Processing Letters, 13(1):43–??, March 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Wylie:2010:LSP

[WGM^{+}10] Brian J. N. Wylie, Markus Geimer, Bernd Mohr, David BöHme, ZoltáN Szebenyi, and Felix Wolf. Large-scale performance analysis of Sweep3D with the Scalasca Toolset. Parallel Processing Letters, 20(4): 397–414, December 2010. CO-DEN PPLTEE. ISSN 0129-6264.

Wittmann:2010:LSC

[WHTW10] Markus Wittmann, Georg Hager, Jan Treibig, and Gerhard Wellein. Leveraging shared caches for parallel temporal blocking of stencil codes on multicore processors and clusters. Parallel Processing Letters, 20(4):359–376, December 2010.

CODEN PPLTEE. ISSN 0129-6264.

Wei:2012:OLL

[WJ12] Zheng Wei and Joseph Jaja. Optimization of linked list prefix computations on multithreaded GPUs using CUDA. Parallel Processing Letters, 22(4): 1250012, December 2012. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Wu:2016:ANG

[WJ16] Jing Wu and Joseph Jaja.

Achieving native GPU performance for out-of-card large dense matrix multiplication. Parallel Processing Letters, 26 (2):1650007, June 2016. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Weems:2011:GEN

[WKR11] Charles C. Weems, Darren J.

Kerbyson, and Ram Rajamony. Guest Editor’s note: Largescale parallel processing. Parallel Processing Letters, 21(3): 275–277, September 2011. CO-DEN PPLTEE. ISSN 0129-6264.

Weems:2013:GEN

[WKR13] Charles C. Weems, Darren J.

Kerbyson, and Ram Rajamony. Guest Editors’ note: Largescale parallel processing. Parallel Processing Letters, 23(4): 1302002, December 2013. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Weems:2014:GEN

[WKR14] Charles C. Weems, Darren J. Kerbyson, and Ram Rajamony. Guest Editors’ note: Special issue on large-scale parallel processing. Parallel Processing Letters, 24(4):1402003, December 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Weems:2010:GEN

[WKRJ10] Charles C. Weems, Darren J. Kerbyson, Ram Rajamony, and Alex K. Jones. Guest Editor’s note: Large-scale parallel processing. Parallel Processing Letters, 20(4):289–291, December 2010. CODEN PPLTEE. ISSN 0129-6264.

Wang:1995:DSU

[WLR95] Jhy-Chun Wang, Tseng-Hui Lin, and S. Ranka. Distributed scheduling of unstructured collective communication on the CM-5. Parallel Processing Letters, 5(4):647–658, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Walker:2011:APG

[WLW11] Coral Walker, Dashan Lu, and David W. Walker. Automatic portal generation based on workflow description. Parallel Processing Letters, 21(2): 155–171, June 2011. CODEN PPLTEE. ISSN 0129-6264.

Wolski:1995:SSH

[Wol95] R. Wolski. Static scheduling of hierarchical program graphs.

Parallel Processing Letters, 5(4): 611–622, December 1995. CO-DEN PPLTEE. ISSN 0129-6264.

Wolski:1999:PCA

[Wol99] Rich Wolski. Predicting CPU availability on the Computational Grid using the Network Weather Service. Parallel Processing Letters, 9(2):227–??, June 1999. CODEN PPLTEE. ISSN 0129-6264.

Wilde:1997:MRA

[WR97] D. Wilde and S. Rajopadhye.

Memory reuse analysis in the polyhedral model. Parallel Processing Letters, 7(2):203–215, June 1997. CODEN PPLTEE. ISSN 0129-6264.

Wu:2003:DFR

[WS03] Jie Wu and Li Sheng. Deadlockfree routing in irregular networks using prefix routing. Parallel Processing Letters, 13(4): 705–??, December 2003. CO-DEN PPLTEE. ISSN 0129-6264.

Wang:2016:OSA

[WSK16] Yang Wang, Wei Shi, and Kenneth B. Kent. On optimal scheduling algorithms for well-structured workflows in the cloud with budget and deadline constraints. Parallel Processing Letters, 26(2):1650009, June 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Wu:1995:TBN

[Wu95a] Jie Wu. Tight bounds on the number of ℓ-nodes in a faulty hypercube. Parallel Processing Letters, 5(2):321–328, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Wu:1995:PIS

[Wu95b]  Min-You Wu.  Parallel in-

cremental scheduling. Parallel Processing Letters, 5(4):659– 670, December 1995. CODEN PPLTEE. ISSN 0129-6264.

Wu:1996:PSA

[Wu96] Min-You Wu. A parallel scheduling algorithm for tree structured interconnection networks. Parallel Processing Letters, 6(1): 45–54, March 1996. CODEN PPLTEE. ISSN 0129-6264.

Wu:1998:FTC

[Wu98] Jie Wu. Fault-tolerant communications in cube-based multiple-bus systems. Parallel Processing Letters, 8(3):307– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264.

Wu:2000:FHP

[Wu00] J. Wu. On finding a Hamiltonian path in a tournament using semi-heap. Parallel Processing Letters, 10(4):279–??, December 2000. CODEN PPLTEE. ISSN 0129-6264. URL http://

                                        ejournals.wspc.com.sg/ppl/
                                        10/1004/S0129626400000275.
                                        html.

Wuensche:2009:CAE

[Wue09] Andrew Wuensche. Cellular Automata Encryption: the reverse algorithm, Z-parameter and chain-rules. Parallel Processing Letters, 19(2):283–297, June 2009. CODEN PPLTEE. ISSN 0129-6264.

Watts:1995:PBM

[WV95] Jerrell Watts and Robert Van De Geijn. A pipelined broadcast for multidimensional meshes. Parallel Processing Letters, 5(2): 281–292, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Wang:2019:NCA

[WW19] Shiying Wang and Mujiangshan Wang. A note on the connectivity of m-ary ndimensional hypercubes. Parallel Processing Letters, 29(04): ??, December 2019. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626419500178.

Wang:2020:CDL

[WXW20] Mujiangshan Wang, Dong Xiang, and Shiying Wang. Connectivity and diagnosability of leaf-sort graphs. Parallel Processing Letters, 30(03): ??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400046.

Wong:1992:MSP

[WY92] W. F. Wong and C. K. Yuen.

A model of speculative parallelism. Parallel Processing Letters, 2(2-3):265–272, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Xu:1994:OPL

[XL94] Cheng-Zhong Xu and Francis C. M. Lau. Optimal parameters for load balancing with the diffusion method in mesh networks. Parallel Processing Letters, 4(1-2):139–147, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Xiao:2012:CDC

[XLP12] Wenjun Xiao, Huomin Liang, and Behrooz Parhami. A class of data-center network models offering symmetry, scalability, and reliability. Parallel Processing Letters, 22(4):1250013, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Xiang:1999:DDB

[XU99] Limin Xiang and Kazuo Ushijima. Decoding and drawing on BSR for a binary tree from its I-P sequence. Parallel Processing Letters, 9(1):103–??, March 1999. CODEN PPLTEE. ISSN 0129-6264.

Xue:1991:SCS

[Xue91] Jingling Xue. Specifying control signals for systolic arrays by uniform recurrence equations. Parallel Processing Letters, 1(2):83– 93, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Xue:1996:GUA

[Xue96] Jingling Xue. Generalising the unimodular approach to restructure imperfectly nested loops. Parallel Processing Letters, 6(3):

401–414, September 1996. CO-DEN PPLTEE. ISSN 0129-6264.

Xue:1997:TLT

[Xue97] Jingling Xue. On tiling as a loop transformation. Parallel Processing Letters, 7(4):409– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Xiang:2002:TOS

[XUZ02] Limin Xiang, Kazuo Ushijima, and Jianjun Zhao. Time optimal n-size matching parentheses and binary tree decoding algorithms on a p-processor BSR. Parallel Processing Letters, 12(3–4):365– ??, September–December 2002. CODEN PPLTEE. ISSN 0129-6264.

Xirouchakis:1993:DPV

[XWF93] P. C. Xirouchakis, P. Y. Wang, and O. Frieder. Data parallel visual reconstruction and partitioning algorithms. Parallel Processing Letters, 3(3):267– 277, September 1993. CODEN PPLTEE. ISSN 0129-6264.

Xu:2020:PQA

[XX20] Guanlei Xu and Xiaogang Xu.

The parallel quantum algorithm for the class of optimization. Parallel Processing Letters, 30(04):??, December 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500140.

Yaikhom:2008:MPM

[Yai08] Gagarine Yaikhom. Message passing without memory copy.

Parallel Processing Letters, 18 (1):87–100, March 2008. CO-DEN PPLTEE. ISSN 0129-6264.

Yang:2005:IPI

[Yan05] Laurence Tianruo Yang. The improved parallel ICGS method for large and sparse unsymmetric linear systems. Parallel Processing Letters, 15(4): 459–467, December 2005. CO-DEN PPLTEE. ISSN 0129-6264.

Yang:1991:IMM

[YDL91] S. B. Yang, S. K. Dhall, and S. Lakshmivarahan. An improved maximal matching algorithm. Parallel Processing Letters, 1(2):135–142, December 1991. CODEN PPLTEE. ISSN 0129-6264.

Yang:1994:PEC

[YDL94] S. B. Yang, S. K. Dhall, and S. Lakshmivarahan. A processor efficient connectivity algorithm on random graphs. Parallel Processing Letters, 4(1-2):29–36, June 1994. CODEN PPLTEE. ISSN 0129-6264.

Yan:2018:UCD

[YF18] Peizhi Yan and Yi Feng. Using convolution and deep learning in Gomoku game artificial intelligence. Parallel Processing Letters, 28(03):??, September 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626418500111.

Yuan:1997:DDD

[YGM97] Xin Yuan, Rajiv Gupta, and Rami Melhem. Demand-driven data flow analysis for communication optimization. Parallel Processing Letters, 7(4):359– ??, December 1997. CODEN PPLTEE. ISSN 0129-6264.

Yamazaki:2014:DIL

[YKLD14] Ichitaro Yamazaki, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Design and implementation of a large scale treebased QR decomposition using a 3D virtual systolic array and a lightweight runtime. Parallel Processing Letters, 24(4): 1442004, December 2014. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Yasim:2010:OSE

[YL10] Shameema Yasim and Shahram Latifi. Optimal subcube embeddability in hypercubes with additional dimensions. Parallel Processing Letters, 20(1):91–99, March 2010. CODEN PPLTEE. ISSN 0129-6264.

Yang:2004:STDa

[YMZL04a] Xiaofan Yang, Graham M. Megson, Shaomln Zhang, and Xiaoping Liu. A solution to the three disjoint path problem on honeycomb meshes. Parallel Processing Letters, 14(3/4):399–??, September/December 2004. CO-DEN PPLTEE. ISSN 0129-6264.

Yang:2004:STDb

[YMZL04b] Xiaofan Yang, Graham M. Mengson, Shaomin Zhang, and Xiaoping Liu. A solution to the three disjoint path problem on honeycomb tori. Parallel Processing Letters, 14(3/4):411–??, September/December 2004. CO-DEN PPLTEE. ISSN 0129-6264.

Yalcinkaya:2005:AVA

[YS05a] Yasemin Yalçinkaya and Trond Steihaug. Aggregating variables for asynchronous iterations. Parallel Processing Letters, 15(1/2):99–114, March/ June 2005. CODEN PPLTEE. ISSN 0129-6264.

Yalcinkaya:2005:EN

[YS05b] Yasemin Yalçinkaya and Trond Steihaug. Editorial note. Parallel Processing Letters, 15(1/2): 115, March/June 2005. CODEN PPLTEE. ISSN 0129-6264.

Yang:1992:PEM

[YTL92] C. S. Yang, Y. M. Tsai, and C. Y. Liu. Performance evaluation of multicast wormhole routing in 2D-Torus multicomputers. Parallel Processing Letters, 2(2-3):161–170, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Yong:2003:AMC

[YWJ03] Xie Yong and Hsu Wen-Jing. Aligned multithreaded computations and their scheduling with FAB performance guarantees. Parallel Processing Letters, 13(3):353–??, September 2003.

CODEN PPLTEE. ISSN 0129-6264.

Zaroliagis:1997:SWP

[Zar97] Christos D. Zaroliagis. Simple and work-efficient parallel algorithms for the minimum spanning tree problem. Parallel Processing Letters, 7(1):25–37, March 1997. CODEN PPLTEE. ISSN 0129-6264.

Zavanella:2001:SBP

[Zav01] A. Zavanella. Skeletons, BSP and performance portability. Parallel Processing Letters, 11 (4):393–??, December 2001. CO-DEN PPLTEE. ISSN 0129-6264.

Zeiser:2009:BAA

[ZHW09] Thomas Zeiser, Georg Hager, and Gerhard Wellein. Benchmark analysis and application results for lattice Boltzmann simulations on NEC SX Vector and Intel Nehalem systems. Parallel Processing Letters, 19(4): 491–511, December 2009. CO-DEN PPLTEE. ISSN 0129-6264.

Ziavras:1995:SMH

[Zia95] Sotirios G. Ziavras. Scalable multifolded hypercubes for versatile parallel computers. Parallel Processing Letters, 5(2): 241–250, June 1995. CODEN PPLTEE. ISSN 0129-6264.

Zhang:2009:NUS

[ZJ09] Yu Zhang and Alex K. Jones.

Non-uniform ‘fat-meshes’ for chip multiprocessors. Parallel Processing Letters, 19(4):595– 617, December 2009. CODEN PPLTEE. ISSN 0129-6264.

Zha:2020:RTL

[ZJSY20] Daolu Zha, Xi Jin, Rui Shang, and Pengfei Yang. A real-time learning-based superresolution system on FPGA. Parallel Processing Letters, 30(04):??, December 2020. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500115.

Zenil:2018:ACR

[ZKmST18] Hector Zenil, Narsis A. Kiani, Ming mei Shang, and Jesper Tegnér. Algorithmic complexity and reprogrammability of chemical structure networks. Parallel Processing Letters, 28(1): 1850005, March 2018. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Zhi:2004:UPS

[ZLL04] Xiaoli Zhi, Rong Lu, and Xinda Lu. Universal parallel solver for convection-diffusion equations. Parallel Processing Letters, 14(1):107–??, March 2004. CODEN PPLTEE. ISSN 0129-6264.

Zhang:2015:PIA

[ZMD15] Xiaoge Zhang, Sankaran Mahadevan, and Yong Deng. Physarum-inspired applications in graph-optimization problems. Parallel Processing Letters, 25 (1):1540005, March 2015. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Zhu:2020:FMP

[ZMZZ20] Bo Zhu, Tianlong Ma, Shuangshuang Zhang, and He Zhang. Fractional matching preclusion for data center networks. Parallel Processing Letters, 30 (02):??, June 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500103.

Zhu:2021:FMP

[ZMZZ21] Bo Zhu, Tianlong Ma, Shuangshuang Zhang, and He Zhang. Fractional matching preclusion for data center networks. Parallel Processing Letters, 31(03):??, September 2021. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500122.

Zhao:2013:PMP

[ZN13] Di Zhao and Shenghua Ni. Parallel multi-proposal and multichain MCMC for calculating pvalue of genome-wide association study. Parallel Processing Letters, 23(3):1350008, September 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Zanikolas:2008:IAA

[ZS08] Serafeim Zanikolas and Rizos Sakellariou. An importanceaware architecture for largescale Grid information services. Parallel Processing Letters, 18 (3):347–370, September 2008.

CODEN PPLTEE. ISSN 0129-6264.

Zheng:2018:SPC

[ZSLC18] Weijian Zheng, Fengguang Song, Lan Lin, and Zizhong Chen. Scaling up parallel computation of tiled QR factorizations by a distributed scheduling runtime system and analytical modeling. Parallel Processing Letters, 28 (1):1850004, March 2018. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Zhang:2021:ODI

[ZSLR21] Junwei Zhang, Li Shi, Yang Liu, and Thomas G. Robertazzi. Optimizing data intensive flows for networks on chips. Parallel Processing Letters, 31(03):??, September 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500134.

Zimeo:2008:CSC

[ZTP^{+}08] Eugenio Zimeo, Alberto Troisi, Harris Papadakis, Paraskevi Fragopoulou, Agostino Forestiero, and Carlo Mastroianni. Cooperative self-composition and discovery of Grid services in P2P networks. Parallel Processing Letters, 18(3):329–346, September 2008. CODEN PPLTEE. ISSN 0129-6264.

Zaichenkov:2016:CBC

[ZTS^{+}16] Pavel Zaichenkov, Olga Tveretina, Alex Shafarenko, Bert Gijsbers, and Clemens Grelck. The cost and benefits of coordination programming: Two case studies in concurrent collections and S-NET. Parallel Processing Letters, 26(3):1650011, September 2016. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Zhang:2020:SNS

[ZXLY20] Shuangshuang Zhang, Yuzhi Xiao, Xia Liu, and Jun Yin. A short note of strong matching preclusion for a class of arrangement graphs. Parallel Processing Letters, 30(01):??, March 2020. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420500012.

Zhai:2020:FTM

[ZXY20] Liyang Zhai, Liqiong Xu, and Weihua Yang. Fault-tolerant maximal local-edge-connectivity of augmented cubes. Parallel Processing Letters, 30(03): ??, September 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626420400010.

Zhai:2021:ECE

[ZXY21] Liyang Zhai, Liqiong Xu, and Shanshan Yin. On the 3extra connectivity of enhanced hypercubes. Parallel Processing Letters, 31(04):??, December 2021. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S012962642150016X.

Zhang:2020:RAG

[ZXYY20] Qifan Zhang, Liqiong Xu, Weihua Yang, and Shanshan Yin. Reliability analysis of the generalized exchanged hypercube. Parallel Processing Letters, 30 (02):??, June 2020. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626420500097.

Zeinalipour-Yazti:2008:MRP

[ZYPGD08] Demetrios Zeinalipour-Yazti, Harris Papadakis, Chryssis Georgiou, and Marios D. Dikaiakos. Metadata ranking and pruning for failure detection in Grids. Parallel Processing Letters, 18(3):371–390, September 2008. CODEN PPLTEE. ISSN 0129-6264.

Zhuang:2021:REG

[ZZL^{+}21] Hongbin Zhuang, Sunjian Zheng, Ximeng Liu, Cheng kuan Lin, and Xiaoyan Li. Reliability evaluation of generalized exchanged hypercubes based on imprecise diagnosis strategies. Parallel Processing Letters, 31(01):??, March 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0129626421500055.

Zhang:2021:FSM

[ZZZY21] He Zhang, Jinyu Zou, Shuangshuang Zhang, and Chengfu Ye. Fractional strong matching preclusion for DHcube. Parallel Processing Letters, 31(01):??, March 2021. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S0129626421500018.